### v2

# Extended-Krylov-subspace methods for trust-region and norm-regularization subproblems

Abstract
We consider an effective new method for solving trust-region and norm-regularization problems that arise as subproblems in many optimization applications. We show that the solutions to such subproblems lie on a manifold of approximately very low rank as a function of their controlling parameters (trust-region radius or regularization weight). Based on this, we build a basis for this manifold using an efficient extended-Krylov-subspace iteration that involves a single matrix factorization. The problems within the subspace using such a basis may be solved at very low cost using effective high-order root-finding methods. This then provides an alternative to common methods using multiple factorizations or standard Krylov subspaces. We provide numerical results to illustrate the effectiveness of our TREK/NREK approach.
1 Introduction
Given an $n$ by $n$ symmetric (Hessian) matrix $A$, an $n$-vector $b$, a scalar “radius” $\Delta>0$, and the Euclidean norm $\|\cdot\|$, the trust-region subproblem is to find
(1.1) $x_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|\leq\Delta.$}
Such problems arise in trust-region methods for (non-convex) nonlinear optimization [ConnGoulToin00, NoceWrig99]. A variety of methods have been proposed to solve them, some based on matrix factorization [GoulRobiThor10:mpc, MoreSore83:sisc], others on approximations from Krylov subspaces [CartGoulToin22, GoulLuciRomaToin99:siopt] and some on reformulation as eigenproblems [AdacIwatNakaTake17:siopt, RendWolk97:mp, RojaSantSore01:siopt, SantSore95]. It is well known [Gay81:sisc, MoreSore83:sisc] that the solution $x(\sigma)$ satisfies the first-order optimality condition
(1.2) $(A+\sigma I)x(\sigma)=b,$}
where $\|x(\sigma)\|\leq\Delta$ and the scalar shift
(1.3) $\sigma\geq\sigma_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A))$}
satisfies the complementarity condition $\sigma(\|x(\sigma)\|-\Delta)=0$; here $\{\lambda_{1}(A),\ldots,\lambda_{n}(A)\}$ are the eigenvalues of $A$, in increasing order, i.e., $\lambda_{i}(A)\leq\lambda_{i+1}(A)$, for $i=1,\ldots,n-1$. That is, either, for positive semi-definite $A$,
(1.4) $Ax(0)=b\;\;\mbox{with}\;\;\|x(0)\|\leq\Delta$}
or, with unrestricted $A$,
(1.5) $(A+\sigma I)x(\sigma)=b\;\;\mbox{with}\;\;\|x(\sigma)\|=\Delta\;\;\mbox{for some}\;\;\sigma\geq\sigma_{\min}.$}
We refer to the first possibility as the trivial—or interior—case, while the second is the general one; there is an additional probability-zero possibility, commonly called the hard case—this occurs if $b$ is orthogonal to all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)$ and the radius is too large—that we shall ignore for the time being. If we can rule out the trivial and hard cases, (1.5) then suggests an iteration in which, for a sequence of positive $\sigma$, $x(\sigma)$ is found by solving the linear system (1.2), and its suitability assessed by examining the size of the residual $\|x(\sigma)\|-\Delta$.
An interesting question is, therefore, what we can say about the manifold ${\cal S}(A,b)=\mbox{span}\{x(\sigma)\}$ as $\sigma$ varies between $\sigma_{\min}$ and $\infty$. There is no loss of generality, from a theoretical perspective, in considering a problem with a diagonal Hessian, since if $A$ has the spectral factorization
(1.6) $A=U\Lambda U^{T},$}
involving an orthogonal matrix $U$ whose columns are its eigenvectors, and a diagonal matrix $\Lambda$ whose entries are its eigenvalues, (1.1) is equivalent to
(1.7) ${\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}n$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}\Lambda y-d^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
under the transformation
$y=U^{T}x,\;\;\mbox{where}\;\;d=U^{T}b.$}
Consider, then, problem (1.7) with $n=10^{4}$, $d$ a vector of ones, eigenvalues $\lambda_{i}$, $i=1,\ldots,n$, in $(0,1]$ that are either (i) equi-spaced at $i/n$, (ii) clustered at the lower end at $i^{2}/n^{2}$, (iii) clustered at the higher end at $1-(i-1)^{2}/n^{2}$, or (iv) logarithmically-distributed in $(10^{-15},1]$ as $\exp(\log(10^{-15})+(i-1)\beta)$, where $\beta=(\log(1)-\log(10^{-15})/(n-1)$, and $p=10n$ evaluation points $\sigma=\sigma_{j}=j/n$ for $j=1,\ldots,p$ in $(0,10]$. In this case, the components of $y(\sigma)$ from the first-order optimality system
$(\Lambda+\sigma I)y(\sigma)=d$}
for (1.7) are $y_{i}(\sigma)=1/(\lambda_{i}+\sigma)$. If we approximate the solution manifold ${\cal S}(\Lambda,d)$ by the discrete set $\{y(\sigma_{j}),j=1,\ldots,p\}$ at the evaluation points, and compute the singular value decomposition of the resulting $n$ by $p$ matrix $S$ with entries
(1.8) $s_{i,j}=d_{i}/(\lambda_{i}+\sigma_{j})\;\;\mbox{for}\;\;i=1,\ldots,n\;\;\mbox{and}\;\;j=1,\ldots,p,$}
we find 37 significant singular values (those that are larger than $10^{-15}$ of the largest) in case (i), and for cases (ii)–(iv) there are 38, 36 and 35 respectively. The precise singular values are given in Table 1.1.
(i): equi-spaced (ii): clustered lower (iii): clustered higher (iv): logarithmically
2.53e+04 1.32e-03 1.34e+05 6.33e-03 1.86e+04 4.84e-04 1.09e+06 4.64e-03
1.44e+04 4.99e-04 4.38e+04 2.48e-03 1.09e+04 1.74e-04 7.32e+04 1.81e-03
6.89e+03 1.87e-04 1.92e+04 9.70e-04 5.12e+03 6.23e-05 2.09e+04 7.04e-04
3.07e+03 7.00e-05 8.81e+03 3.78e-04 2.21e+03 2.22e-05 7.75e+03 2.72e-04
1.33e+03 2.60e-05 3.95e+03 1.46e-04 9.19e+02 7.85e-06 3.23e+03 1.05e-04
5.64e+02 9.64e-06 1.73e+03 5.65e-05 3.76e+02 2.77e-06 1.39e+03 4.04e-05
2.36e+02 3.55e-06 7.42e+02 2.17e-05 1.51e+02 9.70e-07 5.96e+02 1.55e-05
9.78e+01 1.30e-06 3.15e+02 8.33e-06 6.03e+01 3.39e-07 2.52e+02 5.90e-06
4.01e+01 4.76e-07 1.33e+02 3.18e-06 2.38e+01 1.18e-07 1.05e+02 2.24e-06
1.63e+01 1.73e-07 5.54e+01 1.21e-06 9.27e+00 4.07e-08 4.35e+01 8.51e-07
6.55e+00 6.29e-08 2.29e+01 4.59e-07 3.58e+00 1.40e-08 1.79e+01 3.21e-07
2.61e+00 2.27e-08 9.44e+00 1.74e-07 1.37e+00 4.80e-09 7.30e+00 1.21e-07
1.04e+00 8.18e-09 3.86e+00 6.54e-08 5.20e-01 1.64e-09 2.96e+00 4.54e-08
4.07e-01 2.93e-09 1.57e+00 2.46e-08 1.96e-01 5.57e-10 1.19e+00 1.70e-08
1.59e-01 1.05e-09 6.34e-01 9.19e-09 7.33e-02 1.89e-10 4.79e-01 6.35e-09
6.18e-02 3.73e-10 2.55e-01 3.43e-09 2.72e-02 6.37e-11 1.91e-01 2.36e-09
2.38e-02 1.32e-10 1.02e-01 1.28e-09 1.00e-02 2.14e-11 7.60e-02
9.14e-03 4.69e-11 4.05e-02 4.74e-10 3.67e-03 3.01e-02
3.48e-03 1.61e-02 1.76e-10 1.34e-03 1.18e-02
Table 1.1: Singular values of dominant manifold subspace for four eigenvalue distributions
It is then remarkable that the manifold appears to be spanned by so few significant vectors, that is, it is approximately of very low rank. This turns out not to be a coincidence, and we shall exploit this throughout.
This paper is organised as follows. In Section 2, we explain why the solution manifold of interest is (approximately) low rank. We follow this in Section 3 by considering methods that build approximations to this manifold. A particularly appealing approach may be constructed from a so-called extended Krylov subspace, and we show how in Section 4. This then provides the basis for methods aimed at the trust-region subproblem (1.1), and we provide details for positive-definite $A$ in Section 5. We follow this with a description of how we may extend this for general $A$ in Section 6. Numerical experiments follow in Section 7. We show that the same ideas trivially extend to the regularized-norm problem in Section 8, and finally we conclude in Section 9.
2 Cauchy matrices and accurate low-rank approximation
The matrix (1.8) is an example of what is commonly known as a generalised Cauchy—or Cauchy-like—matrix [Cauc24:wiki]. To be more specific, suppose that we know the ordered, increasing eigenvalues $\lambda_{i}\in[\lambda_{1},\lambda_{n}]$ of $A$ (and thus $\Lambda$), and have picked a set of distinct shifts $\sigma_{i}\in[\sigma_{\min}^{+},\sigma_{\max}]$, $i=1,\ldots,p$, where
$\sigma_{\min}^{+}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}+\delta)\;\;\mbox{and}\;\;\sigma_{\max}>\sigma_{\min}^{+}$}
for some $\delta>0$. Then $\sigma_{\min}=\max(0,-\lambda_{1})\leq\sigma_{\min}^{+}$, and moreover $\sigma_{\min}^{+}\geq 0>-\delta\geq-\lambda_{1}$ whenever $\lambda_{1}\geq\delta$ and $\sigma_{\min}^{+}\geq-\lambda_{1}+\delta>-\lambda_{1}$ otherwise. Thus $\sigma_{\min}^{+}<\lambda_{1}$, indeed $\sigma_{\min}^{+}+\lambda_{1}\geq\delta$, and the intervals $\sigma_{\max},-\sigma_{\min}^{+}]$ and $\lambda_{1},\lambda_{n}]$ are non-overlapping.
To investigate the rank of $S$, clearly rows $i$ and $j$ of $S$ are linearly dependent if $\lambda_{i}=\lambda_{j}$, and thus $\mbox{rank}(S)=\mbox{rank}(\bar{S})$, where $\bar{S}$ comprises the rows of $S$ in which each set of such dependencies is replaced by a single representative.
For any $m$ by $n$ matrix $X$ with decreasing singular values $\sigma_{1}()\geq\sigma_{2}(X)\geq\ldots\geq\sigma_{k}(X)$ for increasing $k$, the $\epsilon$-rank, $\mbox{rank}_{\epsilon}(X)$, of $X$ is the smallest integer $k$ for which $\sigma_{k+1}(X)\leq\epsilon\sigma_{1}(X)$. Then Beckermann and Townsend [BeckTown17:simaa, §4.2] show that [^1] Stronger bounds are possible [BeckTown17:simaa] in terms of the so-called Grötzsch ring function, and other bounds are also known, e.g. [Gras04:nlaa].
$\mbox{rank}_{\epsilon}(\bar{S})\leq\left\lceil\log(16\gamma)\log(4/\epsilon)/\pi^{2}\right\rceil,$}
involving the cross-ratio
$\gamma=\left|\frac{(\sigma_{\max}+\lambda_{1})(\sigma_{\min}^{+}+\lambda_{n})}{(\sigma_{\max}+\lambda_{n})(\sigma_{\min}^{+}+\lambda_{1})}\right|,$}
because we have removed repeated eigenvalues, the shifts are distinct and the two sets are non-overlapping. Notice that the cross ratio is bounded from above as $\sigma_{\min}^{+}+\lambda_{1}\geq\delta>0$, and approaches $(\sigma_{\min}^{+}+\lambda_{n})/(\sigma_{\min}^{+}+\lambda_{1})$ as $\sigma_{\max}\rightarrow\infty$; for positive definite $A$, the latter is simply the spectral condition number. In particular, the log dependence on both the cross ratio and the precision $\epsilon$ are significant in keeping the bound small, and its dimension independence is also noteworthy. For the examples illustrated in Table 1.1, this bound is 44, a very reasonable (upper-bound) approximation of the computed values.
3 Manifold approximation from Krylov subspaces
Now that we know that the solution we seek to (1.1) for any $\Delta$ lies in an approximately low-dimensional subspace, our task is to identify a suitable (for convenience, orthogonal) $n$ by $\ell$ basis matrix $V_{\ell}$ for this subspace. Once we do so, we may recover an approximate solution $x_{\ell}=V_{\ell}y_{\ell}$, where
(3.1) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}P_{\ell}y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
with $P_{\ell}=V_{\ell}^{T}AV_{\ell}$ and $b_{\ell}=V_{\ell}^{T}b$. The fact that we expect $\ell$ to be small provides an opportunity to use methods that would be inappropriate has it been larger.
An obvious approach might be to simply select a set of $\sigma_{i}\geq\sigma_{\min}^{+}$ for $i=1,\ldots,\ell$, compute the resulting solutions $x(\sigma_{i})$ and build an orthogonal basis of the set $\{x(\sigma_{i})\}$ by, for example, the modified Gram-Schmidt process. This has two fundamental issues. Firstly, it requires that we solve $\ell$ shifted linear systems, and this is not without cost. Secondly, it is not known a priori how big $\ell$ should be. The latter may be addressed as follows. Suppose that $y_{\ell}$ is the solution to (3.1) with $\|y_{\ell}\|=\Delta$. If so, there is a shift $\sigma_{\ell}\geq 0$, just as in (1.2), for which
(3.2) $0=(P_{\ell}+\sigma_{\ell}I)y_{\ell}-b_{\ell}=V_{\ell}^{T}\left[(A+\sigma_{\ell}I)x_{\ell}-b\right]$}
where $x_{\ell}=V_{\ell}y_{\ell}$ and $\|x_{\ell}\|=\Delta$, since $V_{\ell}$ has orthogonal columns (i.e., $V_{\ell}^{T}V_{\ell}=I$). As the solution we seek satisfies $(A+\sigma I)x=b$, and as $V_{\ell}$ will be an increasingly better approximation to the low-dimensional subspace that contains the required solution, it suffices to monitor the residual $r_{\ell}=(A+\sigma_{\ell}I)x_{\ell}-b$ and to stop when $\|r_{\ell}\|$ is sufficiently small. This, however, does not avoid the cost of multiple factorizations; rival methods, e.g., [GoulRobiThor10:mpc, MoreSore83:sisc], also use factorizations, but experience suggests that relatively few often suffice. Another approach for constructing a good approximation of the manifold is to use a method based on a rational Krylov subspace [Ruhe84:LAA, AldaPali25:ARXIV]. However, this will still require solving $\ell$ shifted linear systems.
The alternative is to try to build a good approximation to the manifold by other means. We shall resort to methods based on the extended Krylov subspace (see Section 4).
To justify this, first note that given a shift $\sigma$, (1.2) and (1.6) give that
$x(\sigma)=(A+\sigma I)^{-1}b=Uy(\sigma),$}
where
$y_{i}(\sigma)=\frac{d_{i}}{\lambda_{i}+\sigma},\;\;i=1,\ldots,n,\;\;\mbox{and}\;\;d=U^{T}b.$}
Consider two cases and first suppose that Then using a Laurent series [Laur24:wiki], we have The smaller the ratio the faster the sequence reaches machine precision and few terms would be sufficient to approximate $y_{i}(\sigma)$. In other words, $y(\sigma)$ can be very well approximated using a polynomial with a moderate degree in , and hence $x(\sigma)$ can be well approximated using a polynomial with a moderate degree in applied to $b$. , when $i=1,\ldots,n$, a similar argument using shows that $x(\sigma)$ can be well approximated using a polynomial with moderate degree in applied to $b$.
An extended Krylov subspace is by definition the space that constructs the sum space of both Krylov subspaces ${\cal K}(A,b)=\text{span}\{A^{i}b\}_{i\geq 0}$ and ${\cal K}(A^{-1},b)=\text{span}\{A^{-i}b\}_{i\geq 0}$, and constructs the aforementioned polynomials required to handle the two ends of the behaviour spectrum. Hence, these observations provide some intuition as to why we choose an extended Krylov method (see Section 4) to solve the trust-region subproblem. Similar arguments have been made by Druskin and Knizhnerman [DrusKniz98:simaa] for different applications.
For a better understanding of the convergence of the extended-Krylov-subspace method when solving shifted linear systems with a symmetric positive definite matrix whose spectrum is contained in $\lambda_{1},\lambda_{n}]$ and shifts lying outside $(-\lambda_{n}-\delta_{n},-\lambda_{1}+\delta_{1})$ for $\delta_{1},\delta_{n}>0$, we provide a convergence analysis in Appendix A that is heavily based on that by Knizhnerman and Simoncini [KnizSimo11:NUMMAT]. This shows that there will be an element $x_{i}(\sigma)$ of the extended Krylov subspace ${\cal K}_{i}(A,b)$, defined in (4.1), for which
$\|x_{i}(\sigma)-x(\sigma)\|\leq c\mu^{i}$
for some constants $c$ and $\mu<1$ and all $i\geq 0$.
4 The EKS algorithm
Our aim is thus to construct efficiently an orthogonal basis $V_{2m+1}$ of the evolving extended Krylov subspaces
(4.1) ${\cal K}_{2m}(A,b)=\text{span}\{b,A^{-1}b,Ab,A^{-2}b,A^{2}b,\ldots,A^{-m}b\}$}
and
${\cal K}_{2m+1}(A,b)=.$}
Fortunately, much of the groundwork is already in place. We suppose, for the time being, that $A$ is positive definite. Then Jagels and Reichel [^2] They also propose a second, more expensive algorithm that copes with indefinite $A$. [JageReic09:laa] have shown how to build such a basis if the order [^3] We shall say why our order is more convenient later. in which the components $A^{-m}b$ and $A^{m}b$ arise is interchanged, that is if instead
${\cal K}^{\prime}_{2m}(A,b)=\text{span}\{b,Ab,A^{-1}b,A^{2}b,A^{-2}b,\ldots,A^{m}b\}$}
and
${\cal K}_{2m+1}(A,b)=.$}
Using the same logic as in [JageReic09:laa, Alg.2.1], that applies to ${\cal K}^{\prime}_{2m}(A,b)$, it is straightforward to derive the following algorithm, Algorithm 1 1, that instead uses the order ${\cal K}_{2m}(A,b)$.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$ and subspace dimension bound $m\geq 1$.
Output: orthonormal $V_{2m+1}=$
     $(v_{0}:v_{-1}:v_{1}:\cdots:v_{k-1}:v_{-k}:v_{k}:v_{-k-1}:v_{k+1}:\cdots:v_{-m}:v_{m})\in\hbox{I\thinspace R}^{n\times(2m+1)}.$
Algorithm:
1
22mm $\delta_{0}:=\|b\|$;
3 if $\delta_{0}=0$ stop, subspace completed;
4 $v_{0}:=b/\delta_{0}$, $u:=A^{-1}v_{0}$;
5 $\beta_{0}:=u^{T}v_{0}$, $u\leftarrow u-\beta_{0}v_{0}$;
6 $\delta_{-1}:=\|u\|$;
7 for $k=1,2,\ldots,m$ do
8 if $\delta_{-k}=0$ stop, subspace completed;
9 $v_{-k}:=u/\delta_{-k}$, $u:=Av_{-k}$;
10 $\alpha_{k-1}:=u^{T}v_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
11 $\alpha_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\alpha_{-k}v_{-k}$;
12 $\delta_{k}:=\|u\|$;
13 if $\delta_{k}=0$ stop, subspace completed;
14 $v_{k}:=u/\delta_{k}$, $u:=A^{-1}v_{k}$;
15 $\beta_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$;
16 $\beta_{k}:=u^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
17 if $k<m$ then
18    $\delta_{-k-1}:=\|u\|$;
19
20
Algorithm 1 The EKS algorithm to find an orthogonal basis of ${\cal K}_{2m+1}(A,b)$
Thus
$\delta_{0}v_{0}$ $=b$
(4.2) $\delta_{-1}v_{-1}$ $=A^{-1}v_{0}-\beta_{0}v_{0}$
(4.3) $\delta_{k}v_{k}$ $=Av_{-k}-\alpha_{-k}v_{-k}-\alpha_{k-1}v_{k-1}\;\;\mbox{for $k=1,\ldots,m$}\;\;$
(4.4) $\delta_{-k-1}v_{-k-1}$ $=A^{-1}v_{k}-\beta_{-k}v_{-k}-\beta_{k}v_{k}\;\;\mbox{for $k=1,\ldots,m-1$}\;\;$
Each iteration of the main loop requires a solve and a product involving $A$, the initialization requires a single solve. The remaining operations simply involve scalar products and vector sums. The decision whether and when to stop in the main loop is the focus of Sections 5 and 6, but observe that the extended Krylov subspace has definitely been completed if ever $\delta_{i}=0$.
We consider the symmetric $\hbox{I\thinspace R}^{(2m+1)\times(2m+1)}$ “projected” matrix
$P_{2m+1}=V_{2m+1}^{T}AV_{2m+1},$
whose $i,j$-entry we denote as $p_{i,j}$, as well as its leading $2k$ and $2k+1$ sub-blocks
$P_{2k}=V_{2k}^{T}AV_{2k}\;\;\mbox{and}\;\;P_{2k+1}=V_{2k+1}^{T}AV_{2k+1},$
where
$V_{2k}=(v_{0}:v_{-1}:v_{1}:\cdots:v_{k-1}:v_{-k})\;\;\mbox{and}\;\;V_{2k+1}=(V_{2k}:v_{k}),$
that play the roles of $A_{k}$ in (3.1). It follows from (4.3) that
(4.5) $Av_{-k}=\alpha_{k-1}v_{k-1}+\alpha_{-k}v_{-k}+\delta_{k}v_{k}\;\;\mbox{for $k=1,\ldots,m$}\;\;$
and the orthogonality of the $v_{i}$ then gives
(4.6) $p_{2k-1,2k}=v_{k-1}^{T}Av_{-k}=\alpha_{k-1},\;\;p_{2k,2k}=v_{-k}^{T}Av_{-k}=\alpha_{-k},\;\;p_{2k+1,2k}=v_{k}^{T}Av_{-k}=\delta_{k}$
and
$p_{i,2k}=v_{i}^{T}Av_{-k}=0\;\;\mbox{for $i\neq 2k-1,-2k$ and $2k+1$.}\;\;$
By symmetry of $P_{2m+1}$, we also have
$p_{2k,2k-1}=\alpha_{k-1},\;\;p_{2k,2k}=\alpha_{-k},\;\;p_{2k,2k+1}=\delta_{k}\;\;\mbox{and}\;\;p_{2k,i}=0\;\;\mbox{for the remaining $i$.}\;\;$
Thus, from (4.5) and (4.6), we have
(4.7) $Av_{-k}=p_{2k,2k-1}v_{k-1}+p_{2k,2k}v_{-k}+p_{2k,2k+1}v_{k}\;\;\mbox{for $k=1,\ldots,m$,}\;\;$}
where formally any $p_{2k,i}=0$ if $i\leq 0$. Similarly, (4.4) gives
$A^{-1}v_{k}=\beta_{-k}v_{-k}+\beta_{k}v_{k}+\delta_{-k-1}v_{-k-1}\;\;\mbox{for $k=1,\ldots,m-1$}\;\;$
and thus that
(4.8) $\begin{array}[]{rl}\beta_{k}Av_{k}\hskip-5.69054pt&=v_{k}-\beta_{-k}Av_{-k}-\delta_{-k-1}Av_{-k-1}\\ &=-\beta_{-k}\alpha_{k-1}v_{k-1}-\beta_{-k}\alpha_{-k}v_{-k}+(1-\beta_{-k}\delta_{k}-\delta_{-k-1}\alpha_{k})v_{k}\\ &\hskip 5.69054pt\hskip 8.53581pt-\delta_{-k-1}\alpha_{-k-1}v_{-k-1}-\delta_{-k-1}\delta_{k+1}v_{k+1}\end{array}$}
using (4.5). Again, the orthogonality of the $v_{i}$ and symmetry implies that [^4] When $\delta_{-k-1}=0$, the last three terms in (4.9) are $p_{2k+1,2k+1}=(1-\beta_{-k}\delta_{k})/\beta_{k}$ and $p_{2k+2,2k+1}=p_{2k+3,2k+1}=0$, and thus do not require that we calculate $\alpha_{k}$, $\alpha_{-k-1}$ and $\delta_{k+1}$.
(4.9) $\begin{array}[]{rl}p_{2k-1,2k+1}\hskip-5.69054pt&=p_{2k+1,2k-1}=v_{k-1}^{T}Av_{k}=-\displaystyle\frac{\beta_{-k}\alpha_{k-1}}{\beta_{k}},\\ p_{2k,2k+1}\hskip-5.69054pt&=p_{2k+1,2k}=v_{-k}^{T}Av_{k}=-\displaystyle\frac{\beta_{-k}\alpha_{-k}}{\beta_{k}},\\ p_{2k+1,2k+1}\hskip-5.69054pt&=v_{k}^{T}Av_{k}=\displaystyle\frac{1-\beta_{-k}\delta_{k}-\delta_{-k-1}\alpha_{k}}{\beta_{k}},\\ p_{2k+2,2k+1}\hskip-5.69054pt&=p_{2k+1,2k+2}=v_{-k-1}^{T}Av_{k}=-\displaystyle\frac{\delta_{-k-1}\alpha_{-k-1}}{\beta_{k}},\\ p_{2k+3,2k+1}\hskip-5.69054pt&=p_{2k+1,2k+3}=v_{k+1}^{T}Av_{k}=-\displaystyle\frac{\delta_{-k-1}\delta_{k+1}}{\beta_{k}}\end{array}$}
and
$p_{i,2k+1}=p_{2k+1,i}=v_{i}^{T}Av_{k}=0\;\;\mbox{for $i\neq-k-1,k-1,k,-k$ and $k+1$.}\;\;$
Thus, it follows from from (4.8) and (4.9) that
(4.10) $\begin{array}[]{rl}Av_{k}\hskip-5.69054pt&=p_{2k-1,2k+1}v_{k-1}+p_{2k,2k+1}v_{-k}+p_{2k+1,2k+1}v_{k}\;+\\ &\hskip 5.69054pt\hskip 5.69054ptp_{2k+2,2k+1}v_{-k-1}+p_{2k+3,2k+1}v_{k+1}\;\;\mbox{for $k=1,\ldots,m-1,$}\end{array}$}
where now formally any $p_{2k+1,i}=0$ if $i\leq 0$. Moreover, note that
$p_{2k,2k+1}=v_{-k}^{T}Av_{k}=\delta_{k}\;\;\mbox{and}\;\;p_{2k+2,2k+1}=v_{-k-1}^{T}Av_{k}=\alpha_{k}$
because of (4.6), and
$\frac{\beta_{-k}\alpha_{k-1}}{\beta_{k}}=-\frac{\delta_{-k}\delta_{k}}{\beta_{k-1}}$
by comparing $p_{2k-1,2k+1}$ with $p_{2k-3,2k-1}$ for subsequent $k$. Finally, if follows from (4.2), (4.5) with $k=1$ and the orthogonality of the $v_{i}$, that [^5] When $\delta_{-1}=0$, $p_{1,1}=1/\beta_{0}$ and $p_{1,2}=p_{1,3}=0$, and thus $\alpha_{0}$, $\alpha_{-1}$ and $\delta_{1}$ need not be calculated.
(4.11) $p_{1,1}=v_{0}^{T}Av_{0}=\frac{1-\delta_{-1}\alpha_{0}}{\beta_{0}},\;\;p_{2,1}=v_{-1}^{T}Av_{0}=-\frac{\delta_{-1}\alpha_{-1}}{\beta_{0}}\equiv\alpha_{0},\;\;p_{3,1}=v_{1}^{T}Av_{0}=-\frac{\delta_{-1}\delta_{1}}{\beta_{0}},$}
where the alternative definition of $p_{1,2}$ follows by symmetry and from (4.6) with $k=0$.
Thus $P_{2k+3}$ is a symmetric, pentadiagonal matrix of the form
$P_{2k+3}=\left[\begin{array}[]{ccccccccc}p_{1,1}&\alpha_{0}&p_{3,1}&0&0&0&0&\cdots\\ \alpha_{0}&\alpha_{-1}&\delta_{1}&0&0&0&0&\cdots\\ p_{3,1}&\delta_{1}&p_{3,3}&\alpha_{1}&p_{5,3}&0&0&\cdots\\ 0&0&\alpha_{1}&\alpha_{-2}&\delta_{2}&0&0&\cdots\\ 0&0&p_{5,3}&\delta_{2}&p_{5,5}&\alpha_{2}&p_{7,5}&\cdots\\ 0&0&0&0&\alpha_{2}&\alpha_{-3}&\delta_{3}&\cdots\\ 0&0&0&0&p_{7,5}&\delta_{3}&p_{7,7}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right]$
[…]
where the entries $p_{2k+1,2k+1}$ and $p_{2k+1,2k-1}$ (etc) are as given above; the leading sub-matrices $P_{2k}$ and $P_{2k+1}$ are the entries in the two “hashed” boxes. In other words, we may grow the (lower-triangular parts of the) submatrices $P_{i}$ of $P_{2m+1}$, $i=1,\ldots,2m+1$, simply by introducing, for each successive column, the entries
(4.13) $\begin{array}[]{rl}p_{2k+1,2k+1}=&\displaystyle\frac{1-\beta_{-k}\delta_{k}-\delta_{-k-1}\alpha_{k}}{\beta_{k}},\;\;p_{2k+2,2k+1}=\alpha_{k}\;\;\mbox{and}\\ p_{2k+3,2k+1}=&-\displaystyle\frac{\delta_{-k-1}\delta_{k+1}}{\beta_{k}}\;\;\mbox{for odd $i=2k+1$}\end{array}$}
and
(4.14) $p_{2k,2k}=\alpha_{-k},\;\;p_{2k+1,2k}=\delta_{k}\;\;\mbox{and}\;\;p_{2k+2,2k}=0\;\;\mbox{for even $i=2k$,}\;\;$}
starting with the first column (4.11) when $i=1$.
Trivially, it also follows from the orthogonality of $V_{k}$ that
(4.15) $V^{T}_{2m+1}b=\|b\|e_{1},$}
where $e_{1}$ is the first unit vector, and thus $b_{k}=\|b\|e_{1}$ in (3.1) for this choice of $V_{k}$.
We may also combine (4.7) and (4.10) to deduce that
(4.16) $AV_{2k}=V_{2k}P_{2k}+p_{2k+1,2k-1}v_{k}e_{2k-1}^{T}+p_{2k+1,2k}v_{k}e_{2k}^{T}$}
and
(4.17) $AV_{2k+1}=V_{2k+1}P_{2k+1}+(p_{2k+2,2k+1}v_{-k-1}+p_{2k+3,2k+1}v_{k+1})e_{2k+1}^{T}$}
These identities will shortly be used to develop suitable stopping rules for our algorithm.
It is perhaps worth mentioning that lines 4–5 and the first part of 8, 9–11 and the first part of 13 and 14–15, 17 and the first part of 8 in Algorithm 1 are mathematically equivalent to
$\begin{array}[]{rl}\beta_{0}&\hskip-5.69054pt:=u^{T}v_{0},\\ \delta_{-1}&:=\sqrt{\|u\|^{2}-\beta_{0}^{2}},\\ v_{-1}&:=[u-\beta_{0}v_{0}]/\delta_{-1}\\ \\ \alpha_{k-1}&\hskip-5.69054pt:=u^{T}v_{k-1},\alpha_{-k}:=u^{T}v_{-k},\\ \delta_{k}&:=\sqrt{\|u\|^{2}-\alpha_{k-1}^{2}-\alpha_{-k}^{2}},\\ v_{k}&:=[u-\alpha_{k-1}v_{k-1}-\alpha_{-k}v_{-k}]/\delta_{k}\end{array}$
and
$\begin{array}[]{rl}\beta_{-k}&\hskip-5.69054pt:=u^{T}v_{-k},\beta_{k}:=u^{T}v_{k}^{T},\\ \delta_{-k-1}&:=\sqrt{\|u\|^{2}-\beta_{-k}^{2}-\beta_{k}^{2}},\\ v_{-k-1}&:=[u-\beta_{-k}v_{-k}-\beta_{k}v_{k}]/\delta_{-k-1}.\end{array}$
However, numerical experience suggests that this form is unstable, in a manner akin to that of Gram-Schmidt orthogonalisation of a set of vectors and how such instability may be cured by its modified form [GoluvanL96, §5.2.7–5.2.8].
5 Solving the convex trust-region subproblem
Having determined an orthogonal basis $V_{\ell}$ of ${\cal K}_{\ell}(A,b)$ using the EKS algorithm, we now turn to the trust-region subproblem (3.1) within this subspace. As we have just seen, the Hessian $P_{\ell}=V_{\ell}^{T}AV_{\ell}$ is symmetric and sparse (pentadiagonal), while the first-order term $b_{\ell}$ is a scalar multiple $\|b\|$ of the $\ell$-dimensional first unit vector. Significantly, we are expecting $\ell$ to be small, and thus the subproblem is low dimensional.
Thus we might apply off-the-shelf trust-region software such as GQT [MoreSore83:sisc] or TRS [GoulRobiThor10:mpc] from GALAHAD [GoulOrbaToin03:toms]. However, the alternative we prefer is to proceed as in (1.6), that is to compute the spectral factorization
(5.1) $P_{\ell}=U_{\ell}\Lambda_{\ell}U_{\ell}^{T}$}
involving matrices of eigenvectors $U_{\ell}$ and (diagonal) eigenvalues $\Lambda_{\ell}$, and to find $y_{\ell}=U_{\ell}z_{\ell}$, where
(5.2) $z_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle z\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{\ell}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}z^{T}\Lambda_{\ell}z-d_{\ell}^{T}z\;\;\mbox{subject to}\;\;\|z\|\leq\Delta$}
and $d_{\ell}=U_{\ell}^{T}b_{\ell}$. Significantly (5.1) is inexpensive since $\ell$ is small, and the Hessian of (5.2) is diagonal. This then enables very fast and accurate solution using high-order, factorization-free iterative schemes, such as TRS_solve_diagonal [GoulRobiThor10:mpc] from GALAHAD.
Having computed the solution $y_{\ell}$ to (3.1) as above (or by any other means), together with the corresponding shift $\sigma_{\ell}$ for which both
(5.3) $(P_{\ell}+\sigma_{\ell}I)y_{\ell}=\|b\|e_{1}$}
and $\|x_{\ell}\|=\Delta$ hold, according to (3.2), it remains to see how big is the residual
$r_{\ell}\stackrel{{\scriptstyle\rm def}}{{=}}(A+\sigma_{\ell}I)x_{\ell}-b,$}
where $x_{\ell}=V_{\ell}y_{\ell}$. When $\ell=2k+1$, we find that
[…]
using (4.17), (5.3), the orthogonality of the columns of $V_{2k+1}$, and the relationship $V_{2k+1}\|b\|e_{1}=b$. Thus
$r_{2k+1}=(p_{2k+2,2k+1}v_{-k-1}+p_{2k+3,2k+1}v_{k+1})e_{2k+1}^{T}y_{2k+1}$
and hence
(5.4) $\|r_{2k+1}\|=|y_{2k+1,2k+1}|\sqrt{p^{2}_{2k+2,2k+1}+p^{2}_{2k+3,2k+1}},$}
where $y_{2k+1,2k}$ and $y_{2k+1,2k+1}$ are the $2k$-th and $2k+1$-st components of $y_{2k+1}$, using the orthogonality of $v_{-k-1}$ and $v_{k+1}$. Thus the norm of the residual may be computed at almost no extra cost using data that is already available. Identical reasoning, when $\ell=2k$ but now using (4.16), shows that
$r_{2k}=(p_{2k+1,2k-1}e_{2k-1}^{T}y_{2k}+p_{2k+1,2k}e_{2k}^{T}y_{2k})v_{k}$
and hence in this case
(5.5) $\|r_{2k}\|=|p_{2k+1,2k-1}y_{2k,2k-1}+p_{2k+1,2k}y_{2k,2k}|.$}
Note that in the exceptional case where $\delta_{i}=0$, the corresponding $\|r_{i}\|=0$, and thus termination will always occur at this step.
The TREK algorithm embeds solutions of (3.1) after every computation of $v_{-k}$ and $v_{k}$ in ; at these stages the matrices $V_{2k}$ and $V_{2k+1}$, respectively, are complete, and we have the data required to solve the appropriate subproblems (3.1). Armed with such solutions, we terminate as soon as $\|r_{2k}\|$ from (5.5) or $\|r_{2k+1}\|$ from (5.4) is deemed small enough. One further important point is that after computing $u:=A^{-1}v_{0}=A^{-1}b/\|b\|$ before we start the main loop in the EKS algorithm, it is trivial to check for the interior case (1.4). Thus if we check and find $\|u\|\|b\|\leq\Delta$ in TREK, the solution is interior, and we exit [^6] Actually, we compute $x=A^{-1}b$, exit with $x_{*}=x$ if $\|x\|\leq\Delta$, and otherwise continue with $u=x/\|b\|$. with $x_{*}=\|b\|u$. For completeness we state the full TREK algorithm (Algorithm 2) 2.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|\leq\Delta.$
2mm Algorithm:
$x:=A^{-1}b$;
if $\|x\|\leq\Delta$ then
 exit with the interior solution $x_{*}=x$, shift $\sigma_{*}=0$ and $k_{\mbox{\scriptsize stop}}=0$;
$\delta_{0}:=\|b\|$, $v_{0}:=b/\delta_{0}$, $u:=x/\delta_{0}$;
$\beta_{0}:=u^{T}v_{0}$, $u\leftarrow u-\beta_{0}v_{0}$;
$\delta_{-1}:=\|u\|$;
for $k=1,2,\ldots,m$ do
 if $\delta_{-k}>0$ then
    $v_{-k}:=u/\delta_{-k}$, $u:=Av_{-k}$;
    $\alpha_{k-1}:=u^{T}v_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
    $\alpha_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\alpha_{-k}v_{-k}$;
    $\delta_{k}:=\|u\|$;
 if $k>1$ then
     update $P_{2k-1}$ from $P_{2k-2}$ using (4.13) and form $b_{2k-1}=(b^{T}_{2k-2},0)^{T}$;
 else
     initialize $P_{1}$ from (4.11) and set $b_{1}=\delta_{0}$;
 $y_{2k-1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k-1}y-b_{2k-1}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ and its optimal shift $\sigma_{2k-1}$;
   compute $\|r_{2k-1}\|$ from (5.4);
 if $\|r_{2k-1}\|\leq\epsilon$ then
    exit with the interior solution $x_{*}=V_{2k-1}y_{2k-1}$, shift $\sigma_{*}=\sigma_{2k-1}$ and $k_{\mbox{\scriptsize stop}}=k$;
 if $\delta_{k}>0$ then
    $v_{k}:=u/\delta_{k}$, $u:=A^{-1}v_{k}$;
    $\beta_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$;
    $\beta_{k}:=u^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
 if $k<m$ then
    if $\delta_{k}>0$ then
       $\delta_{-k-1}:=\|u\|$;
      update $P_{2k}$ from $P_{2k-1}$ using (4.14) and form $b_{2k}=(b^{T}_{2k-1},0)^{T}$;
    $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ together with its optimal shift $\sigma_{2k}$;
      compute $\|r_{2k}\|$ from (5.5);
    if $\|r_{2k}\|\leq\epsilon$ then
       exit with the interior solution $x_{*}=V_{2k}y_{2k}$, shift $\sigma_{*}=\sigma_{2k}$ and $k_{\mbox{\scriptsize stop}}=k$;
Algorithm 2 The TREK algorithm to solve the trust-region subproblem (1.1)
5.1 Resolves
The need to resolve a problem with the same $A$ and $b$, but a smaller $\Delta$ is a common occurrence in trust-region methods, since such a mechanism is used to guarantee convergence when the current solution has proved unsatisfactory [ConnGoulToin00]. This is easy to accommodate within Algorithm 2; the data generated by the previous call to the algorithm is retained, and the ‘”for loop” re-entered with $k$ taking the value $k_{\mbox{\scriptsize stop}}$ that it last had. The motivation is simply that although the segment of the manifold ${\cal S}(A,b)$ we have observed may now have to expand slightly, it may well be that there is little actual difference; the data provided in Appendix C indicates that in only two of ninety six cases examined were any further extended Krylov iterations necessary, and thus that resolves are generally very inexpensive. Examining the bound (A.6) given in Proposition A.3 in Appendix A, and looking at Figure A.1, we see that the convergence rate of the extended Krylov subspace improves for $\sigma$ beyond a certain point, which is consistent with this observation. Although other methods can also take advantage of resolves, all that we are aware of still require significant extra computation.
5.2 Other elliptical norms
Although the trust-region is often defined in terms of the Euclidean norm, it can be that this promotes some variables at the expense of others. To cope with this, it is sometimes more effective to pose the problem as
(5.6) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta,$}
where $P$ is an easily-invertible matrix—a diagonal $P$ is often a simple but effective choice. On writing
(5.7) $x_{\mbox{\tiny P}}=Px,$}
the minimizer of (5.6) is formally $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$, where
(5.8) $x_{\mbox{\tiny P}}{}_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x_{\mbox{\tiny P}}{}^{T}A_{\mbox{\tiny P}}x_{\mbox{\tiny P}}-b_{\mbox{\tiny P}}{}^{T}x_{\mbox{\tiny P}}\;\;\mbox{subject to}\;\;\|x_{\mbox{\tiny P}}\|\leq\Delta$}
and $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and $b_{\mbox{\tiny P}}=P^{-T}b$. Since the TREK algorithm described above merely relies on products with $A_{\mbox{\tiny P}}$ and its inverse $A_{\mbox{\tiny P}}{}^{-1}=PA^{-1}P^{T}$, once again we merely need a factorization of $A$ (to find $A^{-1}v$ for given $v$) so long as means of inverting $P$ and its transpose are available. Commonly a positive-definite scaling matrix $S$ that is (in some sense) close to $A$ is known, and in this case a decomposition, for example a sparse Cholesky factorization, $S=P^{T}P$ gives the appropriate $P$. We provide details of the necessary changes to devise suitable algorithms to solve the resulting problem
(5.9) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|_{S}\leq\Delta,$}
where $\|x\|_{S}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{x^{T}Sx}$, in Appendix B. We illustrate one such algorithm here as Algorithm 3 3. Note that here the generated vectors $v_{i}$ now form an $S$-orthonormal basis for the extended Krylov subspace.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|_{S}\leq\Delta.$
2mm Algorithm:
$x:=A^{-1}b$, $u:=Sx$;
if $\sqrt{u^{T}x}\leq\Delta$ then
 exit with the interior solution $x_{*}=x$, shift $\sigma_{*}=0$ and $k_{\mbox{\scriptsize stop}}=0$;
$w:=S^{-1}b$, $\delta_{0}:=\sqrt{b^{T}w}$;
$v_{0}:=w/\delta_{0}$, $q_{0}:=b/\delta_{0}$, $w:=u/\delta_{0}$, $u:=x/\delta_{0}$;
$\beta_{0}:=u^{T}q_{0}$, $u\leftarrow u-\beta_{0}v_{0}$;
$w=Su$, $\delta_{-1}:=\sqrt{u^{T}w}$;
for $k=1,2,\ldots,m$ do
 if $\delta_{-k}>0$ then
    $v_{-k}:=u/\delta_{-k}$, $q_{-k}:=w/\delta_{-k}$;
    $w:=Av_{-k}$, $u:=S^{-1}w$;
    $\alpha_{-k}:=w^{T}v_{-k}$, $u\leftarrow u-\alpha_{-k}v_{-k}$, $w\leftarrow w-\alpha_{-k}q_{-k}$;
    $\alpha_{k-1}:=w^{T}v_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$, $w\leftarrow w-\alpha_{k-1}q_{k-1}$;
    $\delta_{k}:=\sqrt{u^{T}w}$;
 if $k>1$ then
     update $P_{2k-1}$ from $P_{2k-2}$ using (4.13) and form $b_{2k-1}=(b^{T}_{2k-2},0)^{T}$;
 else
     initialize $P_{1}$ from (4.11) and set $b_{1}=\delta_{0}$;
 $y_{2k-1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k-1}y-b_{2k-1}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ and its optimal shift $\sigma_{2k-1}$;
   compute $\|r_{2k-1}\|$ from (5.4);
 if $\|r_{2k-1}\|\leq\epsilon$ then
    exit with the interior solution $x_{*}=V_{2k-1}y_{2k-1}$, shift $\sigma_{*}=\sigma_{2k-1}$ and $k_{\mbox{\scriptsize stop}}=k$;
 if $\delta_{k}>0$ then
    $v_{k}:=u/\delta_{k}$, $q_{k}:=w/\delta_{k}$;
    $u:=A^{-1}q_{k}$, $w:=Su$;
    $\beta_{-k}:=w^{T}v_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$, $w\leftarrow w-\beta_{-k}q_{-k}$;
    $\beta_{k}:=w^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$, $w\leftarrow w-\beta_{k}q_{k}$;
 if $k<m$ then
    if $\delta_{k}>0$ then
       $\delta_{-k-1}:=\sqrt{u^{T}w}$;
      update $P_{2k}$ from $P_{2k-1}$ using (4.14) and form $b_{2k}=(b^{T}_{2k-1},0)^{T}$;
    $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ together with its optimal shift $\sigma_{2k}$;
      compute $\|r_{2k}\|$ from (5.5);
    if $\|r_{2k}\|\leq\epsilon$ then
       exit with the interior solution $x_{*}=V_{2k}y_{2k}$, shift $\sigma_{*}=\sigma_{2k}$ and $k_{\mbox{\scriptsize stop}}=k$;
Algorithm 3 The TREK algorithm to solve the trust-region subproblem (5.9)
It is also worth noting that the extended Krylov subspace (4.1) generated by $A_{\mbox{\tiny P}}$ and $b_{\mbox{\tiny P}}$,
(5.10) ${\cal K}_{2m+1}(A_{\mbox{\tiny P}},b_{\mbox{\tiny P}})=\text{span}\{b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{-1}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}b_{\mbox{\tiny P}},\ldots,A_{\mbox{\tiny P}}^{-m}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{m}b_{\mbox{\tiny P}}\},$}
is
$P^{-T}\text{span}\{b,(SA^{-1})b,(AS^{-1})b,\ldots,(SA^{-m})b,(AS^{-1})^{m}b\},$}
and it is then possible to derive a suitable variant of Algorithm 1 using $S$ and its inverse rather than $P$ and $P^{-1}$.
6 Solving the general trust-region subproblem
For non-convex problems—indeed if $A$ is not positive definite—it may be awkward (or even impossible if $A$ is singular) to use an extended Krylov subspace involving products with $A^{-1}$. Fortunately, it is straightforward to deal with this possibility using the subspace generated by a diagonally-shifted matrix $A_{\sigma_{\mbox{\tiny S}}}=A+\sigma_{\mbox{\tiny S}}I$, for some suitable $\sigma_{\mbox{\tiny S}}>0$, and its inverse.
To see why, first note that the interior-solution case cannot occur if $A$ is indefinite, and thus the solution we seek must lie on the trust-region boundary. That is, we must satisfy (1.5). We may rewrite this as
(6.1) $(A_{\sigma_{\mbox{\tiny S}}}+\theta I)x_{\sigma_{\mbox{\tiny S}}}(\theta)=b\;\;\mbox{with}\;\;\|x_{\sigma_{\mbox{\tiny S}}}(\theta)\|=\Delta,$}
where $x(\sigma)\equiv x_{\sigma_{\mbox{\tiny S}}}(\theta)$ and
$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A)-\sigma_{\mbox{\tiny S}}).$}
Thus we seek the unique root of (6.1) but now over the interval $\theta\in[\sigma_{\min}-\sigma_{\mbox{\tiny S}},\infty)$. Given $\sigma_{\mbox{\tiny S}}$, it is trivial to adapt any bracketing root-finding procedure to account for this extended interval.
We have in mind that the shift should chosen be so that $A_{\sigma_{\mbox{\tiny S}}}$ is positive definite. The smallest shift for which $A_{\sigma_{\mbox{\tiny S}}}$ is guaranteed to be positive semi-definite is $\sigma_{\min}$ from (1.3) so any value slightly larger than this is possible. But a far cheaper option that does not rely on knowing the spectrum of $A$, once $A$ has been found not to be positive definite, is to use the value
$\sigma_{\mbox{\tiny G}}=-\min_{1\leq i\leq n}\left[a_{ii}-\sum_{1\leq j\neq i\leq n}|a_{i,j}|\right]+\epsilon_{\mbox{\tiny G}}$}
from the Gershgorin theorem, requiring a single pass through the values of $A$; we have found that this with the heuristic value $\epsilon_{\mbox{\tiny G}}=\sqrt{\epsilon_{\mbox{\tiny M}}}\max_{1\leq i,j\leq n}|a_{i,j}|$ where $\epsilon_{\mbox{\tiny M}}$ is the machine precision is effective.
As before, we generate the extended Krylov subspace ${\cal K}_{\ell}(A_{\sigma_{\mbox{\tiny S}}},b)$ using Algorithm 1. If the $\sigma$ we seek lies to the right of $\sigma_{\mbox{\tiny S}}$ (i.e., $\sigma\geq\sigma_{\mbox{\tiny S}}$ and hence $\theta\geq 0$), the Beckermann-Townsend [BeckTown17:simaa, §4.2] analysis outlined in Section 2 continues to hold, and we can expect that $x_{\sigma_{\mbox{\tiny S}}}(\theta)$ will lie in an approximately low-dimensional subspace that will be uncovered by Algorithm 1.
It is also worth noting that although we have now generated $P_{\ell}=V_{\ell}^{T}A_{\sigma_{\mbox{\tiny S}}}V_{\ell}$, we equally have
$V_{\ell}^{T}AV_{\ell}=V_{\ell}^{T}A_{\sigma_{\mbox{\tiny S}}}V_{\ell}-\sigma_{\mbox{\tiny S}}V_{\ell}^{T}V_{\ell}=P_{\ell}-\sigma_{\mbox{\tiny S}}I$
since the columns of are orthonormal. Thus subspace trust-region subproblems of the form (3.1) can be formulated as
(6.2) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}(P_{\ell}-\sigma_{\mbox{\tiny S}}I)y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
More particularly, if we rely on the spectral factorization (1.6), the diagonal subproblem (5.2) becomes
$z_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle z\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{\ell}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}z^{T}(\Lambda_{\ell}-\sigma_{\mbox{\tiny S}}I)z-d_{\ell}^{T}z\;\;\mbox{subject to}\;\;\|z\|\leq\Delta.$
Finally, we return briefly to the thorny issue of the so-called hard case, that is when $b$ is orthogonal to the subspace ${\cal E}_{1}$ of all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)<0$, and the radius is too large. In this case, the solution we seek is of the form [^7] Note that $x_{+}$ exists since $b$ lies in the range of $A-\lambda_{1}(A)I$ as $b$ is orthogonal to ${\cal E}_{1}$.
$x_{+}+\theta y,\;\;\mbox{where}\;\;x_{+}=\hskip-5.69054pt\hskip-5.69054pt\lim_{\sigma\stackrel{{\scriptstyle+}}{{\rightarrow}}-\lambda_{1}(A)}\hskip-5.69054pt\hskip-5.69054ptx(\sigma),\;\;y\in{\cal E}_{1},\;\;\mbox{and}\;\;\|x_{+}+\theta y\|=\Delta,$
when $\|x_{+}\|<\Delta$. No Krylov method based on $A$ and $b$ (extended or otherwise) will see the eigenspace ${\cal E}_{1}$ in the hard case in exact arithmetic, although rounding errors can gradually introduce it. Since we have never observed the hard case in practice for anything other than contrived examples, our only precaution is to add a tiny (pseudo-random) perturbation to $b$ if requested. We certainly do this when $b=0$.
6.1 Other elliptical norms
For the general problem defined in terms of an elliptical norm $\|x\|_{S}=\sqrt{x^{T}Sx}$ for a given symmetric positive definite $S=P^{T}P$, a similar approach is possible. Transforming via (5.7) as in Subsection 5.2, the (non-interior) optimality condition (1.2) becomes
(6.3) $(A_{\mbox{\tiny P}}+\sigma I)x_{\mbox{\tiny P}}(\sigma)=b_{\mbox{\tiny P}}\;\;\mbox{with}\;\;\|x_{\mbox{\tiny P}}(\sigma)\|=\Delta$}
for some
$\sigma\geq\sigma_{\mbox{\tiny P}}{}_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A_{\mbox{\tiny P}})),$}
or equivalently
$(A+\sigma S)x(\sigma)=b\;\;\mbox{with}\;\;\|x(\sigma)\|_{S}=\Delta;$}
note that $\lambda_{1}(A_{\mbox{\tiny P}})$ is equivalently the leftmost eigenvalue $\lambda_{1}(A,S)$ of the symmetric matrix pencil $A+\lambda S$. The equivalent version of (6.1) is then
(6.4) $(A_{\sigma_{\mbox{\tiny S}}}+\theta S)x_{\sigma_{\mbox{\tiny S}}}(\theta)=b\;\;\mbox{with}\;\;\|x_{\sigma_{\mbox{\tiny S}}}(\theta)\|_{S}=\Delta,$}
where
$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\mbox{\tiny P}}{}_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A,S)-\sigma_{\mbox{\tiny S}}).$}
Although it is less obvious how to find a lower bound on $\lambda_{1}(A,S)$ in general, it is possible to use a variant of the inexpensive Gershgorin bound in the commonly-occurring case where $S$ is strictly diagonally dominant (see [GoulRobiThor10:mpc, §3.3.2] for details). Our software packages make the strict diagonal dominance of $S$ a prerequisite.
7 Numerical experiments
We now provide some evidence that our new approach is a useful alternative to common existing methods. We consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables. For these, we generate the gradient and Hessian at the provided initial point, and use these for b-b and $A$ respectively.
We compare our new method (Algorithm 2 implemented as TREK), using the initial (Gershgorin-based) shifting strategy outlined in Section 6, with the multi-factorization method [GoulRobiThor10:mpc] available as TRS within GALAHAD [GoulOrbaToin03:toms] and the iterative method GLTR [GoulLuciRomaToin99:siopt] from the same library; GLTR only requires Hessian-vector products which can be an asset in some cases. Since all three methods can take advantage of information gathered when solving a problem with one radius and are now faced with another with identical data but a smaller radius, we consider each problem with radii, $\Delta=1$ and then $0.5$. TREK uses an iteration bound $m=100$, but almost always [^8] In four instances, $m$ was too small, but increasing the value to $400$ cured the problem. terminates with far-fewer $k$. Default, and roughly equivalent stopping rules were applied; TREK terminates as soon as (5.4) or (5.5) is smaller than $10^{-10}$.
We perform our experiments on four cores of a PC with sixteen Intel Core i9-9900 CPU 3.10GHz processors, and thirty two Gbytes of memory. The codes are all from the GALAHAD library compiled with -Ofast optimization by gfortran, and use is made of OpenMP-tuned BLAS and LAPACK. Factorization is performed, when needed, by the HSL_MA57 sparse symmetric linear solver [DuffReid96b:toms], enabled via GALAHAD’s SLS package, and relies on the tuned BLAS for good performance. On some examples, particularly those whose Hessians are banded, further experiments not listed here revealed that LAPACK’s PBTR band solver is a good alternative.
We summarize our findings in Figure 7.1 using the performance profile [dolamore02:mp] in which relative times are ranked. The profile is built using the data given in Appendix C.
Figure 7.1: Performance profile comparing three trust-region subproblem solvers on larger CUTEst examples.
We did not expect any of the methods tested to be an overall winner, and indeed this is the case. Clearly, the factorization-free method GLTR is frequently very effective, particularly when factorization is a significant overhead. But we were encouraged to see that TREK performs well in other cases, and the requirement of a single factorization often proves decisive when compared to the multi-factorization approach used by TRS. [^9] TRS also failed in three cases because the shifted matrix $A+\sigma I$ became too close to singularity to progress.
8 Regularization subproblems
Another important subproblem employed by methods for nonlinear optimization uses norm-regularization [CartGoulToin11:mp, CartGoulToin22, NestPoly06:mp]. Again, given an $n$ by $n$ symmetric (Hessian) matrix $A$, an $n$-vector $b$ and the Euclidean norm $\|\cdot\|$, but now a regularization weight $\rho>0$ and a power $r\geq 2$, the norm-regularization subproblem is to find
(8.1) $x_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x+\frac{\rho}{r}\|x\|^{r}.$}
Once again, the solution $x(\sigma)$ satisfies the first-order optimality condition (1.2), but now
(8.2) $\sigma=\rho\|x(\sigma)\|^{r-2}$}
for some $\sigma\geq\sigma_{\min}$ [CartGoulToin11:mp, NestPoly06:mp]. Thus the manifold of solutions is identical to that for the trust-region subproblem, but the one of interest satisfies (8.2) instead of (1.5); there is no interior-solution possibility in this case.
Since this is the case, it is trivial to recast the TREK algorithm for (8.1), the only differences are that the steps that calculate $y_{2k}$ and $y_{2k+1}$ and their shifts find instead
(8.3) $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y+\frac{\rho}{r}\|y\|^{r}\;\;\mbox{and its optimal shift $\sigma_{2k}$}\;\;$}
and
(8.4) $y_{2k+1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k+1}y-b_{2k+1}^{T}y+\frac{\rho}{r}\|y\|^{r}\;\;\mbox{and its optimal shift $\sigma_{2k+1},$}\;\;$}
and that the check for an interior solution before the main iteration loop is omitted. We refer to the resulting algorithm (and the GALAHAD package that implements it) as NREK. The subproblems (8.3) and (8.4) are solved just as those for the trust-region case by finding the spectral decompositions of $P_{2k}$ and $P_{2k+1}$, and transforming so that the resulting (small) subproblem is diagonal. Thereafter the high-order, factorization-free diagonal regularization subproblem subroutine RQS_solve_diagonal [GoulRobiThor10:mpc] from GALAHAD is employed. Other elliptical norms may be handled just as described in Subsection 5.2.
9 Conclusions
We have provided an alternative to currently popular methods for solving quadratic trust-region and regularization problems. This lies somewhere between multiple factorization methods that were first popularised by Moré and Sorensen [MoreSore83:sisc] and CG/Lanczos-based factorization-free methods such as GLTR [GoulLuciRomaToin99:siopt]. The novel features are firstly that it is shown that the manifold of all possible solutions lies on an (approximately) very low-rank manifold, and secondly that by extending the Krylov subspace to include products with $A^{-1}$ as well as $A$, this manifold may be computed very efficiently. This then adds a further tool to the arsenal of solvers that trust-region and cubic-regularization methods for general optimization methods rely on. Unsurprisingly, the new method is not always the best, but in some cases it is. It is also remarkable that empirically that the extended Krylov-subspace approach requires so few iterations to build a good approximation of the solution manifold; examination of the results in Appendix C illustrates only a few instances where more than 20 iterations are required. We provide analysis to explain why this should be so, although this may well be tightened in future.
Software that implements the algorithms will be available very shortly [^10] The TREK algorithm with its documentation is now available as a beta release, and once C, Python, Julia and Matlab interfaces have been finished, it will be simple to extend these to the norm regularisation case discussed in Section 8 as the NREK package. as part of the GALAHAD [GoulOrbaToin03:toms] fortran library, and there are interfaces to other popular languages such as C, Python and Julia, and tools such as Matlab. The codes are designed to solve not only a single instance, but a sequence in which the regularization is tightened, without recourse to expensive re-evaluations; the current subspace $V_{k}$ is extended if it is insufficient as the regularization changes. The new packages will be rolled out as optional subproblem solvers in other GALAHAD packages in due course.
Appendix A Convergence analysis
The first $j$-th matrix products of the EKS algorithm (Algorithm 1) generate an orthogonal basis for the extended Krylov subspace ${\cal K}_{j}(A,b)$. A natural question is then, how closely do the elements of this subspace lie to components of the manifold $x(\sigma)$. That is to say, for a given $\sigma$, is there an entry in ${\cal K}_{j}(A,b)$ close to $x(\sigma)$, and if so, how close?
Consider the symmetric positive-definite matrix $A$ with smallest and largest eigenvalues $\lambda_{1}=\lambda_{1}(A)$ and $\lambda_{n}=\lambda_{n}(A)$, and consider the interval ${\cal W}=[\lambda_{1},\lambda_{n}]\subset\mathbb{R_{+}^{*}}$. In this section, we repeat for completeness, with minor variations, the theory developed by Knizhnerman and Simoncini [KnizSimo11:NUMMAT] to prove that the convergence of the extended-Krylov-subspace method to build an approximate basis for the solution manifold $\{(A+\sigma I)^{-1}b\}$, for all $\sigma\notin-{\cal W}-(-\delta_{1},\delta_{n})=-(\lambda_{1}-\delta_{1},\lambda_{n}+\delta_{n})$, and $\delta_{1},\delta_{n}>0$, is linear at worst. The proof relies heavily on 19th to mid 20th century complex analysis, and we refer the reader to [KnizSimo11:NUMMAT] for the details; the technical buildup may appear somewhat fearsome to the uninitiated.
Let $D$ denote the closed unit disk, let $\Psi:\bar{\mathbb{C}}\setminus D\rightarrow\bar{\mathbb{C}}\setminus W$ be the Riemann mapping for $\cal W$} that preserves infinity and has positive derivatives there and let $\Phi=\Psi^{-1}$ be its inverse. That is, $\Psi=\Phi^{-1}$ and both $\Phi$ and $\Psi$ are holomorphic on their domains. For the specific set $\cal W$}, we have that $\Phi(t)$ and $\Psi(t)$ are real $t\notin{\cal W}$, and explicitly
$\Psi(t)=c+\frac{r}{2}\left(t+\frac{1}{t}\right)\text{ and}$
$\Phi(t)=\begin{cases}\displaystyle\frac{t-c}{r}+\sqrt{\left(\displaystyle\frac{t-c}{r}\right)^{2}-1}&\text{ for }t\leq-\lambda_{n}\\ \left(\displaystyle\frac{t-c}{r}+\sqrt{\left(\displaystyle\frac{t-c}{r}\right)^{2}-1}\right)^{-1}&\text{ for }t\geq-\lambda_{1}\end{cases}$
where $c=(\lambda_{n}+\lambda_{1})/2$ and $r=(\lambda_{n}-\lambda_{1})/2$. We then define the function (a finite Blaschke product)
$B(w)=w\frac{1-\overline{\Phi(0)}w}{\Phi(0)-w},$
and note that this function satisfies $|B(w)|=1$ for $|w|=1$ and
(A.1) B(w)<1{B(w)<1}
for $|w|<1$ (see, [KnizSimo11:NUMMAT]).
Armed with these definitions, let $\phi_{2\ell}$ and $\phi_{2\ell+1}$ be the (Takenaka–Malmquist) set of rational functions
(A.2) $\phi_{2\ell}(w)=B(w)^{\ell}\text{ and }\phi_{2\ell+1}(w)=-\frac{\sqrt{1-|\Phi(0)|^{-2}}}{1-\Phi(0)^{-1}w}wB(w)^{\ell}.$}
for $\ell\in\mathbb{N}$. For these, we have
$\frac{1}{2\pi}\int_{|w|=1}\overline{\phi_{p}(w)}\phi_{m}(w)|dw|=\delta_{m,p}$
and
$\max_{m\in\mathbb{N}}\max_{|w|=1}|\phi_{m}(w)|\leq c_{\phi}$
for all $m,p\in\mathbb{N}$ and some finite constant $c_{\phi}$.
Now consider the Faber-Dzhrbashyan rational functions
$M_{m}(z)=\frac{1}{2\pi i}\int_{\Gamma_{R}}\frac{\phi_{m}(\Phi(\xi))}{\xi-z}d\xi\text{ for }z\notin\overline{G_{R}}\text{ and }m\in\mathbb{N},$
where
$\Gamma_{R}=\{z\in\mathbb{C}\setminus W:|\Phi(z)|=R\},\;\;G_{R}=W\cup\{z\in\mathbb{C}\setminus W:|\Phi(z)|<R\}$
and $R>1$. In particular, $M_{m}$ is a Faber transformation of $\phi_{m}$, and it then follows that
(A.3) $M_{2k}(z)=\frac{p_{2k}(z)}{z^{k}}\text{ and }M_{2k+1}(z)=\frac{p_{2k+1}(z)}{z^{k+1}}$}
where $p_{m}$ is a polynomial of degree $m$. Furthermore,
(A.4) $|M_{m}(z)|\leq c_{M}$}
for some finite constant $c_{M}$ for all $m$ and all $z\in{\cal W}$.
Proposition A.1.
Given $A$, $\Phi$, $\Psi$, $\phi_{k}$, $M_{k}$ and $\cal W$} as above, we have that
(A.5) $(zI-A)^{-1}=\frac{1}{\Phi(z)\Psi^{\prime}(\Phi(z))}\sum_{k=0}^{\infty}\phi_{k}(\Phi(z)^{-1})M_{k}(A)\text{ for every }z\notin{\cal W}$}
Proof A.2.
In general, we have the identity [Koch58]
$\frac{\Psi^{\prime}(w)}{\Psi(w)-u}=\frac{1}{w}\sum_{k=0}^{\infty}\overline{\phi_{k}\left(\frac{1}{\bar{w}}\right)}M_{k}(u).$
The special (matrix-valued) case for which $u=A$ and $\Psi(w)=zI$ leads directly to (A.5).
Proposition A.3.
Let $\sigma\in\mathbb{R}\setminus(-\lambda_{n}-\delta_{n},-\lambda_{1}+\delta_{1})$. The EKS algorithm (Algorithm 1) used to approximate $x(\sigma)=(A+\sigma I)^{-1}b$ converges at worst linearly with a rate
(A.6) $\mu=\sqrt{|E(\sigma)|}<1,$}
where $E:\mathbb{R}\setminus[-\lambda_{n},-\lambda_{1}]\rightarrow\mathbb{R}$ is
$E(\sigma)\stackrel{{\scriptstyle\rm def}}{{=}}B\left(\Phi(-\sigma)^{-1}\right).$
Furthermore, if $\sigma\geq 0$,
(A.7) $\mu\leq\frac{\kappa^{1/4}-1}{\kappa^{1/4}+1},$}
where $\kappa=\lambda_{n}/\lambda_{1}$.
Proof A.4.
Employing (A.5) with $z=-\sigma$ immediately gives
(A.8) $\begin{array}[]{rl}x(\sigma)\hskip-5.69054pt&=-((-\sigma)I-A)^{-1}b=\displaystyle\frac{1}{\Phi(-\sigma)\Psi^{\prime}(\Phi(-\sigma))}\displaystyle\sum_{i=0}^{\infty}\phi_{i}(\Phi(-\sigma)^{-1})M_{i}(A)b\\ &=\displaystyle\sum_{i=0}^{\infty}a_{i}M_{i}(A)b,\end{array}$}
where
$a_{i}=\frac{\phi_{i}(\Phi(-\sigma)^{-1})}{\Phi(-\sigma)\Psi^{\prime}(\Phi(-\sigma))}.$
Using (A.2) with $w=-\sigma$, we may bound
(A.9) $|a_{i}|\leq\frac{|\phi_{i}(\Phi(-\sigma)^{-1})|}{|\Phi(-\sigma)||\Psi^{\prime}(\Phi(-\sigma))|}\leq\frac{|B\left(\Phi(-\sigma)^{-1}\right)^{i/2}|}{|\Phi(-\sigma)||\Psi^{\prime}(\Phi(-\sigma))|}\leq c_{1}\mu^{i},$}
where
$c_{1}=\frac{1}{\Psi^{\prime}\left(\Phi(-\lambda_{1}-\delta_{1})\right)}$
is a constant, and
(A.10) $\mu=\sqrt{|B\left(\Phi(-\sigma)^{-1}\right)|}\equiv\sqrt{|E(\sigma)|}<1$}
from (A.1) as $|\Phi(-\sigma)^{-1}|<1$.
Since $M_{i}(A)b\in{\cal K}_{i}(A,b)$ for $i=1,\ldots,2k$ by definition (4.1) of ${\cal K}_{i}(A,b)$ and from (A.3), the estimate
$x_{j}\stackrel{{\scriptstyle\rm def}}{{=}}\sum_{i=0}^{j}a_{i}M_{i}(A)b$
is in ${\cal K}_{j}(A,b)$ for $j=1,\ldots 2k$. But then we have
$\|x(\sigma)-x_{j}\|\leq\sum_{i=j+1}^{\infty}|a_{i}|\|M_{i}(A)\|\|b\|\leq c\mu^{j}.$
from (A.8), (A.4) and and (A.9) respectively, where $c=c_{M}c_{1}\|b\|\mu/(1-\mu)$ as $\mu<1$.
A simple analysis, e.g., [KnizSimo11:NUMMAT, Prop.4.1] shows that the function $E$ evolves as follows:
$\sigma$ -$\infty$ $\lambda_{n}$ $\lambda_{1}$ 0 $\sqrt{\lambda_{1}\lambda_{n}}$ $\infty$
$E(\sigma)$ 0 $\searrow$ 1-1 1-1 $\nearrow$ 0 $\nearrow$ $\left(\frac{\kappa^{1/4}-1}{\kappa^{1/4}+1}\right)^{2}$ $\searrow$ 0
where $\kappa=\frac{\lambda_{n}}{\lambda_{1}}$; as an illustration, Figure A.1 depicts the function $E$ on $\mathbb{R}\setminus(-\lambda_{n},-\lambda_{1})$ with $\lambda_{1}=5$ and $\lambda_{n}=100$. Therefore, for symmetric positive-definite AA and a shift $\sigma\notin[-\lambda_{n},-\lambda_{1}]$, the convergence rate of EKS in approximating the solution of the shifted linear system $(A+\sigma I)x=b$ is $\sqrt{|E(\sigma)|}$, and if $\sigma\geq 0$, we have $E(\sigma)\leq\left(\frac{\kappa^{1/4}-1}{\kappa^{1/4}+1}\right)^{2}$, and thus (A.7) follows from (A.10).
Figure A.1: Upper bound on the convergence rate of extended Krylov method to solve the shifted linear system with a shift $\sigma$. The spectrum of the matrix is in $[5,100]$.
Theorem A.5.
The extended Krylov method to approximate a solution manifold
$\{x(\sigma)\}_{\sigma\notin(-\lambda_{n}-\delta_{n},-\lambda_{1}+\delta_{1})}$ converges at worst linearly with a convergence rate
$\mu=\max\left(\sqrt{|E(-\lambda_{n}-\delta_{n})|},\sqrt{|E(-\lambda_{1}+\delta_{1})|},\left(\frac{\kappa^{1/4}-1}{\kappa^{1/4}+1}\right)\right),$
where $E$ is defined in Proposition A.3 and $\kappa=\lambda_{n}/\lambda_{1}$.
Proof A.6.
The proof follows straightforwardly from Proposition A.3, and the discussion of the evolution of $E(\sigma)$ given there.
Appendix B Modified norm trust region subproblem solver
Here we provide a detailed description of the modifications to the TREK algorithm (Algorithm 2) that are necessary when using the trust region norm for which $\|x\|_{S}^{2}=x^{T}Sx$, with $S=P^{T}P$, as discussed in Section 5.2. As we stated there, it suffices to apply Algorithm 2 with the matrix $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and vector $b_{\mbox{\tiny P}}=P^{-T}b$, and to recover the solution via $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$. Since $S^{-1}=P^{-1}P^{-T}$, this is formally as Algorithm 4 4.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta.$
2mm Algorithm:
$x:=A^{-1}b$, $w:=Px$;
if $\|w\|\leq\Delta$ then
 exit with the interior solution $x_{*}=x$ and shift $\sigma_{*}=0$;
$\delta_{0}:=\|P^{-T}b\|$, $v_{0}:=P^{-T}b/\delta_{0}$, $u:=w/\delta_{0}$;
$\beta_{0}:=u^{T}v_{0}$;
$u\leftarrow u-\beta_{0}v_{0}$;
$\delta_{-1}:=\|u\|$;
for $k=1,2,\ldots,m$ do
 if $\delta_{-k}>0$ then
    $v_{-k}:=u/\delta_{-k}$;
    $u:=P^{-T}AP^{-1}v_{-k}$;
    $\alpha_{k-1}:=u^{T}v_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
    $\alpha_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\alpha_{-k}v_{-k}$;
    $\delta_{k}:=\|u\|$;
 if $k>1$ then
     update $P_{2k-1}$ from $P_{2k-2}$ using (4.13) and form $b_{2k-1}=(b^{T}_{2k-2},0)^{T}$;
 else
     initialize $P_{1}$ from (4.11) and set $b_{1}=\delta_{0}$;
 $y_{2k-1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k-1}y-b_{2k-1}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ and its optimal shift $\sigma_{2k-1}$;
   compute $\|r_{2k-1}\|$ from (5.4);
 if $\|r_{2k-1}\|\leq\epsilon$ then
    exit with the interior solution $x_{*}=V_{2k-1}y_{2k-1}$, shift $\sigma_{*}=\sigma_{2k-1}$ and $k_{\mbox{\scriptsize stop}}=k$;
 if $\delta_{k}>0$ then
    $v_{k}:=u/\delta_{k}$, $u:=PA^{-1}P^{T}v_{k}$;
    $\beta_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$;
    $\beta_{k}:=u^{T}v_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
 if $k<m$ then
    if $\delta_{k}>0$ then
       $\delta_{-k-1}:=\|u\|$;
      update $P_{2k}$ from $P_{2k-1}$ using (4.14) and form $b_{2k}=(b^{T}_{2k-1},0)^{T}$;
    $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ together with its optimal shift $\sigma_{2k}$;
      compute $\|r_{2k}\|$ from (5.5);
    if $\|r_{2k}\|\leq\epsilon$ then
       exit with the interior solution $x_{*}=V_{2k}y_{2k}$, shift $\sigma_{*}=\sigma_{2k}$ and $k_{\mbox{\scriptsize stop}}=k$;
Algorithm 4 The TREK algorithm to solve the trust-region subproblem (5.6)
Under the transformation $w_{k}=P^{-1}v_{k}$ i.e., $v_{k}=Pw_{k}$, and $u=Pv$, i.e., $v=P^{-1}u$ (and then replacing $v_{k}\rightarrow w_{k}$ and $u\rightarrow v$ so that the resulting changes may be easily compared with Algorithm 2), and introducing temporary workspace vectors $z$ and $s$, we may rewrite Algorithm 4 as Algorithm 5 5.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|_{S}\leq\Delta.$
2mm Algorithm:
$x:=A^{-1}b$, $u:=Sx$;
if $\sqrt{u^{T}x}\leq\Delta$ then
 exit with the interior solution $x_{*}=x$ and shift $\sigma_{*}=0$;
$w:=S^{-1}b$, $\delta_{0}:=\sqrt{b^{T}w}$;
$v_{0}:=w/\delta_{0}$, $w:=u/\delta_{0}$, $u:=x/\delta_{0}$;
$\beta_{0}:=v_{0}^{T}Su$, $u\leftarrow u-\beta_{0}v_{0}$;
$w=Su$, $\delta_{-1}:=\sqrt{u^{T}w}$;
for $k=1,2,\ldots,m$ do
 if $\delta_{-k}>0$ then
    $v_{-k}:=u/\delta_{-k}$, $w:=Av_{-k}$, $u:=S^{-1}w$;
    $\alpha_{-k}:=v_{-k}^{T}Su$, $u\leftarrow u-\alpha_{-k}v_{-k}$;
    $\alpha_{k-1}:=v_{k-1}^{T}Su$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
    $w=Su$, $\delta_{k}:=\sqrt{u^{T}w}$;
 if $k>1$ then
     update $P_{2k-1}$ from $P_{2k-2}$ using (4.13) and form $b_{2k-1}=(b^{T}_{2k-2},0)^{T}$;
 else
     initialize $P_{1}$ from (4.11) and set $b_{1}=\delta_{0}$;
 $y_{2k-1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k-1}y-b_{2k-1}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ and its optimal shift $\sigma_{2k-1}$;
   compute $\|r_{2k-1}\|$ from (5.4);
 if $\|r_{2k-1}\|\leq\epsilon$ then
    exit with the interior solution $x_{*}=V_{2k-1}y_{2k-1}$, shift $\sigma_{*}=\sigma_{2k-1}$ and $k_{\mbox{\scriptsize stop}}=k$;
 if $\delta_{k}>0$ then
    $v_{k}:=u/\delta_{k}$, $w:=Sv_{k}$, $u:=A^{-1}w$;
    $\beta_{-k}:=v_{-k}^{T}Su$, $u\leftarrow u-\beta_{-k}v_{-k}$;
    $\beta_{k}:=v_{k}^{T}Su$, $u\leftarrow u-\beta_{k}v_{k}$;
 if $k<m$ then
    if $\delta_{k}>0$ then
       $w=Su$, $\delta_{-k-1}:=\sqrt{u^{T}w}$;
      update $P_{2k}$ from $P_{2k-1}$ from (4.12) and form $b_{2k}=(b^{T}_{2k-1},0)^{T}$;
    $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ together with its optimal shift $\sigma_{2k}$;
      compute $\|r_{2k}\|$ from (5.5);
    if $\|r_{2k}\|\leq\epsilon$ then
       exit with the interior solution $x_{*}=V_{2k}y_{2k}$, shift $\sigma_{*}=\sigma_{2k}$ and $k_{\mbox{\scriptsize stop}}=k$;
Algorithm 5 The TREK algorithm to solve the trust-region subproblem (5.9)
While it might appear that Algorithm 5 requires multiple products of $S$ with selected vectors, as needed to calculate the terms $v_{i}^{T}Su$ that define $\alpha_{i}$ and $\beta_{i}$, fortunately this is not the case. There are two ways to avoid this. In the first, we maintain appropriate vectors $q_{i}\stackrel{{\scriptstyle\rm def}}{{=}}Sv_{i}$. We record this as Algorithm 6 6.
Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|_{S}\leq\Delta.$
2mm Algorithm:
$x:=A^{-1}b$, $u:=Sx$;
if $\sqrt{u^{T}x}\leq\Delta$ then
 exit with the interior solution $x_{*}=x$ and shift $\sigma_{*}=0$;
$w:=S^{-1}b$, $\delta_{0}:=\sqrt{b^{T}w}$;
$v_{0}:=w/\delta_{0}$, $q_{0}:=b/\delta_{0}$, $w:=u/\delta_{0}$, $u:=x/\delta_{0}$;
$\beta_{0}:=u^{T}q_{0}$, $u\leftarrow u-\beta_{0}v_{0}$;
$w=Su$, $\delta_{-1}:=\sqrt{u^{T}w}$;
for $k=1,2,\ldots,m$ do
 if $\delta_{-k}>0$ then
    $v_{-k}:=u/\delta_{-k}$, $q_{-k}:=w/\delta_{-k}$;
    $w:=Av_{-k}$, $u:=S^{-1}w$;
    $\alpha_{-k}:=u^{T}q_{-k}$, $u\leftarrow u-\alpha_{-k}v_{-k}$;
    $\alpha_{k-1}:=u^{T}q_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
    $w=Su$, $\delta_{k}:=\sqrt{u^{T}w}$;
 if $k>1$ then
     update $P_{2k-1}$ from $P_{2k-2}$ using (4.13) and form $b_{2k-1}=(b^{T}_{2k-2},0)^{T}$;
 else
     initialize $P_{1}$ from (4.11) and set $b_{1}=\delta_{0}$;
 $y_{2k-1}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k-1}y-b_{2k-1}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ and its optimal shift $\sigma_{2k-1}$;
   compute $\|r_{2k-1}\|$ from (5.4);
 if $\|r_{2k-1}\|\leq\epsilon$ then
    exit with the interior solution $x_{*}=V_{2k-1}y_{2k-1}$, shift $\sigma_{*}=\sigma_{2k-1}$ and $k_{\mbox{\scriptsize stop}}=k$;
 if $\delta_{k}>0$ then
    $v_{k}:=u/\delta_{k}$, $q_{k}:=w/\delta_{k}$;
    $u:=A^{-1}q_{k}$;
    $\beta_{-k}:=u^{T}q_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$;
    $\beta_{k}:=u^{T}q_{k}$, $u\leftarrow u-\beta_{k}v_{k}$;
 if $k<m$ then
    if $\delta_{k}>0$ then
       $w=Su$, $\delta_{-k-1}:=\sqrt{u^{T}w}$;
      update $P_{2k}$ from $P_{2k-1}$ using (4.14) and form $b_{2k}=(b^{T}_{2k-1},0)^{T}$;
    $y_{2k}:=\mbox{arg min}{\scriptstyle\frac{1}{2}}y^{T}P_{2k}y-b_{2k}^{T}y\;\mbox{s.t.}\;\|y\|\leq\Delta$ together with its optimal shift $\sigma_{2k}$;
      compute $\|r_{2k}\|$ from (5.5);
    if $\|r_{2k}\|\leq\epsilon$ then
       exit with the interior solution $x_{*}=V_{2k}y_{2k}$, shift $\sigma_{*}=\sigma_{2k}$ and $k_{\mbox{\scriptsize stop}}=k$;
Algorithm 6 The TREK algorithm to solve the trust-region subproblem (5.9)
In practice, use two arrays $q_{+}$ and $q_{-}$, and store and successively overwrite $q_{+}=q_{k}$ for $k\geq 0$ and $q_{-}=q_{k}$ for $k<0$ with $k=1,2,\ldots,m$.
The second way to avoid multiple products with $S$ is to use $q_{i}$ as before, but now to update $w\stackrel{{\scriptstyle\rm def}}{{=}}Su$ as it appears. We record this variant as Algorithm 3 3. Once again, in practice, use two arrays $q_{+}$ and $q_{-}$, and store and successively overwrite $q_{+}=q_{k}$ for $k\geq 0$ and $q_{-}=q_{k}$ for $k<0$.
Computationally, the main difference is that Algorithm 6 requires an additional $S$ product per cycle, while Algorithm 3 has an additional four vector subtractions.
Appendix C Detailed numerical experiments
Here we summarize the results from our numerical experiments that lead to Figure 7.1. As we said, we consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables, evaluate the gradient and Hessian to provide b-b and $A$, and for each a pair of radii $\Delta=1$ and then $0.5$. We apply the GALAHAD solvers TREK (Algorithm 2), TRS [GoulRobiThor10:mpc] and GLTR [GoulLuciRomaToin99:siopt] in each case; when $\Delta=0.5$, we use data accumulated from the $\Delta=1$ case to “warm start” the second solve as each solver provides for this.
In Table LABEL:table-a1 we report the name of each test problem, the size of the instance chosen, $n$ (many CUTEst examples have multiple dimensions), the radius value, $\Delta$ and the optimal objective function found. For TREK, we give the terminal iteration count, $k_{\mbox{\scriptsize stop}}$, which is the number of solves with $A$ required, and the clock time required. For TRS, we report the number of factorizations and clock time required, while for GLTR, we state the number of vector products involving $A$ and again the time. The symbol $\dagger$} indicates an instance that the iteration bound $m=100$ was reached, while and $\ddagger$} indicates that a factorization failed through singularity.
Table C.1: Subproblem solves for CUTEst examples: Complete results
TREK TRS GLTR
name $n$ $\Delta$ $f(x_{*})$ $k_{\mbox{\tiny stop}}$ time facts time prods time
ARWHEAD 5000 1.0 -9.99800000E+03 0 0.00 2 0.00 2 0.00
ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
BA-L16LS 66462 1.0 -4.00647969E+09 10 1.18 6 1.94 69 2.76
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 1.43 34 1.42
BA-L16LS 66462 1.0 -4.00647969E+09 10 0.82 6 0.94 69 2.55
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.68 34 1.34
ARWHEAD 5000 1.0 -9.99800000E+03 -1 0.00 2 0.00 2 0.00
ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
BA-L16LS 66462 1.0 -4.00647969E+09 10 0.84 6 0.94 69 2.60
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.73 34 1.32
BA-L21LS 34134 1.0 -1.43350953E+08 100† 0.88 8 0.39 1129 18.09
BA-L21LS 34134 0.5 -1.27987567E+08 100† 0.00 12 0.27 564 8.95
BA-L49LS 23769 1.0 -2.32581840E+10 6 0.25 8 0.45 27 0.40
BA-L49LS 23769 0.5 -5.82477997E+09 6 0.00 11 0.20 13 0.21
BA-L52LS 192627 1.0 -1.78720629E+10 31 4.65 11 9.18 173 27.82
BA-L52LS 192627 0.5 -1.26606809E+10 31 0.01 16 4.79 86 13.82
BA-L73LS 33753 1.0 -9.88938047E+10 4 0.37 10 0.93 25 0.56
BA-L73LS 33753 0.5 -2.47259422E+10 4 0.00 13 0.35 12 0.29
BDQRTIC 5000 1.0 -4.70328224E+05 4 0.01 4 0.01 13 0.00
BDQRTIC 5000 0.5 -4.37969301E+05 4 0.00 8 0.01 6 0.00
BOX 10000 1.0 -4.99665149E+01 1 0.01 3 0.01 3 0.00
BOX 10000 0.5 -2.49841935E+01 1 0.00 6 0.01 1 0.00
BOXPOWER 20000 1.0 -1.27930301E+05 3 0.01 3 0.02 5 0.00
BOXPOWER 20000 0.5 -7.39842920E+04 3 0.00 6 0.02 2 0.00
BROYDN3DLS 5000 1.0 -5.47141790E+02 6 0.01 3 0.02 19 0.00
BROYDN3DLS 5000 0.5 -2.78620552E+02 6 0.00 6 0.01 9 0.00
BROYDN7D 5000 1.0 -1.04198440E+03 4 0.01 3 0.01 11 0.00
BROYDN7D 5000 0.5 -5.29359803E+02 4 0.00 6 0.01 5 0.00
BROYDNBDLS 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
BROYDNBDLS 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
BRYBND 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
BRYBND 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
CHAINWOO 4000 1.0 -4.18808861E+05 4 0.01 2 0.01 9 0.00
CHAINWOO 4000 0.5 -2.10745351E+05 4 0.00 4 0.01 4 0.00
COSINE 10000 1.0 -7.33802606E+01 5 0.03 2 0.02 11 0.00
COSINE 10000 0.5 -3.63234227E+01 5 0.00 4 0.01 5 0.00
CRAGGLVY 5000 1.0 -2.70357001E+05 2 0.02 3 0.01 5 0.00
CRAGGLVY 5000 0.5 -1.38612789E+05 2 0.00 6 0.01 2 0.00
CURLY10 10000 1.0 -2.55466912E+03 52 0.11 3 0.05 157 0.04
CURLY10 10000 0.5 -6.72371854E+02 52 0.00 6 0.04 78 0.02
CURLY20 10000 1.0 -9.12107858E+03 49 0.11 3 0.05 205 0.07
CURLY20 10000 0.5 -2.35575388E+03 49 0.00 6 0.04 102 0.03
CURLY30 10000 1.0 -1.97300910E+04 41 0.15 3 0.10 235 0.11
CURLY30 10000 0.5 -5.06068395E+03 41 0.00 6 0.06 117 0.04
DIXMAANB 3000 1.0 -1.94571746E+03 4 0.01 3 0.01 11 0.00
DIXMAANB 3000 0.5 -9.82394549E+02 4 0.00 6 0.01 5 0.00
DIXMAANC 3000 1.0 -3.67420974E+03 4 0.02 3 0.01 11 0.00
DIXMAANC 3000 0.5 -1.85594245E+03 4 0.00 6 0.01 5 0.00
DIXMAAND 3000 1.0 -7.40783941E+03 4 0.01 3 0.01 11 0.00
DIXMAAND 3000 0.5 -3.74285037E+03 4 0.00 6 0.01 5 0.00
DIXMAANE1 3000 1.0 -1.03848735E+03 4 0.01 3 0.00 11 0.00
DIXMAANE1 3000 0.5 -5.25114487E+02 4 0.00 6 0.00 5 0.00
DIXMAANF 3000 1.0 -1.83737329E+03 4 0.01 3 0.01 11 0.00
DIXMAANF 3000 0.5 -9.28137690E+02 4 0.00 6 0.01 5 0.00
DIXMAANG 3000 1.0 -3.56193345E+03 4 0.01 3 0.01 11 0.00
DIXMAANG 3000 0.5 -1.79971801E+03 4 0.00 6 0.01 5 0.00
DIXMAANH 3000 1.0 -7.28769893E+03 4 0.01 3 0.01 11 0.00
DIXMAANH 3000 0.5 -3.68269064E+03 4 0.00 6 0.01 5 0.00
DIXMAANI1 3000 1.0 -1.00064454E+03 4 0.01 3 0.00 11 0.00
DIXMAANI1 3000 0.5 -5.06141189E+02 4 0.00 6 0.00 5 0.00
DIXMAANJ 3000 1.0 -1.79984433E+03 4 0.01 3 0.01 11 0.00
DIXMAANJ 3000 0.5 -9.09324798E+02 4 0.00 6 0.01 5 0.00
DIXMAANK 3000 1.0 -3.52376478E+03 4 0.01 3 0.01 11 0.00
DIXMAANK 3000 0.5 -1.78058448E+03 4 0.00 6 0.01 5 0.00
DIXMAANL 3000 1.0 -7.24829529E+03 4 0.01 3 0.01 11 0.00
DIXMAANL 3000 0.5 -3.66293885E+03 4 0.00 6 0.01 5 0.00
DIXMAANM1 3000 1.0 -4.27493497E+02 5 0.01 3 0.01 13 0.00
DIXMAANM1 3000 0.5 -2.16348948E+02 5 0.00 6 0.01 6 0.00
DIXMAANN 3000 1.0 -9.97842749E+02 5 0.01 3 0.01 13 0.00
DIXMAANN 3000 0.5 -5.05235340E+02 5 0.00 6 0.01 6 0.00
DIXMAANO 3000 1.0 -1.90145642E+03 5 0.01 3 0.01 13 0.00
DIXMAANO 3000 0.5 -9.63139983E+02 5 0.00 6 0.01 6 0.00
DIXMAANP 3000 1.0 -3.85352076E+03 5 0.01 3 0.01 13 0.00
DIXMAANP 3000 0.5 -1.95233964E+03 5 0.00 6 0.01 6 0.00
DIXON3DQ 10000 1.0 -4.35180402E+00 19 0.02 5 0.03 45 0.01
DIXON3DQ 10000 0.5 -2.48148299E+00 19 0.00 8 0.02 22 0.00
DQDRTIC 5000 1.0 -8.50546818E+04 3 0.00 2 0.00 7 0.00
DQDRTIC 5000 0.5 -4.25775883E+04 3 0.00 4 0.00 3 0.00
DQRTIC 5000 1.0 -1.33489191E+13 2 0.00 2 0.00 5 0.00
DQRTIC 5000 0.5 -6.67448869E+12 2 0.00 4 0.00 2 0.00
EDENSCH 2000 1.0 -9.90061935E+04 3 0.00 2 0.00 7 0.00
EDENSCH 2000 0.5 -4.96303271E+04 3 0.00 4 0.00 3 0.00
EG2 1000 1.0 -1.73066127E+02 0 0.00 2 0.00 1 0.00
EG2 1000 0.5 -1.64667237E+02 1 0.00 4 0.00 0 0.00
EIGENALS 2550 1.0 -1.01221801E+03 7 0.75 3 0.85 25 0.04
EIGENALS 2550 0.5 -4.77552997E+02 7 0.00 6 0.79 12 0.02
EIGENBLS 2550 1.0 -3.52937908E+01 8 0.70 3 0.89 25 0.04
EIGENBLS 2550 0.5 -1.80972156E+01 8 0.00 5 0.62 12 0.02
EIGENCLS 2652 1.0 -4.70591220E+02 9 0.78 3 0.91 31 0.04
EIGENCLS 2652 0.5 -2.33514688E+02 9 0.00 6 0.93 15 0.02
ENGVAL1 5000 1.0 -8.67081566E+03 3 0.01 2 0.01 9 0.00
ENGVAL1 5000 0.5 -4.35940622E+03 3 0.00 4 0.01 4 0.00
EXTROSNB 1000 1.0 -3.66203611E+04 4 0.00 3 0.00 11 0.00
EXTROSNB 1000 0.5 -1.86350898E+04 4 0.00 6 0.00 5 0.00
FLETBV3M 5000 1.0 -4.36166522E+01 3 0.01 2 0.01 7 0.00
FLETBV3M 5000 0.5 -2.18337960E+01 3 0.00 4 0.01 3 0.00
FLETCBV2 5000 1.0 -3.88209445E-06 13 0.01 4 0.01 9737 412.07
FLETCBV2 5000 0.5 -2.02683952E-06 16 0.00 8 0.01 4868 0.53
FLETCBV3 5000 1.0 -4.36172752E+01 3 0.01 2 0.01 7 0.04
FLETCBV3 5000 0.5 -2.18341109E+01 3 0.00 4 0.01 3 0.00
FLETCHBV 5000 1.0 -2.74868337E+09 3 0.01 2 0.01 9 0.00
FLETCHBV 5000 0.5 -1.37709990E+09 3 0.00 4 0.01 4 0.00
FLETCHCR 1000 1.0 -1.08786732E+01 1 0.00 3 0.00 3 0.00
FLETCHCR 1000 0.5 -1.05114880E+01 1 0.00 7 0.00 1 0.00
FMINSRF2 5625 1.0 -2.74210232E-01 44 0.04 4 0.03 107 0.03
FMINSRF2 5625 0.5 -1.42608286E-01 44 0.00 8 0.03 53 0.01
FMINSURF 5625 1.0 -2.74396123E-01 44 5.28 4 8.28 107 0.03
FMINSURF 5625 0.5 -1.42697922E-01 44 0.00 8 8.05 53 0.01
FREUROTH 5000 1.0 -5.51793805E+04 3 0.01 3 0.01 7 0.00
FREUROTH 5000 0.5 -2.75854293E+04 3 0.00 6 0.01 3 0.00
GENHUMPS 5000 1.0 -6.64118303E+03 6 0.03 3 0.02 13 0.00
GENHUMPS 5000 0.5 -3.16552923E+03 6 0.00 6 0.01 6 0.00
INDEF 5000 1.0 -2.10493952E+03 6 0.01 103‡ 0.12 13 0.00
INDEF 5000 0.5 -5.26825583E+02 6 0.00 207‡ 0.25 6 0.00
INDEFM 100000 1.0 -4.20748078E+04 3 0.09 11 0.38 11 0.05
INDEFM 100000 0.5 -1.05192638E+04 3 0.00 16 0.23 5 0.03
JIMACK 3549 1.0 -4.37368349E-01 100† 0.45 26 0.57 3201 17.07
JIMACK 3549 0.5 -2.41925217E-01 100† 0.00 31 0.14 1600 0.62
LIARWHD 5000 1.0 -4.61798034E+05 1 0.00 3 0.00 3 0.01
LIARWHD 5000 0.5 -2.36034628E+05 1 0.00 6 0.00 1 0.00
MODBEALE 20000 1.0 -3.04981491E+05 2 0.05 2 0.03 7 0.01
MODBEALE 20000 0.5 -1.52960464E+05 2 0.00 4 0.03 3 0.00
MOREBV 5000 1.0 -7.72418940E-14 9 0.01 107‡ 0.13 295 0.17
MOREBV 5000 0.5 -7.72418940E-14 54 0.03 116 0.02 294 0.02
MSQRTALS 1024 1.0 -3.70040042E+02 4 0.16 4 0.14 27 0.01
MSQRTALS 1024 0.5 -1.75608597E+02 5 0.00 7 0.11 13 0.01
MSQRTBLS 1024 1.0 -3.69312117E+02 4 0.16 4 0.14 27 0.01
MSQRTBLS 1024 0.5 -1.75279355E+02 5 0.00 7 0.12 13 0.01
NCB20 5010 1.0 -2.77922014E+02 6 0.03 4 0.02 27 0.07
NCB20 5010 0.5 -1.39893162E+02 6 0.00 7 0.02 13 0.03
NCB20B 5000 1.0 -2.77950996E+02 6 0.03 4 0.02 27 0.05
NCB20B 5000 0.5 -1.39907519E+02 6 0.00 7 0.02 13 0.02
NONCVXU2 5000 1.0 -3.33554888E+06 2 0.21 2 0.34 5 0.00
NONCVXU2 5000 0.5 -1.66777663E+06 2 0.00 4 0.34 2 0.00
NONCVXUN 5000 1.0 -3.56003262E+06 2 0.02 2 0.02 5 0.00
NONCVXUN 5000 0.5 -1.78001885E+06 2 0.00 4 0.02 2 0.00
NONDIA 5000 1.0 -1.49970308E+06 1 0.01 3 0.01 3 0.00
NONDIA 5000 0.5 -8.75226609E+05 1 0.00 6 0.01 1 0.00
NONDQUAR 5000 1.0 -3.33683482E+03 43 0.03 10 0.02 117 0.01
NONDQUAR 5000 0.5 -3.33513581E+03 43 0.00 15 0.01 58 0.00
NONMSQRT 4900 1.0 -1.15116171E+03 8 0.07 5 0.03 41 0.17
NONMSQRT 4900 0.5 -5.50172607E+02 8 0.00 9 0.03 20 0.08
OSCIGRAD 100000 1.0 -5.34690566E+08 8 0.31 5 0.23 43 0.13
OSCIGRAD 100000 0.5 -4.31017580E+08 8 0.00 9 0.20 21 0.09
PENALTY1 1000 1.0 -2.43960328E+13 1 0.07 2 0.07 1 0.00
PENALTY1 1000 0.5 -1.21985172E+13 1 0.00 4 0.05 0 0.00
POWELLSG 5000 1.0 -1.57803913E+04 2 0.01 3 0.01 7 0.00
POWELLSG 5000 0.5 -7.99992041E+03 2 0.00 6 0.01 3 0.00
POWER 10000 1.0 -1.13407283E+14 5 13.59 3 29.47 11 0.00
POWER 10000 0.5 -5.72243018E+13 5 0.00 6 29.24 5 0.00
QUARTC 5000 1.0 -1.33489191E+13 2 0.00 2 0.00 5 0.00
QUARTC 5000 0.5 -6.67448869E+12 2 0.00 4 0.00 2 0.00
SBRYBND 5000 1.0 -9.63032705E+11 20 0.03 16 0.03 179 0.06
SBRYBND 5000 0.5 -2.40760941E+11 20 0.00 21 0.02 89 0.03
SCHMVETT 5000 1.0 -7.28426352E+01 10 0.01 3 0.01 35 0.01
SCHMVETT 5000 0.5 -3.68824395E+01 10 0.00 6 0.01 17 0.01
SCOSINE 5000 1.0 -8.25379341E+10 13 0.03 11 0.02 187 0.02
SCOSINE 5000 0.5 -2.06344856E+10 13 0.00 13 0.01 93 0.01
SCURLY10 10000 1.0 -1.28077199E+30 4 0.04 3 0.03 11 0.00
SCURLY10 10000 0.5 -6.43068284E+29 4 0.00 6 0.03 5 0.00
SCURLY20 10000 1.0 -1.63614356E+31 4 0.07 3 0.04 11 0.00
SCURLY20 10000 0.5 -8.21427988E+30 4 0.00 6 0.04 5 0.00
SCURLY30 10000 1.0 -7.47542209E+31 4 0.09 3 0.06 9 0.01
SCURLY30 10000 0.5 -3.75274481E+31 4 0.00 6 0.06 4 0.00
SINQUAD 5000 1.0 -7.12672063E+03 2 0.00 4 0.00 5 0.00
SINQUAD 5000 0.5 -2.75259696E+03 2 0.00 8 0.00 2 0.00
SPARSINE 5000 1.0 -2.92330968E+06 4 0.48 3 1.16 11 0.00
SPARSINE 5000 0.5 -1.46942254E+06 4 0.00 6 1.18 5 0.00
SPARSQUR 10000 1.0 -1.19555581E+06 7 3.09 2 5.80 15 0.01
SPARSQUR 10000 0.5 -6.09140674E+05 7 0.00 4 5.86 7 0.00
SPMSRTLS 4999 1.0 -8.16740777E+01 7 0.02 3 0.01 21 0.01
SPMSRTLS 4999 0.5 -3.97487352E+01 7 0.00 6 0.01 10 0.00
SROSENBR 5000 1.0 -1.15690489E+02 9 0.01 4 0.01 5 0.00
SROSENBR 5000 0.5 -6.52884475E+01 9 0.00 8 0.01 2 0.00
SSBRYBND 5000 1.0 -5.98063290E+06 17 0.03 12 0.04 145 0.09
SSBRYBND 5000 0.5 -1.51010118E+06 17 0.00 15 0.02 72 0.04
SSCOSINE 5000 1.0 -5.13429696E+05 12 0.02 16 0.02 155 0.02
SSCOSINE 5000 0.5 -1.28378413E+05 12 0.00 19 0.01 77 0.01
TOINTGSS 5000 1.0 -4.23179011E+02 4 0.02 2 0.02 9 0.00
TOINTGSS 5000 0.5 -2.11839554E+02 4 0.00 4 0.01 4 0.00
TQUARTIC 5000 1.0 -2.91590249E-02 1 0.00 3 0.00 3 0.00
TQUARTIC 5000 0.5 -1.66434134E-02 1 0.00 6 0.00 1 0.00
TRIDIA 10000 1.0 -1.14762126E+06 7 0.02 3 0.02 15 0.00
TRIDIA 10000 0.5 -5.75687744E+05 7 0.00 6 0.02 7 0.00
VAREIGVL 5000 1.0 -1.03187418E+04 3 2.68 3 4.55 9 0.00
VAREIGVL 5000 0.5 -5.19645395E+03 3 0.00 6 4.58 4 0.00
WOODS 4000 1.0 -5.13132992E+05 2 0.01 3 0.01 7 0.00
WOODS 4000 0.5 -2.57913812E+05 2 0.00 6 0.01 3 0.00
YATP1LS 10200 1.0 -2.47600401E+06 1 4.29 3 11.48 3 0.01
YATP1LS 10200 0.5 -1.24008348E+06 1 0.00 5 7.65 1 0.00
YATP2LS 10200 1.0 -6.17000891E+04 1 8.21 3 11.89 3 0.00
YATP2LS 10200 0.5 -3.11097632E+04 1 0.00 5 7.77 1 0.00
Table C.1: Subproblem solves for CUTEst examples: Complete results (continued)

### v3

@@
-Extended-Krylov-subspace methods for trust-region and norm-regularization subproblems
+arXiv:2511.11135v3[math.NA] 02 Mar 2026
 Abstract
 We consider an effective new method for solving trust-region and norm-regularization problems that arise as subproblems in many optimization applications.
-We show that the solutions to such subproblems lie on a manifold of approximately very low rank as a function of their controlling parameters (trust-region radius or regularization weight).
-Based on this, we build a basis for this manifold using an efficient extended-Krylov-subspace iteration that involves a single matrix factorization.
+We show that the solutions to such subproblems as a function of their controlling parameters (trust-region radius or regularization weight).
+Based on this, we build a basis using an efficient extended-Krylov-subspace iteration that involves a single matrix factorization.
 The problems within the subspace using such a basis may be solved at very low cost using effective high-order root-finding methods.
@@
 (1.5) $(A+\sigma I)x(\sigma)=b\;\;\mbox{with}\;\;\|x(\sigma)\|=\Delta\;\;\mbox{for some}\;\;\sigma\geq\sigma_{\min}.$}
-We refer to the first possibility as the trivial—or interior—case, while the second is the general one; there is an additional probability-zero possibility, commonly called the hard case—this occurs if $b$ is orthogonal to all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)$ and the radius is too large—that we shall ignore for the time being.
+We refer to the first possibility as the trivial—or interior—case, while the second is the general one; there is an additional probability-zero possibility, commonly called the hard case—this occurs if $b$ is orthogonal to all the eigenvectors for the leftmost eigenvalue $\lambda_{1}(A)$ and the radius is too large—that we shall ignore for the time being [^1] A few methods are specifically designed to cope with this possibility [BeckVasi18:siopt, CarmDuch20:sirev]..
 If we can rule out the trivial and hard cases, (1.5) then suggests an iteration in which, for a sequence of positive $\sigma$, $x(\sigma)$ is found by solving the linear system (1.2), and its suitability assessed by examining the size of the residual $\|x(\sigma)\|-\Delta$.
-An interesting question is, therefore, what we can say about the manifold ${\cal S}(A,b)=\mbox{span}\{x(\sigma)\}$ as $\sigma$ varies between $\sigma_{\min}$ and $\infty$.
+An interesting question is, therefore, what we can say about the manifold ${\cal M}(A,b)=\{x\mid x=x(\sigma)\;\;\mbox{for some}\;\;\sigma\in[\sigma_{\min},\infty)\}$ or more specifically the ${\cal S}(A,b)=\mbox{span}\{x(\sigma)\}$, that contains ${\cal M}(A,b)$, as $\sigma$ varies between $\sigma_{\min}$ and $\infty$.
 There is no loss of generality, from a theoretical perspective, in considering a problem with a diagonal Hessian, since if $A$ has the spectral factorization
@@
 involving an orthogonal matrix $U$ whose columns are its eigenvectors, and a diagonal matrix $\Lambda$ whose entries are its eigenvalues, (1.1) is equivalent to
-(1.7) ${\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}n$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}\Lambda y-d^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
+(1.7) ${\scriptstyle\frac{1}{2}}y^{T}\Lambda y-d^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
 under the transformation
@@
 $(\Lambda+\sigma I)y(\sigma)=d$}
-for (1.7) are $y_{i}(\sigma)=1/(\lambda_{i}+\sigma)$.
-If we approximate the solution manifold ${\cal S}(\Lambda,d)$ by the discrete set $\{y(\sigma_{j}),j=1,\ldots,p\}$ at the evaluation points, and compute the singular value decomposition of the resulting $n$ by $p$ matrix $S$ with entries
-(1.8) $s_{i,j}=d_{i}/(\lambda_{i}+\sigma_{j})\;\;\mbox{for}\;\;i=1,\ldots,n\;\;\mbox{and}\;\;j=1,\ldots,p,$}
+for (1.7) are $1/(\lambda_{i}+\sigma)$.
+If we approximate the ${\cal S}(\Lambda,d)$ by the discrete set $\{y(\sigma_{j}),j=1,\ldots,p\}$ at the evaluation points, and compute the singular value decomposition of the resulting $n$ by $p$ matrix $S$ with entries
+(1.8) $(\lambda_{i}+\sigma_{j})\;\;\mbox{for}\;\;i=1,\ldots,n\;\;\mbox{and}\;\;j=1,\ldots,p,$}
 we find 37 significant singular values (those that are larger than $10^{-15}$ of the largest) in case (i), and for cases (ii)–(iv) there are 38, 36 and 35 respectively.
@@
 3.48e-03 1.61e-02 1.76e-10 1.34e-03 1.18e-02
-Table 1.1: Singular values of dominant manifold subspace for four eigenvalue distributions
-It is then remarkable that the manifold appears to be spanned by so few significant vectors, that is, it is approximately of very low rank.
+Table 1.1: singular values of for four eigenvalue distributions
+It is then remarkable that appears to be spanned by so few significant vectors, that is, it is approximately of very low rank.
 This turns out not to be a coincidence, and we shall exploit this throughout.
 This paper is organised as follows.
-In Section 2, we explain why the solution manifold of interest is (approximately) low rank.
-We follow this in Section 3 by considering methods that build approximations to this manifold.
+In Section 2, we explain why the of interest is (approximately) low We follow this in Section 3 by considering methods that build approximations to this .
 A particularly appealing approach may be constructed from a so-called extended Krylov subspace, and we show how in Section 4.
@@
 Thus $\sigma_{\min}^{+}<\lambda_{1}$, indeed $\sigma_{\min}^{+}+\lambda_{1}\geq\delta$, and the intervals $\sigma_{\max},-\sigma_{\min}^{+}]$ and $\lambda_{1},\lambda_{n}]$ are non-overlapping.
-To investigate the rank of $S$, clearly rows $i$ and $j$ of $S$ are linearly dependent if $\lambda_{i}=\lambda_{j}$, and thus $\mbox{rank}(S)=\mbox{rank}(\bar{S})$, where $\bar{S}$ comprises the rows of $S$ in which each set of such dependencies is replaced by a single representative.
 For any $m$ by $n$ matrix $X$ with decreasing singular values $\sigma_{1}()\geq\sigma_{2}(X)\geq\ldots\geq\sigma_{k}(X)$ for increasing $k$, the $\epsilon$-rank, $\mbox{rank}_{\epsilon}(X)$, of $X$ is the smallest integer $k$ for which $\sigma_{k+1}(X)\leq\epsilon\sigma_{1}(X)$.
@@
 For the examples illustrated in Table 1.1, this bound is 44, a very reasonable (upper-bound) approximation of the computed values.
-3 Manifold approximation from Krylov subspaces
+3 approximation from Krylov subspaces
 Now that we know that the solution we seek to (1.1) for any $\Delta$ lies in an approximately low-dimensional subspace, our task is to identify a suitable (for convenience, orthogonal) $n$ by $\ell$ basis matrix $V_{\ell}$ for this subspace.
 Once we do so, we may recover an approximate solution $x_{\ell}=V_{\ell}y_{\ell}$, where
-(3.1) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}P_{\ell}y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
+(3.1) $y_{\ell}={\scriptstyle\frac{1}{2}}y^{T}P_{\ell}y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta$}
 with $P_{\ell}=V_{\ell}^{T}AV_{\ell}$ and $b_{\ell}=V_{\ell}^{T}b$. The fact that we expect $\ell$ to be small provides an opportunity to use methods that would be inappropriate has it been larger.
@@
 This, however, does not avoid the cost of multiple factorizations; rival methods, e.g., [GoulRobiThor10:mpc, MoreSore83:sisc], also use factorizations, but experience suggests that relatively few often suffice.
-Another approach for constructing a good approximation of the manifold is to use a method based on a rational Krylov subspace [Ruhe84:LAA, AldaPali25:ARXIV].
+Another approach for constructing a good approximation of is to use a method based on a rational Krylov subspace [Ruhe84:LAA, AldaPali25:ARXIV].
 However, this will still require solving $\ell$ shifted linear systems.
-The alternative is to try to build a good approximation to the manifold by other means. We shall resort to methods based on the extended Krylov subspace (see Section 4).
+The alternative is to try to build a good approximation to by other means. We shall resort to methods based on the extended Krylov subspace (see Section 4).
 To justify this, first note that given a shift $\sigma$, (1.2) and (1.6) give that
@@
 where
-$y_{i}(\sigma)=\frac{d_{i}}{\lambda_{i}+\sigma},\;\;i=1,\ldots,n,\;\;\mbox{and}\;\;d=U^{T}b.$}
+$\frac{}{\lambda_{i}+\sigma},\;\;i=1,\ldots,n,\;\;\mbox{and}\;\;d=U^{T}b.$}
 Consider two cases and first suppose that Then using a Laurent series [Laur24:wiki], we have The smaller the ratio the faster the sequence reaches machine precision and few terms would be sufficient to approximate $y_{i}(\sigma)$.
@@
 Fortunately, much of the groundwork is already in place.
-We suppose, for the time being, that $A$ is positive definite.
-Then Jagels and Reichel [^2] They also propose a second, more expensive algorithm that copes with indefinite $A$.
-[JageReic09:laa] have shown how to build such a basis if the order [^3] We shall say why our order is more convenient later.
+We suppose, for this section and the next, that $A$ is positive definite333.
+Then Jagels and Reichel [^4] They also propose a second, more expensive algorithm that copes with indefinite $A$, [JageReic09:laa] have shown how to build such a basis if the order [^5] We shall say why our order is more convenient later.
 in which the components $A^{-m}b$ and $A^{m}b$ arise is interchanged, that is if instead
@@
 ${\cal K}_{2m+1}(A,b)=.$}
-Using the same logic as in [JageReic09:laa, Alg.2.1], that applies to ${\cal K}^{\prime}_{2m}(A,b)$, it is straightforward to derive the following algorithm, Algorithm 1 1, that instead uses the order ${\cal K}_{2m}(A,b)$.
+Using the same logic as in [JageReic09:laa, Alg.2.1], that applies to ${\cal K}^{\prime}_{2m}(A,b)$, it is straightforward to derive the following algorithm, Algorithm 1 1, that instead uses the order ${\cal K}_{2m}(A,b)$; for completeness, we provide details of an alternative justification, that avoids Laurent polynomials, in Appendix B.
 Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$ and subspace dimension bound $m\geq 1$.
@@
 $V_{2k}=(v_{0}:v_{-1}:v_{1}:\cdots:v_{k-1}:v_{-k})\;\;\mbox{and}\;\;V_{2k+1}=(V_{2k}:v_{k}),$
-that play the roles of $A_{k}$ in (3.1). It follows from (4.3) that
+that play the roles of in (3.1). It follows from (4.3) that
 (4.5) $Av_{-k}=\alpha_{k-1}v_{k-1}+\alpha_{-k}v_{-k}+\delta_{k}v_{k}\;\;\mbox{for $k=1,\ldots,m$}\;\;$
@@
 and
-$p_{i,2k}=v_{i}^{T}Av_{-k}=0\;\;\mbox{for $i\neq 2k-1,-2k$ and $2k+1$.}\;\;$
+$p_{i,2k}=v_{i}^{T}Av_{-k}=0\;\;\mbox{for $i\neq 2k-1,$ and $2k+1$.}\;\;$
 By symmetry of $P_{2m+1}$, we also have
@@
 (4.8) $\begin{array}[]{rl}\beta_{k}Av_{k}\hskip-5.69054pt&=v_{k}-\beta_{-k}Av_{-k}-\delta_{-k-1}Av_{-k-1}\\ &=-\beta_{-k}\alpha_{k-1}v_{k-1}-\beta_{-k}\alpha_{-k}v_{-k}+(1-\beta_{-k}\delta_{k}-\delta_{-k-1}\alpha_{k})v_{k}\\ &\hskip 5.69054pt\hskip 8.53581pt-\delta_{-k-1}\alpha_{-k-1}v_{-k-1}-\delta_{-k-1}\delta_{k+1}v_{k+1}\end{array}$}
-using (4.5).
-Again, the orthogonality of the $v_{i}$ and symmetry implies that [^4] When $\delta_{-k-1}=0$, the last three terms in (4.9) are $p_{2k+1,2k+1}=(1-\beta_{-k}\delta_{k})/\beta_{k}$ and $p_{2k+2,2k+1}=p_{2k+3,2k+1}=0$, and thus do not require that we calculate $\alpha_{k}$, $\alpha_{-k-1}$ and $\delta_{k+1}$.
+using (4.5). Again, the orthogonality of the $v_{i}$ and symmetry implies that
 (4.9) $\begin{array}[]{rl}p_{2k-1,2k+1}\hskip-5.69054pt&=p_{2k+1,2k-1}=v_{k-1}^{T}Av_{k}=-\displaystyle\frac{\beta_{-k}\alpha_{k-1}}{\beta_{k}},\\ p_{2k,2k+1}\hskip-5.69054pt&=p_{2k+1,2k}=v_{-k}^{T}Av_{k}=-\displaystyle\frac{\beta_{-k}\alpha_{-k}}{\beta_{k}},\\ p_{2k+1,2k+1}\hskip-5.69054pt&=v_{k}^{T}Av_{k}=\displaystyle\frac{1-\beta_{-k}\delta_{k}-\delta_{-k-1}\alpha_{k}}{\beta_{k}},\\ p_{2k+2,2k+1}\hskip-5.69054pt&=p_{2k+1,2k+2}=v_{-k-1}^{T}Av_{k}=-\displaystyle\frac{\delta_{-k-1}\alpha_{-k-1}}{\beta_{k}},\\ p_{2k+3,2k+1}\hskip-5.69054pt&=p_{2k+1,2k+3}=v_{k+1}^{T}Av_{k}=-\displaystyle\frac{\delta_{-k-1}\delta_{k+1}}{\beta_{k}}\end{array}$}
 and
-$p_{i,2k+1}=p_{2k+1,i}=v_{i}^{T}Av_{k}=0\;\;\mbox{for $i\neq-k-1,k-1,k,-k$ and $k+1$.}\;\;$
+$p_{i,2k+1}=p_{2k+1,i}=v_{i}^{T}Av_{k}=0\;\;\mbox{for $i\neq-k-1,k-1,k,-k$ and $k+1$,}\;\;$
 Thus, it follows from from (4.8) and (4.9) that
@@
 $\frac{\beta_{-k}\alpha_{k-1}}{\beta_{k}}=-\frac{\delta_{-k}\delta_{k}}{\beta_{k-1}}$
-by comparing $p_{2k-1,2k+1}$ with $p_{2k-3,2k-1}$ for subsequent $k$.
-Finally, if follows from (4.2), (4.5) with $k=1$ and the orthogonality of the $v_{i}$, that [^5] When $\delta_{-1}=0$, $p_{1,1}=1/\beta_{0}$ and $p_{1,2}=p_{1,3}=0$, and thus $\alpha_{0}$, $\alpha_{-1}$ and $\delta_{1}$ need not be calculated.
+by comparing $p_{2k-1,2k+1}$ with $p_{2k-3,2k-1}$ for subsequent $k$. Finally, if follows from (4.2), (4.5) with $k=1$ and the orthogonality of the $v_{i}$, that
 (4.11) $p_{1,1}=v_{0}^{T}Av_{0}=\frac{1-\delta_{-1}\alpha_{0}}{\beta_{0}},\;\;p_{2,1}=v_{-1}^{T}Av_{0}=-\frac{\delta_{-1}\alpha_{-1}}{\beta_{0}}\equiv\alpha_{0},\;\;p_{3,1}=v_{1}^{T}Av_{0}=-\frac{\delta_{-1}\delta_{1}}{\beta_{0}},$}
@@
 (5.5) $\|r_{2k}\|=|p_{2k+1,2k-1}y_{2k,2k-1}+p_{2k+1,2k}y_{2k,2k}|.$}
-Note that in the exceptional case where $\delta_{i}=0$, the corresponding $\|r_{i}\|=0$, and thus termination will always occur at this step.
+Note that in the exceptional case where $\delta_{\ell}=0$, the corresponding $\|r_{\ell}\|=0$, and thus termination will always occur at this step;
 The TREK algorithm embeds solutions of (3.1) after every computation of $v_{-k}$ and $v_{k}$ in ; at these stages the matrices $V_{2k}$ and $V_{2k+1}$, respectively, are complete, and we have the data required to solve the appropriate subproblems (3.1).
@@
 The need to resolve a problem with the same $A$ and $b$, but a smaller $\Delta$ is a common occurrence in trust-region methods, since such a mechanism is used to guarantee convergence when the current solution has proved unsatisfactory [ConnGoulToin00].
-This is easy to accommodate within Algorithm 2; the data generated by the previous call to the algorithm is retained, and the ‘”for loop” re-entered with $k$ taking the value $k_{\mbox{\scriptsize stop}}$ that it last had.
-The motivation is simply that although the segment of the manifold ${\cal S}(A,b)$ we have observed may now have to expand slightly, it may well be that there is little actual difference; the data provided in Appendix C indicates that in only two of ninety six cases examined were any further extended Krylov iterations necessary, and thus that resolves are generally very inexpensive.
+This is easy to accommodate within Algorithm 2; the data generated by the previous call to the algorithm is retained, and the “for loop” re-entered with $k$ taking the value $k_{\mbox{\scriptsize stop}}$ that it last had.
+The motivation is simply that although the to expand slightly, it may well be that there is little actual difference; the data provided in Appendix D indicates that in only cases examined were any further extended Krylov iterations necessary, and thus that resolves are generally very inexpensive.
 Examining the bound (A.6) given in Proposition A.3 in Appendix A, and looking at Figure A.1, we see that the convergence rate of the extended Krylov subspace improves for $\sigma$ beyond a certain point, which is consistent with this observation.
@@
 To cope with this, it is sometimes more effective to pose the problem as
-(5.6) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta,$}
-where $P$ is an easily-invertible matrix—a diagonal $P$ is often a simple but effective choice. On writing
-(5.7) $x_{\mbox{\tiny P}}=Px,$}
-the minimizer of (5.6) is formally $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$, where
-(5.8) $x_{\mbox{\tiny P}}{}_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x_{\mbox{\tiny P}}{}^{T}A_{\mbox{\tiny P}}x_{\mbox{\tiny P}}-b_{\mbox{\tiny P}}{}^{T}x_{\mbox{\tiny P}}\;\;\mbox{subject to}\;\;\|x_{\mbox{\tiny P}}\|\leq\Delta$}
-and $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and $b_{\mbox{\tiny P}}=P^{-T}b$.
-Since the TREK algorithm described above merely relies on products with $A_{\mbox{\tiny P}}$ and its inverse $A_{\mbox{\tiny P}}{}^{-1}=PA^{-1}P^{T}$, once again we merely need a factorization of $A$ (to find $A^{-1}v$ for given $v$) so long as means of inverting $P$ and its transpose are available.
-Commonly a positive-definite scaling matrix $S$ that is (in some sense) close to $A$ is known, and in this case a decomposition, for example a sparse Cholesky factorization, $S=P^{T}P$ gives the appropriate $P$.
+(5.6) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|\leq\Delta,$}
+where is an easily-invertible matrix—a diagonal is often a simple but effective choice. On writing
+(5.7) $x_{\mbox{\tiny}}=x,$}
+the minimizer of (5.6) is formally $x_{*}=^{-1}x_{\mbox{\tiny}}{}_{*}$, where
+(5.8) $x_{\mbox{\tiny}}{}_{*}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x_{\mbox{\tiny}}{}^{T}A_{\mbox{\tiny}}x_{\mbox{\tiny}}-b_{\mbox{\tiny}}{}^{T}x_{\mbox{\tiny}}\;\;\mbox{subject to}\;\;\|x_{\mbox{\tiny}}\|\leq\Delta$}
+and $A_{\mbox{\tiny}}=^{-T}A^{-1}$ and $b_{\mbox{\tiny}}=^{-T}b$.
+Since the TREK algorithm described above merely relies on products with $A_{\mbox{\tiny}}$ and its inverse $A_{\mbox{\tiny}}{}^{-1}=A^{-1}{}^{T}$, once again we merely need a factorization of $A$ (to find $A^{-1}v$ for given $v$) so long as means of inverting and its transpose are available.
+Commonly a positive-definite scaling matrix $S$ that is (in some sense) close to $A$ is known, and in this case a decomposition, for example a sparse Cholesky factorization, $S=^{T}$ gives the appropriate .
 We provide details of the necessary changes to devise suitable algorithms to solve the resulting problem
 (5.9) ${\begin{array}[t]{c}\mbox{minimize}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|_{S}\leq\Delta,$}
-where $\|x\|_{S}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{x^{T}Sx}$, in Appendix B.
+where $\|x\|_{S}\stackrel{{\scriptstyle\rm def}}{{=}}\sqrt{x^{T}Sx}$, in Appendix C.
 We illustrate one such algorithm here as Algorithm 3 3.
@@
 Algorithm 3 The TREK algorithm to solve the trust-region subproblem (5.9)
-It is also worth noting that the extended Krylov subspace (4.1) generated by $A_{\mbox{\tiny P}}$ and $b_{\mbox{\tiny P}}$,
-(5.10) ${\cal K}_{2m+1}(A_{\mbox{\tiny P}},b_{\mbox{\tiny P}})=\text{span}\{b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{-1}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}b_{\mbox{\tiny P}},\ldots,A_{\mbox{\tiny P}}^{-m}b_{\mbox{\tiny P}},A_{\mbox{\tiny P}}^{m}b_{\mbox{\tiny P}}\},$}
+It is also worth noting that the extended Krylov subspace (4.1) generated by $A_{\mbox{\tiny}}$ and $b_{\mbox{\tiny}}$,
+(5.10) ${\cal K}_{2m+1}(A_{\mbox{\tiny}},b_{\mbox{\tiny}})=\text{span}\{b_{\mbox{\tiny}},A_{\mbox{\tiny}}^{-1}b_{\mbox{\tiny}},A_{\mbox{\tiny}}b_{\mbox{\tiny}},\ldots,A_{\mbox{\tiny}}^{-m}b_{\mbox{\tiny}},A_{\mbox{\tiny}}^{m}b_{\mbox{\tiny}}\},$}
 is
-$P^{-T}\text{span}\{b,(SA^{-1})b,(AS^{-1})b,\ldots,(SA^{-m})b,(AS^{-1})^{m}b\},$}
-and it is then possible to derive a suitable variant of Algorithm 1 using $S$ and its inverse rather than $P$ and $P^{-1}$.
+$^{-T}\text{span}\{b,(SA^{-1})b,(AS^{-1})b,\ldots,b,(AS^{-1})^{m}b\},$}
+and it is then possible to derive a suitable variant of Algorithm 1 using $S$ and its inverse rather than and -1.
 6 Solving the general trust-region subproblem
@@
 Given $\sigma_{\mbox{\tiny S}}$, it is trivial to adapt any bracketing root-finding procedure to account for this extended interval.
-We have in mind that the shift should chosen be so that $A_{\sigma_{\mbox{\tiny S}}}$ is positive definite.
+We have in mind that the shift should so that $A_{\sigma_{\mbox{\tiny S}}}$ is positive definite.
 The smallest shift for which $A_{\sigma_{\mbox{\tiny S}}}$ is guaranteed to be positive semi-definite is $\sigma_{\min}$ from (1.3) so any value slightly larger than this is possible.
 But a far cheaper option that does not rely on knowing the spectrum of $A$, once $A$ has been found not to be positive definite, is to use the value
-$\sigma_{\mbox{\tiny G}}=-\min_{1\leq i\leq n}\left[a_{ii}-\sum_{1\leq j\neq i\leq n}|a_{i,j}|\right]+\epsilon_{\mbox{\tiny G}}$}
-from the Gershgorin theorem, requiring a single pass through the values of $A$; we have found that this with the heuristic value $\epsilon_{\mbox{\tiny G}}=\sqrt{\epsilon_{\mbox{\tiny M}}}\max_{1\leq i,j\leq n}|a_{i,j}|$ where $\epsilon_{\mbox{\tiny M}}$ is the machine precision is effective.
+$\sigma_{\mbox{\tiny G}}=-\min_{1\leq i\leq n}\left[-\sum_{1\leq j\neq i\leq n}||\right]+\epsilon_{\mbox{\tiny G}}$}
+from the Gershgorin theorem, requiring a single pass through the values of $A$; we have found that this with the heuristic value $\epsilon_{\mbox{\tiny G}}=\sqrt{\epsilon_{\mbox{\tiny M}}}\max_{1\leq i,j\leq n}|$ where $\epsilon_{\mbox{\tiny M}}$ is the machine precision is effective.
 As before, we generate the extended Krylov subspace ${\cal K}_{\ell}(A_{\sigma_{\mbox{\tiny S}}},b)$ using Algorithm 1.
@@
 since the columns of are orthonormal. Thus subspace trust-region subproblems of the form (3.1) can be formulated as
-(6.2) $y_{\ell}={\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle y\in\mbox{\raisebox{-0.6pt}{\scriptsize I\thinspace R\thinspace}}^{k}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}y^{T}(P_{\ell}-\sigma_{\mbox{\tiny S}}I)y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
+(6.2) $y_{\ell}={\scriptstyle\frac{1}{2}}y^{T}(P_{\ell}-\sigma_{\mbox{\tiny S}}I)y-b_{\ell}^{T}y\;\;\mbox{subject to}\;\;\|y\|\leq\Delta.$}
 More particularly, if we rely on the spectral factorization (1.6), the diagonal subproblem (5.2) becomes
@@
 6.1 Other elliptical norms
-For the general problem defined in terms of an elliptical norm $\|x\|_{S}=\sqrt{x^{T}Sx}$ for a given symmetric positive definite $S=P^{T}P$, a similar approach is possible.
-Transforming via (5.7) as in Subsection 5.2, the (non-interior) optimality condition (1.2) becomes
-(6.3) $(A_{\mbox{\tiny P}}+\sigma I)x_{\mbox{\tiny P}}(\sigma)=b_{\mbox{\tiny P}}\;\;\mbox{with}\;\;\|x_{\mbox{\tiny P}}(\sigma)\|=\Delta$}
+For the general problem defined in terms of an elliptical norm $\|x\|_{S}=\sqrt{x^{T}Sx}$ for a given symmetric positive definite $S=^{T}$, a similar approach is possible.
+Transforming via (5.7) as in Subsection 5.2, the (non-interior) optimality condition becomes
+(6.3) $(A_{\mbox{\tiny}}+\sigma I)x_{\mbox{\tiny}}(\sigma)=b_{\mbox{\tiny}}\;\;\mbox{with}\;\;\|x_{\mbox{\tiny}}(\sigma)\|=\Delta$}
 for some
-$\sigma\geq\sigma_{\mbox{\tiny P}}{}_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A_{\mbox{\tiny P}})),$}
+$\sigma\geq\sigma_{\mbox{\tiny}}{}_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A_{\mbox{\tiny}})),$}
 or equivalently
 $(A+\sigma S)x(\sigma)=b\;\;\mbox{with}\;\;\|x(\sigma)\|_{S}=\Delta;$}
-note that $\lambda_{1}(A_{\mbox{\tiny P}})$ is equivalently the leftmost eigenvalue $\lambda_{1}(A,S)$ of the symmetric matrix pencil $A+\lambda S$. The equivalent version of (6.1) is then
+note that $\lambda_{1}(A_{\mbox{\tiny}})$ is equivalently the leftmost eigenvalue $\lambda_{1}(A,S)$ of the symmetric matrix pencil [GoluvanL96, §7,7].
+$A-\lambda S$.
+The equivalent version of (6.1) is then
 (6.4) $(A_{\sigma_{\mbox{\tiny S}}}+\theta S)x_{\sigma_{\mbox{\tiny S}}}(\theta)=b\;\;\mbox{with}\;\;\|x_{\sigma_{\mbox{\tiny S}}}(\theta)\|_{S}=\Delta,$}
 where
-$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\mbox{\tiny P}}{}_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A,S)-\sigma_{\mbox{\tiny S}}).$}
+$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\mbox{\tiny}}{}_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A,S)-\sigma_{\mbox{\tiny S}}).$}
 Although it is less obvious how to find a lower bound on $\lambda_{1}(A,S)$ in general, it is possible to use a variant of the inexpensive Gershgorin bound in the commonly-occurring case where $S$ is strictly diagonally dominant (see [GoulRobiThor10:mpc, §3.3.2] for details).
@@
 We now provide some evidence that our new approach is a useful alternative to common existing methods.
-We consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables.
+We consider the set of unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables.
 For these, we generate the gradient and Hessian at the provided initial point, and use these for b-b and $A$ respectively.
 We compare our new method (Algorithm 2 implemented as TREK), using the initial (Gershgorin-based) shifting strategy outlined in Section 6, with the multi-factorization method [GoulRobiThor10:mpc] available as TRS within GALAHAD [GoulOrbaToin03:toms] and the iterative method GLTR [GoulLuciRomaToin99:siopt] from the same library; GLTR only requires Hessian-vector products which can be an asset in some cases.
-Since all three methods can take advantage of information gathered when solving a problem with one radius and are now faced with another with identical data but a smaller radius, we consider each problem with radii, $\Delta=1$ and then $0.5$.
-TREK uses an iteration bound $m=100$, but almost always [^8] In four instances, $m$ was too small, but increasing the value to $400$ cured the problem.
+Since all three methods can take advantage of information gathered when solving a problem with one radius and are now faced with another with identical data but a smaller radius, we consider each problem with TREK uses an iteration bound , but almost always [^8] In instances, $m$ was too small, but increasing the value to $400$ cured the problem.
 terminates with far-fewer $k$.
@@
 On some examples, particularly those whose Hessians are banded, further experiments not listed here revealed that LAPACK’s PBTR band solver is a good alternative.
-We summarize our findings in Figure 7.1 using the performance profile [dolamore02:mp] in which relative times are ranked. The profile is built using the data given in Appendix C.
+We summarize our findings in Figure 7.1 using the performance profile [dolamore02:mp] in which relative times are ranked. The profile is built using the data given in Appendix D.
 Figure 7.1: Performance profile comparing three trust-region subproblem solvers on larger CUTEst examples.
@@
 But we were encouraged to see that TREK performs well in other cases, and the requirement of a single factorization often proves decisive when compared to the multi-factorization approach used by TRS.
-[^9] TRS also failed in three cases because the shifted matrix $A+\sigma I$ became too close to singularity to progress.
+[^9] TRS also failed in cases because the shifted matrix $A+\sigma I$ became too close to singularity to progress.
 8 Regularization subproblems
@@
 (8.2) $\sigma=\rho\|x(\sigma)\|^{r-2}$}
-for some $\sigma\geq\sigma_{\min}$ [CartGoulToin11:mp, NestPoly06:mp].
-Thus the manifold of solutions is identical to that for the trust-region subproblem, but the one of interest satisfies (8.2) instead of (1.5); there is no interior-solution possibility in this case.
+for some $\sigma\geq\sigma_{\min}$ Thus the of solutions is identical to that for the trust-region subproblem, but the one of interest satisfies (8.2) instead of (1.5); there is no interior-solution possibility in this case.
 Since this is the case, it is trivial to recast the TREK algorithm for (8.1), the only differences are that the steps that calculate $y_{2k}$ and $y_{2k+1}$ and their shifts find instead
@@
 This lies somewhere between multiple factorization methods that were first popularised by Moré and Sorensen [MoreSore83:sisc] and CG/Lanczos-based factorization-free methods such as GLTR [GoulLuciRomaToin99:siopt].
-The novel features are firstly that it is shown that the manifold of all possible solutions lies on an (approximately) very low-rank manifold, and secondly that by extending the Krylov subspace to include products with $A^{-1}$ as well as $A$, this manifold may be computed very efficiently.
+The novel features are firstly that it is shown that the of all possible solutions lies in an (approximately) very , and secondly that by extending the Krylov subspace to include products with $A^{-1}$ as well as $A$, this may be computed very efficiently.
 This then adds a further tool to the arsenal of solvers that trust-region and cubic-regularization methods for general optimization methods rely on.
 Unsurprisingly, the new method is not always the best, but in some cases it is.
-It is also remarkable that empirically that the extended Krylov-subspace approach requires so few iterations to build a good approximation of the solution manifold; examination of the results in Appendix C illustrates only a few instances where more than 20 iterations are required.
+It is also remarkable that empirically that the extended Krylov-subspace approach requires so few iterations to good approximation of the solution ; examination of the results in Appendix D illustrates only a few instances where more than 20 iterations are required.
 We provide analysis to explain why this should be so, although this may well be tightened in future.
-Software that implements the algorithms will be available very shortly [^10] The TREK algorithm with its documentation is now available as a beta release, and once C, Python, Julia and Matlab interfaces have been finished, it will be simple to extend these to the norm regularisation case discussed in Section 8 as the NREK package.
-as part of the GALAHAD [GoulOrbaToin03:toms] fortran library, and there are interfaces to other popular languages such as C, Python and Julia, and tools such as Matlab.
+Software that implements the algorithms is now available 101010 as part of the GALAHAD [GoulOrbaToin03:toms] fortran library, and there are interfaces to other popular languages such as C, Python and Julia, and tools such as Matlab.
 The codes are designed to solve not only a single instance, but a sequence in which the regularization is tightened, without recourse to expensive re-evaluations; the current subspace $V_{k}$ is extended if it is insufficient as the regularization changes.
 The new packages will be rolled out as optional subproblem solvers in other GALAHAD packages in due course.
+Acknowledgement
 Appendix A Convergence analysis
 The first $j$-th matrix products of the EKS algorithm (Algorithm 1) generate an orthogonal basis for the extended Krylov subspace ${\cal K}_{j}(A,b)$.
-A natural question is then, how closely do the elements of this subspace lie to components of the manifold $x(\sigma)$.
+A natural question is then, how closely do the elements of lie to .
 That is to say, for a given $\sigma$, is there an entry in ${\cal K}_{j}(A,b)$ close to $x(\sigma)$, and if so, how close?
 Consider the symmetric positive-definite matrix $A$ with smallest and largest eigenvalues $\lambda_{1}=\lambda_{1}(A)$ and $\lambda_{n}=\lambda_{n}(A)$, and consider the interval ${\cal W}=[\lambda_{1},\lambda_{n}]\subset\mathbb{R_{+}^{*}}$.
-In this section, we repeat for completeness, with minor variations, the theory developed by Knizhnerman and Simoncini [KnizSimo11:NUMMAT] to prove that the convergence of the extended-Krylov-subspace method to build an approximate basis for the solution manifold $\{(A+\sigma I)^{-1}b\}$, for all $\sigma\notin-{\cal W}-(-\delta_{1},\delta_{n})=-(\lambda_{1}-\delta_{1},\lambda_{n}+\delta_{n})$, and $\delta_{1},\delta_{n}>0$, is linear at worst.
+In this section, we repeat for completeness, with minor variations, the theory developed by Knizhnerman and Simoncini [KnizSimo11:NUMMAT] to prove that the convergence of the extended-Krylov-subspace method to build an approximate basis for the ${\cal S}(A,b)$, for all $\sigma\notin-{\cal W}-(-\delta_{1},\delta_{n})=-(\lambda_{1}-\delta_{1},\lambda_{n}+\delta_{n})$, and $\delta_{1},\delta_{n}>0$, is linear at worst.
 The proof relies heavily on 19th to mid 20th century complex analysis, and we refer the reader to [KnizSimo11:NUMMAT] for the details; the technical buildup may appear somewhat fearsome to the uninitiated.
@@
 $\frac{\Psi^{\prime}(w)}{\Psi(w)-u}=\frac{1}{w}\sum_{k=0}^{\infty}\overline{\phi_{k}\left(\frac{1}{\bar{w}}\right)}M_{k}(u).$
-The special (matrix-valued) case for which $u=A$ and $\Psi(w)=zI$ leads directly to (A.5).
+The special (matrix-valued) case for which $u=A$ and $\Psi(w)=zI$, leads directly to (A.5).
 Proposition A.3.
@@
 Theorem A.5.
-The extended Krylov method to approximate a solution manifold
+The extended Krylov method to approximate a
 $\{x(\sigma)\}_{\sigma\notin(-\lambda_{n}-\delta_{n},-\lambda_{1}+\delta_{1})}$ converges at worst linearly with a convergence rate
@@
 The proof follows straightforwardly from Proposition A.3, and the discussion of the evolution of $E(\sigma)$ given there.
-Appendix B Modified norm trust region subproblem solver
-Here we provide a detailed description of the modifications to the TREK algorithm (Algorithm 2) that are necessary when using the trust region norm for which $\|x\|_{S}^{2}=x^{T}Sx$, with $S=P^{T}P$, as discussed in Section 5.2.
-As we stated there, it suffices to apply Algorithm 2 with the matrix $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and vector $b_{\mbox{\tiny P}}=P^{-T}b$, and to recover the solution via $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$.
-Since $S^{-1}=P^{-1}P^{-T}$, this is formally as Algorithm 4 4.
+Appendix B Computing an orthogonal basis of ${\cal K}_{\ell}(A,b)$
+Given the $n$ by $n$ symmetric matrix $A$, the $n$-vector $b$, we define a sequences of extended Krylov subspaces
+${\cal K}_{2k}(A,b)={\rm span}\{A^{-k}b,\ldots,A^{k-1}b\}$}
+and
+${\cal K}_{2k+1}(A,b)={\rm span}\{A^{-k}b,\ldots,A^{k}b\}={\rm span}\{{\cal K}_{2k}(A,b),A^{k}b\}$}
+where ${\cal K}_{0}(A,b)={\rm span}\{b\}$, and thus
+${\cal K}_{2k}(A,b)={\rm span}\{{\cal K}_{2k-1}(A,b),A^{-k}b\}.$}
+We build a sequence of orthogonal basis for these so that
+${\cal K}_{2k}(A,b)={\rm span}\{v_{-k},\ldots,v_{k-1}\}\;\;\mbox{and}\;\;{\cal K}_{2k+1}(A,b)={\rm span}\{v_{-k},\ldots,v_{k}\}$}
+with ${\cal K}_{0}(A,b)=\{b/\|b\|\}$. Note that
+$\begin{array}[]{rl}{\cal K}_{2k-1}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-1)}b,\ldots,A^{k-1}b\}={\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\},\\ {\cal K}_{2k-2}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-1)}b,\ldots,A^{k-2}b\}={\rm span}\{v_{-(k-1)},\ldots,v_{k-2}\}\\ \;\;\mbox{and}\;\;{\cal K}_{2k-3}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-2)}b,\ldots,A^{k-2}b\}={\rm span}\{v_{-(k-2)},\ldots,v_{k-2}\}\end{array}$}
+We find $v_{j}$ by orthogonalizing $Av_{-j}$ with respect to columns of $V_{2j}$, in which case
+(B.11) $Av_{-j}\in{\cal K}_{2j+1},$}
+and $v_{-(j+1)}$ by orthogonalizing $A^{-1}v_{j}$ with respect to the columns of $V_{2j+1}$, so that
+(B.12) $A^{-1}v_{j}\in V_{2j+2}.$}
+By construction, it then follows that
+(B.13) $\delta_{k}v_{k}=Av_{-k}-\sum_{j=-k}^{k-1}\alpha_{j}v_{j}=Av_{-k}-\alpha_{-k}v_{-k}-\alpha_{k-1}v_{k-1}-\!\!\!\sum_{j=-k+1}^{k-2}\alpha_{j}v_{j}$}
+for some normalizing constant $\delta_{k}$, which implies that
+$0=\delta_{k}v_{i}^{T}v_{-k}=v_{j}^{T}Av_{-k}-\alpha_{j},$}
+i.e.
+(B.14) $\alpha_{j}=v_{j}^{T}Av_{-k}=v_{-k}^{T}Av_{j}$}
+for $j=-k,\ldots,k-1$. Consider first $j=-k+1,\ldots,-1$, and for these
+$\alpha_{j}\equiv\alpha_{-i}=v_{-i}^{T}Av_{-k}=v_{-k}^{T}Av_{-i}$}
+for $i=1,\ldots,k-1$. But then, it follows from (B.11) that
+$Av_{-i}\in{\cal K}_{2i+1}(A,b)\subseteq{\cal K}_{2k-1}(A,b),$}
+and thus $\alpha_{-i}=0$ since $v_{-k}$ is orthogonal to ${\cal K}_{2k-1}(A,b)={\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\}$.
+By contrast, consider $j=0,\ldots,k-2$, and note that in this case $v_{j}\in{\cal K}_{2k-3}(A,b)$.
+But then
+$Av_{j}\in{\rm span}\{A^{-(k-3)}b,\ldots,A^{k-1}b\}\subset{\rm span}\{A^{-(k-1)}b,\ldots,A^{k-1}b\}={\cal K}_{2k-1}(A,b)$}
+and thus as before $\alpha_{j}=0$ since $v_{-k}$ is orthogonal to ${\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\}$. Thus (B.13) simplifies to a short-term recurrence
+(B.15) $\delta_{k}v_{k}=Av_{-k}-\alpha_{-k}v_{-k}-\alpha_{k-1}v_{k-1}.$}
+Similarly, by construction, we have that
+(B.16) $\delta_{-(k+1)}v_{-(k+1)}=A^{-1}v_{k}-\sum_{j=-k}^{k}\beta_{j}v_{j}=A^{-1}v_{k}-\beta_{-k}v_{-k}-\beta_{k}v_{k}-\!\!\!\sum_{j=-k+1}^{k-1}\beta_{j}v_{j}$}
+for another normalizing constant $\delta_{-(k+1)}$, which implies that
+$0=\delta_{-(k+1)}v_{i}^{T}v_{-(k+1)}=v_{j}^{T}A^{-1}v_{-k}-\beta_{j},$}
+i.e.
+(B.17) $\beta_{j}=v_{j}^{T}A^{-1}v_{k}=v_{k}^{T}A^{-1}v_{j}$}
+for $j=-k,\ldots,k$. Consider first $j=0,\ldots,k-1$, and for these (B.12) implies that
+$A^{-1}v_{j}\in{\cal K}_{2i+2}(A,b)\subseteq{\cal K}_{2k}(A,b).$}
+But then $\beta_{j}=0$ as $v_{k}$ does not lie in ${\rm span}\{v_{-k},\ldots,v_{k-1}\}$. Whereas, if $j=-k+1,\ldots,-1$,
+$\beta_{j}\equiv\beta_{-i}=v_{-i}^{T}A^{-1}v_{k}=v_{k}^{T}A^{-1}v_{-i}$}
+for $i=1,\ldots,k-1$. Note that in this case $v_{-i}\in{\cal K}_{2i}(A,b)\subseteq{\cal K}_{2k-2}(A,b)$, and then it follows that
+$A^{-1}v_{-i}\in{\rm span}\{A^{-k}b,\ldots,A^{k-3}b\}\subset{\rm span}\{A^{-k}b,\ldots,A^{k-1}b\}={\cal K}_{2k}(A,b)$}
+Hence, once again, $\beta_{j}=0$ as $v_{k}$ is orthogonal to ${\rm span}\{v_{-k},\ldots,v_{k-1}\}$. Thus, as before, (B.16) reduces to a short-term recurrence
+(B.18) $\delta_{-(k+1)}v_{-(k+1)}=A^{-1}v_{k}-\beta_{-k}v_{-k}-\beta_{k}v_{k}.$}
+The identities (B.14), (B.15), (B.17) and (B.18) then lead to Algorithm 1.
+Appendix C Modified norm trust region subproblem solver
+Here we provide a detailed description of the modifications to the TREK algorithm (Algorithm 2) that are necessary when using the trust region norm for which $\|x\|_{S}^{2}=x^{T}Sx$, with $S=^{T}$, as discussed in Section 5.2.
+As we stated there, it suffices to apply Algorithm 2 with the matrix $A_{\mbox{\tiny}}=^{-T}A^{-1}$ and vector $b_{\mbox{\tiny}}=^{-T}b$, and to recover the solution via $x_{*}=^{-1}x_{\mbox{\tiny}}{}_{*}$.
+Since $S^{-1}=^{-1}{}^{-T}$, this is formal as Algorithm 4 4.
 Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
-Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta.$
+Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|x\|\leq\Delta.$
 2mm Algorithm:
-$x:=A^{-1}b$, $w:=Px$;
+$x:=A^{-1}b$, $w:=x$;
 if $\|w\|\leq\Delta$ then
  exit with the interior solution $x_{*}=x$ and shift $\sigma_{*}=0$;
-$\delta_{0}:=\|P^{-T}b\|$, $v_{0}:=P^{-T}b/\delta_{0}$, $u:=w/\delta_{0}$;
+$\delta_{0}:=\|^{-T}b\|$, $v_{0}:=^{-T}b/\delta_{0}$, $u:=w/\delta_{0}$;
 $\beta_{0}:=u^{T}v_{0}$;
@@
     $v_{-k}:=u/\delta_{-k}$;
-    $u:=P^{-T}AP^{-1}v_{-k}$;
+    $u:=^{-T}A^{-1}v_{-k}$;
     $\alpha_{k-1}:=u^{T}v_{k-1}$, $u\leftarrow u-\alpha_{k-1}v_{k-1}$;
@@
  if $\delta_{k}>0$ then
-    $v_{k}:=u/\delta_{k}$, $u:=PA^{-1}P^{T}v_{k}$;
+    $v_{k}:=u/\delta_{k}$, $u:=A^{-1}{}^{T}v_{k}$;
     $\beta_{-k}:=u^{T}v_{-k}$, $u\leftarrow u-\beta_{-k}v_{-k}$;
@@
 Algorithm 4 The TREK algorithm to solve the trust-region subproblem (5.6)
-Under the transformation $w_{k}=P^{-1}v_{k}$ i.e., $v_{k}=Pw_{k}$, and $u=Pv$, i.e., $v=P^{-1}u$ (and then replacing $v_{k}\rightarrow w_{k}$ and $u\rightarrow v$ so that the resulting changes may be easily compared with Algorithm 2), and introducing temporary workspace vectors $z$ and $s$, we may rewrite Algorithm 4 as Algorithm 5 5.
+Under the transformation $w_{k}=^{-1}v_{k}$ i.e., $v_{k}=w_{k}$, and $u=v$, i.e., $v=^{-1}u$ (and then replacing $v_{k}\rightarrow w_{k}$ and $u\rightarrow v$ so that the resulting changes may be easily compared with Algorithm 2), and introducing temporary workspace vectors $z$ and $s$, we may rewrite Algorithm 4 as Algorithm 5 5.
 Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
@@
 In the first, we maintain appropriate vectors $q_{i}\stackrel{{\scriptstyle\rm def}}{{=}}Sv_{i}$.
-We record this as Algorithm 6 6.
+We this as Algorithm 6 6.
 Input: symmetric $A\in\hbox{I\thinspace R}^{n\times n}$, $b\in\hbox{I\thinspace R}^{n}$, symmetric, positive-definite $S\in\hbox{I\thinspace R}^{n\times n}$, $\Delta>0$, a stopping tolerance $\epsilon>0$ and an iteration bound $m\geq 1$.
@@
 Computationally, the main difference is that Algorithm 6 requires an additional $S$ product per cycle, while Algorithm 3 has an additional four vector subtractions.
-Appendix C Detailed numerical experiments
+Appendix D Detailed numerical experiments
 Here we summarize the results from our numerical experiments that lead to Figure 7.1.
-As we said, we consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables, evaluate the gradient and Hessian to provide b-b and $A$, and for each a pair of radii $\Delta=1$ and then $0.5$.
-We apply the GALAHAD solvers TREK (Algorithm 2), TRS [GoulRobiThor10:mpc] and GLTR [GoulLuciRomaToin99:siopt] in each case; when $\Delta=0.5$, we use data accumulated from the $\Delta=1$ case to “warm start” the second solve as each solver provides for this.
+As we said, we consider the set of unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables, evaluate the gradient and Hessian to provide b-b and $A$, We apply the GALAHAD solvers TREK (Algorithm 2), TRS [GoulRobiThor10:mpc] and GLTR [GoulLuciRomaToin99:siopt] in each case
 In Table LABEL:table-a1 we report the name of each test problem, the size of the instance chosen, $n$ (many CUTEst examples have multiple dimensions), the radius value, $\Delta$ and the optimal objective function found.
@@
 For TRS, we report the number of factorizations and clock time required, while for GLTR, we state the number of vector products involving $A$ and again the time.
-The symbol $\dagger$} indicates an instance that the iteration bound $m=100$ was reached, while and $\ddagger$} indicates that a factorization failed through singularity.
-Table C.1: Subproblem solves for CUTEst examples: Complete results
+Both of the $\Delta$ regimes (a) and (b) are included for each solver.
+The symbol $\dagger$} indicates an instance that the iteration bound was reached, while and $\ddagger$} indicates that a factorization failed through singularity.
+Table D.1: Subproblem solves for CUTEst examples: Complete results
 TREK TRS GLTR
-name $n$ $\Delta$ $f(x_{*})$ $k_{\mbox{\tiny stop}}$ time facts time prods time
-ARWHEAD 5000 1.0 -9.99800000E+03 0 0.00 2 0.00 2 0.00
-ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
-BA-L16LS 66462 1.0 -4.00647969E+09 10 1.18 6 1.94 69 2.76
-BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 1.43 34 1.42
-BA-L16LS 66462 1.0 -4.00647969E+09 10 0.82 6 0.94 69 2.55
-BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.68 34 1.34
-ARWHEAD 5000 1.0 -9.99800000E+03 -1 0.00 2 0.00 2 0.00
-ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
-BA-L16LS 66462 1.0 -4.00647969E+09 10 0.84 6 0.94 69 2.60
-BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.73 34 1.32
-BA-L21LS 34134 1.0 -1.43350953E+08 100† 0.88 8 0.39 1129 18.09
-BA-L21LS 34134 0.5 -1.27987567E+08 100† 0.00 12 0.27 564 8.95
-BA-L49LS 23769 1.0 -2.32581840E+10 6 0.25 8 0.45 27 0.40
-BA-L49LS 23769 0.5 -5.82477997E+09 6 0.00 11 0.20 13 0.21
-BA-L52LS 192627 1.0 -1.78720629E+10 31 4.65 11 9.18 173 27.82
-BA-L52LS 192627 0.5 -1.26606809E+10 31 0.01 16 4.79 86 13.82
-BA-L73LS 33753 1.0 -9.88938047E+10 4 0.37 10 0.93 25 0.56
-BA-L73LS 33753 0.5 -2.47259422E+10 4 0.00 13 0.35 12 0.29
-BDQRTIC 5000 1.0 -4.70328224E+05 4 0.01 4 0.01 13 0.00
-BDQRTIC 5000 0.5 -4.37969301E+05 4 0.00 8 0.01 6 0.00
-BOX 10000 1.0 -4.99665149E+01 1 0.01 3 0.01 3 0.00
-BOX 10000 0.5 -2.49841935E+01 1 0.00 6 0.01 1 0.00
-BOXPOWER 20000 1.0 -1.27930301E+05 3 0.01 3 0.02 5 0.00
-BOXPOWER 20000 0.5 -7.39842920E+04 3 0.00 6 0.02 2 0.00
-BROYDN3DLS 5000 1.0 -5.47141790E+02 6 0.01 3 0.02 19 0.00
-BROYDN3DLS 5000 0.5 -2.78620552E+02 6 0.00 6 0.01 9 0.00
-BROYDN7D 5000 1.0 -1.04198440E+03 4 0.01 3 0.01 11 0.00
-BROYDN7D 5000 0.5 -5.29359803E+02 4 0.00 6 0.01 5 0.00
-BROYDNBDLS 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
-BROYDNBDLS 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
-BRYBND 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
-BRYBND 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
-CHAINWOO 4000 1.0 -4.18808861E+05 4 0.01 2 0.01 9 0.00
-CHAINWOO 4000 0.5 -2.10745351E+05 4 0.00 4 0.01 4 0.00
-COSINE 10000 1.0 -7.33802606E+01 5 0.03 2 0.02 11 0.00
-COSINE 10000 0.5 -3.63234227E+01 5 0.00 4 0.01 5 0.00
-CRAGGLVY 5000 1.0 -2.70357001E+05 2 0.02 3 0.01 5 0.00
-CRAGGLVY 5000 0.5 -1.38612789E+05 2 0.00 6 0.01 2 0.00
-CURLY10 10000 1.0 -2.55466912E+03 52 0.11 3 0.05 157 0.04
-CURLY10 10000 0.5 -6.72371854E+02 52 0.00 6 0.04 78 0.02
-CURLY20 10000 1.0 -9.12107858E+03 49 0.11 3 0.05 205 0.07
-CURLY20 10000 0.5 -2.35575388E+03 49 0.00 6 0.04 102 0.03
-CURLY30 10000 1.0 -1.97300910E+04 41 0.15 3 0.10 235 0.11
-CURLY30 10000 0.5 -5.06068395E+03 41 0.00 6 0.06 117 0.04
-DIXMAANB 3000 1.0 -1.94571746E+03 4 0.01 3 0.01 11 0.00
-DIXMAANB 3000 0.5 -9.82394549E+02 4 0.00 6 0.01 5 0.00
-DIXMAANC 3000 1.0 -3.67420974E+03 4 0.02 3 0.01 11 0.00
-DIXMAANC 3000 0.5 -1.85594245E+03 4 0.00 6 0.01 5 0.00
-DIXMAAND 3000 1.0 -7.40783941E+03 4 0.01 3 0.01 11 0.00
-DIXMAAND 3000 0.5 -3.74285037E+03 4 0.00 6 0.01 5 0.00
-DIXMAANE1 3000 1.0 -1.03848735E+03 4 0.01 3 0.00 11 0.00
-DIXMAANE1 3000 0.5 -5.25114487E+02 4 0.00 6 0.00 5 0.00
-DIXMAANF 3000 1.0 -1.83737329E+03 4 0.01 3 0.01 11 0.00
-DIXMAANF 3000 0.5 -9.28137690E+02 4 0.00 6 0.01 5 0.00
-DIXMAANG 3000 1.0 -3.56193345E+03 4 0.01 3 0.01 11 0.00
-DIXMAANG 3000 0.5 -1.79971801E+03 4 0.00 6 0.01 5 0.00
-DIXMAANH 3000 1.0 -7.28769893E+03 4 0.01 3 0.01 11 0.00
-DIXMAANH 3000 0.5 -3.68269064E+03 4 0.00 6 0.01 5 0.00
-DIXMAANI1 3000 1.0 -1.00064454E+03 4 0.01 3 0.00 11 0.00
-DIXMAANI1 3000 0.5 -5.06141189E+02 4 0.00 6 0.00 5 0.00
-DIXMAANJ 3000 1.0 -1.79984433E+03 4 0.01 3 0.01 11 0.00
-DIXMAANJ 3000 0.5 -9.09324798E+02 4 0.00 6 0.01 5 0.00
-DIXMAANK 3000 1.0 -3.52376478E+03 4 0.01 3 0.01 11 0.00
-DIXMAANK 3000 0.5 -1.78058448E+03 4 0.00 6 0.01 5 0.00
-DIXMAANL 3000 1.0 -7.24829529E+03 4 0.01 3 0.01 11 0.00
-DIXMAANL 3000 0.5 -3.66293885E+03 4 0.00 6 0.01 5 0.00
-DIXMAANM1 3000 1.0 -4.27493497E+02 5 0.01 3 0.01 13 0.00
-DIXMAANM1 3000 0.5 -2.16348948E+02 5 0.00 6 0.01 6 0.00
-DIXMAANN 3000 1.0 -9.97842749E+02 5 0.01 3 0.01 13 0.00
-DIXMAANN 3000 0.5 -5.05235340E+02 5 0.00 6 0.01 6 0.00
-DIXMAANO 3000 1.0 -1.90145642E+03 5 0.01 3 0.01 13 0.00
-DIXMAANO 3000 0.5 -9.63139983E+02 5 0.00 6 0.01 6 0.00
-DIXMAANP 3000 1.0 -3.85352076E+03 5 0.01 3 0.01 13 0.00
-DIXMAANP 3000 0.5 -1.95233964E+03 5 0.00 6 0.01 6 0.00
-DIXON3DQ 10000 1.0 -4.35180402E+00 19 0.02 5 0.03 45 0.01
-DIXON3DQ 10000 0.5 -2.48148299E+00 19 0.00 8 0.02 22 0.00
-DQDRTIC 5000 1.0 -8.50546818E+04 3 0.00 2 0.00 7 0.00
-DQDRTIC 5000 0.5 -4.25775883E+04 3 0.00 4 0.00 3 0.00
-DQRTIC 5000 1.0 -1.33489191E+13 2 0.00 2 0.00 5 0.00
-DQRTIC 5000 0.5 -6.67448869E+12 2 0.00 4 0.00 2 0.00
-EDENSCH 2000 1.0 -9.90061935E+04 3 0.00 2 0.00 7 0.00
-EDENSCH 2000 0.5 -4.96303271E+04 3 0.00 4 0.00 3 0.00
-EG2 1000 1.0 -1.73066127E+02 0 0.00 2 0.00 1 0.00
-EG2 1000 0.5 -1.64667237E+02 1 0.00 4 0.00 0 0.00
-EIGENALS 2550 1.0 -1.01221801E+03 7 0.75 3 0.85 25 0.04
-EIGENALS 2550 0.5 -4.77552997E+02 7 0.00 6 0.79 12 0.02
-EIGENBLS 2550 1.0 -3.52937908E+01 8 0.70 3 0.89 25 0.04
-EIGENBLS 2550 0.5 -1.80972156E+01 8 0.00 5 0.62 12 0.02
-EIGENCLS 2652 1.0 -4.70591220E+02 9 0.78 3 0.91 31 0.04
-EIGENCLS 2652 0.5 -2.33514688E+02 9 0.00 6 0.93 15 0.02
-ENGVAL1 5000 1.0 -8.67081566E+03 3 0.01 2 0.01 9 0.00
-ENGVAL1 5000 0.5 -4.35940622E+03 3 0.00 4 0.01 4 0.00
-EXTROSNB 1000 1.0 -3.66203611E+04 4 0.00 3 0.00 11 0.00
-EXTROSNB 1000 0.5 -1.86350898E+04 4 0.00 6 0.00 5 0.00
-FLETBV3M 5000 1.0 -4.36166522E+01 3 0.01 2 0.01 7 0.00
-FLETBV3M 5000 0.5 -2.18337960E+01 3 0.00 4 0.01 3 0.00
-FLETCBV2 5000 1.0 -3.88209445E-06 13 0.01 4 0.01 9737 412.07
-FLETCBV2 5000 0.5 -2.02683952E-06 16 0.00 8 0.01 4868 0.53
-FLETCBV3 5000 1.0 -4.36172752E+01 3 0.01 2 0.01 7 0.04
-FLETCBV3 5000 0.5 -2.18341109E+01 3 0.00 4 0.01 3 0.00
-FLETCHBV 5000 1.0 -2.74868337E+09 3 0.01 2 0.01 9 0.00
-FLETCHBV 5000 0.5 -1.37709990E+09 3 0.00 4 0.01 4 0.00
-FLETCHCR 1000 1.0 -1.08786732E+01 1 0.00 3 0.00 3 0.00
-FLETCHCR 1000 0.5 -1.05114880E+01 1 0.00 7 0.00 1 0.00
-FMINSRF2 5625 1.0 -2.74210232E-01 44 0.04 4 0.03 107 0.03
-FMINSRF2 5625 0.5 -1.42608286E-01 44 0.00 8 0.03 53 0.01
-FMINSURF 5625 1.0 -2.74396123E-01 44 5.28 4 8.28 107 0.03
-FMINSURF 5625 0.5 -1.42697922E-01 44 0.00 8 8.05 53 0.01
-FREUROTH 5000 1.0 -5.51793805E+04 3 0.01 3 0.01 7 0.00
-FREUROTH 5000 0.5 -2.75854293E+04 3 0.00 6 0.01 3 0.00
-GENHUMPS 5000 1.0 -6.64118303E+03 6 0.03 3 0.02 13 0.00
-GENHUMPS 5000 0.5 -3.16552923E+03 6 0.00 6 0.01 6 0.00
-INDEF 5000 1.0 -2.10493952E+03 6 0.01 103‡ 0.12 13 0.00
-INDEF 5000 0.5 -5.26825583E+02 6 0.00 207‡ 0.25 6 0.00
-INDEFM 100000 1.0 -4.20748078E+04 3 0.09 11 0.38 11 0.05
-INDEFM 100000 0.5 -1.05192638E+04 3 0.00 16 0.23 5 0.03
-JIMACK 3549 1.0 -4.37368349E-01 100† 0.45 26 0.57 3201 17.07
-JIMACK 3549 0.5 -2.41925217E-01 100† 0.00 31 0.14 1600 0.62
-LIARWHD 5000 1.0 -4.61798034E+05 1 0.00 3 0.00 3 0.01
-LIARWHD 5000 0.5 -2.36034628E+05 1 0.00 6 0.00 1 0.00
-MODBEALE 20000 1.0 -3.04981491E+05 2 0.05 2 0.03 7 0.01
-MODBEALE 20000 0.5 -1.52960464E+05 2 0.00 4 0.03 3 0.00
-MOREBV 5000 1.0 -7.72418940E-14 9 0.01 107‡ 0.13 295 0.17
-MOREBV 5000 0.5 -7.72418940E-14 54 0.03 116 0.02 294 0.02
-MSQRTALS 1024 1.0 -3.70040042E+02 4 0.16 4 0.14 27 0.01
-MSQRTALS 1024 0.5 -1.75608597E+02 5 0.00 7 0.11 13 0.01
-MSQRTBLS 1024 1.0 -3.69312117E+02 4 0.16 4 0.14 27 0.01
-MSQRTBLS 1024 0.5 -1.75279355E+02 5 0.00 7 0.12 13 0.01
-NCB20 5010 1.0 -2.77922014E+02 6 0.03 4 0.02 27 0.07
-NCB20 5010 0.5 -1.39893162E+02 6 0.00 7 0.02 13 0.03
-NCB20B 5000 1.0 -2.77950996E+02 6 0.03 4 0.02 27 0.05
-NCB20B 5000 0.5 -1.39907519E+02 6 0.00 7 0.02 13 0.02
-NONCVXU2 5000 1.0 -3.33554888E+06 2 0.21 2 0.34 5 0.00
-NONCVXU2 5000 0.5 -1.66777663E+06 2 0.00 4 0.34 2 0.00
-NONCVXUN 5000 1.0 -3.56003262E+06 2 0.02 2 0.02 5 0.00
-NONCVXUN 5000 0.5 -1.78001885E+06 2 0.00 4 0.02 2 0.00
-NONDIA 5000 1.0 -1.49970308E+06 1 0.01 3 0.01 3 0.00
-NONDIA 5000 0.5 -8.75226609E+05 1 0.00 6 0.01 1 0.00
-NONDQUAR 5000 1.0 -3.33683482E+03 43 0.03 10 0.02 117 0.01
-NONDQUAR 5000 0.5 -3.33513581E+03 43 0.00 15 0.01 58 0.00
-NONMSQRT 4900 1.0 -1.15116171E+03 8 0.07 5 0.03 41 0.17
-NONMSQRT 4900 0.5 -5.50172607E+02 8 0.00 9 0.03 20 0.08
-OSCIGRAD 100000 1.0 -5.34690566E+08 8 0.31 5 0.23 43 0.13
-OSCIGRAD 100000 0.5 -4.31017580E+08 8 0.00 9 0.20 21 0.09
-PENALTY1 1000 1.0 -2.43960328E+13 1 0.07 2 0.07 1 0.00
-PENALTY1 1000 0.5 -1.21985172E+13 1 0.00 4 0.05 0 0.00
-POWELLSG 5000 1.0 -1.57803913E+04 2 0.01 3 0.01 7 0.00
-POWELLSG 5000 0.5 -7.99992041E+03 2 0.00 6 0.01 3 0.00
-POWER 10000 1.0 -1.13407283E+14 5 13.59 3 29.47 11 0.00
-POWER 10000 0.5 -5.72243018E+13 5 0.00 6 29.24 5 0.00
-QUARTC 5000 1.0 -1.33489191E+13 2 0.00 2 0.00 5 0.00
-QUARTC 5000 0.5 -6.67448869E+12 2 0.00 4 0.00 2 0.00
-SBRYBND 5000 1.0 -9.63032705E+11 20 0.03 16 0.03 179 0.06
-SBRYBND 5000 0.5 -2.40760941E+11 20 0.00 21 0.02 89 0.03
-SCHMVETT 5000 1.0 -7.28426352E+01 10 0.01 3 0.01 35 0.01
-SCHMVETT 5000 0.5 -3.68824395E+01 10 0.00 6 0.01 17 0.01
-SCOSINE 5000 1.0 -8.25379341E+10 13 0.03 11 0.02 187 0.02
-SCOSINE 5000 0.5 -2.06344856E+10 13 0.00 13 0.01 93 0.01
-SCURLY10 10000 1.0 -1.28077199E+30 4 0.04 3 0.03 11 0.00
-SCURLY10 10000 0.5 -6.43068284E+29 4 0.00 6 0.03 5 0.00
-SCURLY20 10000 1.0 -1.63614356E+31 4 0.07 3 0.04 11 0.00
-SCURLY20 10000 0.5 -8.21427988E+30 4 0.00 6 0.04 5 0.00
-SCURLY30 10000 1.0 -7.47542209E+31 4 0.09 3 0.06 9 0.01
-SCURLY30 10000 0.5 -3.75274481E+31 4 0.00 6 0.06 4 0.00
-SINQUAD 5000 1.0 -7.12672063E+03 2 0.00 4 0.00 5 0.00
-SINQUAD 5000 0.5 -2.75259696E+03 2 0.00 8 0.00 2 0.00
-SPARSINE 5000 1.0 -2.92330968E+06 4 0.48 3 1.16 11 0.00
-SPARSINE 5000 0.5 -1.46942254E+06 4 0.00 6 1.18 5 0.00
-SPARSQUR 10000 1.0 -1.19555581E+06 7 3.09 2 5.80 15 0.01
-SPARSQUR 10000 0.5 -6.09140674E+05 7 0.00 4 5.86 7 0.00
-SPMSRTLS 4999 1.0 -8.16740777E+01 7 0.02 3 0.01 21 0.01
-SPMSRTLS 4999 0.5 -3.97487352E+01 7 0.00 6 0.01 10 0.00
-SROSENBR 5000 1.0 -1.15690489E+02 9 0.01 4 0.01 5 0.00
-SROSENBR 5000 0.5 -6.52884475E+01 9 0.00 8 0.01 2 0.00
-SSBRYBND 5000 1.0 -5.98063290E+06 17 0.03 12 0.04 145 0.09
-SSBRYBND 5000 0.5 -1.51010118E+06 17 0.00 15 0.02 72 0.04
-SSCOSINE 5000 1.0 -5.13429696E+05 12 0.02 16 0.02 155 0.02
-SSCOSINE 5000 0.5 -1.28378413E+05 12 0.00 19 0.01 77 0.01
-TOINTGSS 5000 1.0 -4.23179011E+02 4 0.02 2 0.02 9 0.00
-TOINTGSS 5000 0.5 -2.11839554E+02 4 0.00 4 0.01 4 0.00
-TQUARTIC 5000 1.0 -2.91590249E-02 1 0.00 3 0.00 3 0.00
-TQUARTIC 5000 0.5 -1.66434134E-02 1 0.00 6 0.00 1 0.00
-TRIDIA 10000 1.0 -1.14762126E+06 7 0.02 3 0.02 15 0.00
-TRIDIA 10000 0.5 -5.75687744E+05 7 0.00 6 0.02 7 0.00
-VAREIGVL 5000 1.0 -1.03187418E+04 3 2.68 3 4.55 9 0.00
-VAREIGVL 5000 0.5 -5.19645395E+03 3 0.00 6 4.58 4 0.00
-WOODS 4000 1.0 -5.13132992E+05 2 0.01 3 0.01 7 0.00
-WOODS 4000 0.5 -2.57913812E+05 2 0.00 6 0.01 3 0.00
-YATP1LS 10200 1.0 -2.47600401E+06 1 4.29 3 11.48 3 0.01
-YATP1LS 10200 0.5 -1.24008348E+06 1 0.00 5 7.65 1 0.00
-YATP2LS 10200 1.0 -6.17000891E+04 1 8.21 3 11.89 3 0.00
-YATP2LS 10200 0.5 -3.11097632E+04 1 0.00 5 7.77 1 0.00
-Table C.1: Subproblem solves for CUTEst examples: Complete results (continued)
+(a) (b) (a) (b) (a) (b)
+name $n$ $\Delta$ $f(x_{*})$ $k_{\mbox{\tiny s}}$ time $k_{\mbox{\tiny s}}$ time fac time fac time pr time pr time
+ARWHEAD 5000 10 -9.99800000E+03 0 0.01 2 0.00 2 0.00
+ARWHEAD 5000 .1 -3.59936000E+03 1 0.00 0 0.01 4 0.00 2 0.00 1 0.00 2 0.00
+ARWHEAD 5000 .01 -3.95930600E+02 1 0.00 1 0.00 6 0.00 4 0.00 1 0.00 1 0.00
+BA-L16LS 66462 10 -3.46265942E+11 12 0.91 9 1.28 75 2.11
+BA-L16LS 66462 1 -4.00647969E+09 12 0.00 10 0.82 13 0.79 6 1.06 37 1.08 69 1.96
+BA-L16LS 66462 .1 -2.71798703E+08 12 0.00 10 0.00 19 0.98 12 1.02 37 1.09 34 1.01
+BA-L21LS 34134 10 -1.09802044E+09 284 11.60 15 0.70 1999 22.79
+BA-L21LS 34134 1 -1.43350953E+08 284 0.15 196 3.34 21 0.39 8 0.43 999 11.38 1129 13.46
+BA-L21LS 34134 1 -1.43350953E+08 284 0.15 196 0.05 22 0.15 14 0.39 999 11.36 564 6.62
+BA-L49LS 23769 10 -2.32220041E+12 6 0.25 9 0.57 29 0.32
+BA-L49LS 23769 1 -2.32581840E+10 6 0.00 6 0.24 12 0.25 8 0.59 14 0.17 27 0.31
+BA-L49LS 23769 .1 -2.36543531E+08 6 0.00 6 0.00 15 0.26 11 0.28 14 0.17 13 0.17
+BA-L52LS 192627 10 -5.93753593E+11 34 5.24 9 8.15 221 25.43
+BA-L52LS 192627 1 -1.78720629E+10 34 0.01 31 4.98 14 5.33 11 10.32 110 12.86 173 20.84
+BA-L52LS 192627 .1 -5.16366984E+09 34 0.01 31 0.01 18 4.42 15 4.81 110 12.84 86 10.49
+BA-L73LS 33753 10 -9.88855550E+12 5 0.34 12 1.12 27 0.44
+BA-L73LS 33753 1 -9.88938047E+10 5 0.00 4 0.32 15 0.42 10 0.97 13 0.24 25 0.41
+BA-L73LS 33753 .1 -9.90136989E+08 5 0.00 4 0.00 18 0.42 13 0.46 13 0.24 12 0.22
+BDQRTIC 5000 10 -6.53953444E+05 6 0.00 3 0.01 29 0.01
+BDQRTIC 5000 1 -4.70328224E+05 6 0.00 4 0.00 7 0.02 4 0.02 14 0.00 13 0.00
+BDQRTIC 5000 .1 -1.37454488E+05 6 0.00 4 0.00 10 0.01 7 0.01 14 0.00 6 0.00
+BOX 10000 10 -4.99340865E+02 1 0.03 3 0.01 3 0.00
+BOX 10000 1 -4.99665149E+01 1 0.00 1 0.04 6 0.01 3 0.02 1 0.00 3 0.00
+BOX 10000 .1 -4.99704061E+00 1 0.00 1 0.00 9 0.01 6 0.01 1 0.00 1 0.00
+BOXPOWER 20000 10 -1.76398331E+05 24 0.15 6 0.04 9 0.00
+BOXPOWER 20000 1 -1.27930301E+05 24 0.00 2 0.18 9 0.02 3 0.02 4 0.00 5 0.00
+BOXPOWER 20000 .1 -1.63999277E+04 24 0.00 2 0.00 11 0.02 5 0.02 4 0.00 2 0.00
+BROYDN3DLS 5000 10 -3.66408186E+03 6 0.00 3 0.01 45 0.00
+BROYDN3DLS 5000 1 -5.47141790E+02 6 0.00 6 0.01 6 0.01 3 0.01 22 0.00 19 0.00
+BROYDN3DLS 5000 .1 -5.65333513E+01 6 0.00 6 0.00 9 0.01 6 0.01 22 0.00 9 0.00
+BROYDN7D 5000 10 -7.40764118E+03 7 0.00 3 0.01 31 0.00
+BROYDN7D 5000 1 -1.04198440E+03 7 0.00 4 0.01 6 0.01 3 0.01 15 0.00 11 0.00
+BROYDN7D 5000 .1 -1.07210780E+02 7 0.00 4 0.00 8 0.01 5 0.01 15 0.00 5 0.00
+BROYDNBDLS 5000 10 -5.46819825E+04 4 0.01 4 0.02 31 0.01
+BROYDNBDLS 5000 1 -7.54822794E+03 5 0.00 5 0.01 7 0.02 3 0.02 15 0.00 15 0.00
+BROYDNBDLS 5000 .1 -7.75634313E+02 5 0.00 5 0.00 9 0.01 5 0.01 15 0.00 7 0.00
+BRYBND 5000 10 -5.46819825E+04 4 0.01 4 0.02 31 0.01
+BRYBND 5000 1 -7.54822794E+03 5 0.00 5 0.01 7 0.02 3 0.02 15 0.00 15 0.00
+BRYBND 5000 .1 -7.75634313E+02 5 0.00 5 0.00 9 0.01 5 0.01 15 0.00 7 0.00
+CHAINWOO 4000 10 -3.70797995E+06 5 0.00 3 0.01 15 0.00
+CHAINWOO 4000 1 -4.18808861E+05 5 0.00 4 0.00 5 0.01 2 0.00 7 0.00 9 0.00
+CHAINWOO 4000 .1 -4.23637069E+04 5 0.00 4 0.00 7 0.01 4 0.00 7 0.00 4 0.00
+COSINE 10000 10 -8.65819784E+02 12 0.01 3 0.01 27 0.00
+COSINE 10000 1 -7.33802606E+01 12 0.00 5 0.01 5 0.01 2 0.01 13 0.00 11 0.00
+COSINE 10000 .1 -7.20601140E+00 12 0.00 5 0.00 7 0.01 4 0.01 13 0.00 5 0.00
+CRAGGLVY 5000 10 -1.47282263E+06 2 0.00 6 0.01 11 0.00
+CRAGGLVY 5000 1 -2.70357001E+05 2 0.00 2 0.00 9 0.01 3 0.01 5 0.00 5 0.00
+CRAGGLVY 5000 .1 -2.82720573E+04 2 0.00 2 0.00 11 0.01 5 0.01 5 0.00 2 0.00
+CURLY10 10000 10 -2.43341009E+05 62 0.09 3 0.08 519 0.10
+CURLY10 10000 1 -2.55466912E+03 62 0.00 52 0.07 7 0.05 3 0.05 259 0.05 157 0.03
+CURLY10 10000 .1 -3.76837984E+01 62 0.00 52 0.00 10 0.05 6 0.05 259 0.05 78 0.01
+CURLY20 10000 10 -8.84977907E+05 36 0.07 4 0.13 677 0.21
+CURLY20 10000 1 -9.12107858E+03 49 0.02 49 0.09 8 0.11 3 0.10 338 0.09 205 0.05
+CURLY20 10000 .1 -1.18405527E+02 49 0.00 49 0.00 11 0.10 6 0.09 338 0.09 102 0.03
+CURLY30 10000 10 -1.92700208E+06 26 0.08 4 0.18 769 0.28
+CURLY30 10000 1 -1.97300910E+04 40 0.02 40 0.10 8 0.14 3 0.14 384 0.13 235 0.09
+CURLY30 10000 .1 -2.43499809E+02 40 0.00 40 0.00 11 0.12 6 0.14 384 0.14 117 0.04
+DIXMAANB 3000 10 -1.60339163E+04 7 0.00 3 0.01 23 0.00
+DIXMAANB 3000 1 -1.94571746E+03 7 0.00 4 0.00 6 0.01 3 0.01 11 0.00 11 0.00
+DIXMAANB 3000 .1 -1.98005001E+02 7 0.00 4 0.00 8 0.01 5 0.01 11 0.00 5 0.00
+DIXMAANC 3000 10 -2.99803804E+04 7 0.00 3 0.01 25 0.00
+DIXMAANC 3000 1 -3.67420974E+03 7 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANC 3000 .1 -3.74203237E+02 7 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAAND 3000 10 -6.01051338E+04 8 0.00 3 0.01 25 0.00
+DIXMAAND 3000 1 -7.40783941E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAAND 3000 .1 -7.54800526E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANE1 3000 10 -8.27329007E+03 6 0.00 3 0.00 21 0.00
+DIXMAANE1 3000 1 -1.03848735E+03 6 0.00 4 0.00 6 0.00 3 0.00 10 0.00 11 0.00
+DIXMAANE1 3000 .1 -1.05962269E+02 6 0.00 4 0.00 8 0.00 5 0.00 10 0.00 5 0.00
+DIXMAANF 3000 10 -1.49809432E+04 8 0.00 3 0.01 25 0.00
+DIXMAANF 3000 1 -1.83737329E+03 8 0.00 4 0.01 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANF 3000 .1 -1.87140059E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANG 3000 10 -2.88885934E+04 8 0.00 3 0.01 25 0.00
+DIXMAANG 3000 1 -3.56193345E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANG 3000 .1 -3.62944535E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANH 3000 10 -5.89358136E+04 8 0.00 3 0.01 25 0.00
+DIXMAANH 3000 1 -7.28769893E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANH 3000 .1 -7.42754250E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANI1 3000 10 -7.91454338E+03 6 0.00 3 0.00 21 0.00
+DIXMAANI1 3000 1 -1.00064454E+03 6 0.00 4 0.00 6 0.00 3 0.00 10 0.00 11 0.00
+DIXMAANI1 3000 .1 -1.02159327E+02 6 0.00 4 0.00 8 0.00 5 0.00 10 0.00 5 0.00
+DIXMAANJ 3000 10 -1.46232627E+04 8 0.00 3 0.01 25 0.00
+DIXMAANJ 3000 1 -1.79984433E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANJ 3000 .1 -1.83369741E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANK 3000 10 -2.85248120E+04 8 0.01 3 0.01 25 0.00
+DIXMAANK 3000 1 -3.52376478E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANK 3000 .1 -3.59109964E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANL 3000 10 -5.85599682E+04 8 0.00 3 0.01 25 0.00
+DIXMAANL 3000 1 -7.24829529E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
+DIXMAANL 3000 .1 -7.38795901E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
+DIXMAANM1 3000 10 -3.35446625E+03 9 0.00 3 0.00 33 0.00
+DIXMAANM1 3000 1 -4.27493497E+02 9 0.00 5 0.00 6 0.00 3 0.00 16 0.00 13 0.00
+DIXMAANM1 3000 .1 -4.36867895E+01 9 0.00 5 0.00 8 0.00 5 0.00 16 0.00 6 0.00
+DIXMAANN 3000 10 -7.76833364E+03 10 0.00 3 0.01 33 0.00
+DIXMAANN 3000 1 -9.97842749E+02 10 0.00 5 0.00 6 0.01 3 0.01 16 0.00 13 0.00
+DIXMAANN 3000 .1 -1.02059568E+02 10 0.00 5 0.00 8 0.01 5 0.01 16 0.00 6 0.00
+DIXMAANO 3000 10 -1.46773625E+04 11 0.00 3 0.01 33 0.00
+DIXMAANO 3000 1 -1.90145642E+03 11 0.00 5 0.00 6 0.01 3 0.01 16 0.00 13 0.00
+DIXMAANO 3000 .1 -1.94618568E+02 11 0.00 5 0.00 8 0.01 5 0.01 16 0.00 6 0.00
+DIXMAANP 3000 10 -2.96054573E+04 11 0.00 3 0.01 35 0.00
+DIXMAANP 3000 1 -3.85352076E+03 11 0.00 5 0.00 6 0.01 3 0.01 17 0.00 13 0.00
+DIXMAANP 3000 .1 -3.94570550E+02 11 0.00 5 0.00 8 0.01 5 0.01 17 0.00 6 0.00
+DIXON3DQ 10000 10 -7.95918012E+00 89 0.14 9 0.03 2359 0.23
+DIXON3DQ 10000 1 -4.35180402E+00 89 0.01 19 0.01 15 0.02 5 0.02 1179 0.10 45 0.00
+DIXON3DQ 10000 .1 -5.50941460E-01 89 0.01 19 0.00 18 0.01 8 0.01 1179 0.11 22 0.00
+DQDRTIC 5000 10 -8.32457765E+05 3 0.00 3 0.00 9 0.00
+DQDRTIC 5000 1 -8.50546818E+04 3 0.00 3 0.00 5 0.00 2 0.00 4 0.00 7 0.00
+DQDRTIC 5000 .1 -8.52355726E+03 3 0.00 3 0.00 7 0.00 4 0.00 4 0.00 3 0.00
+DQRTIC 5000 10 -1.33478697E+14 3 0.00 2 0.00 5 0.00
+DQRTIC 5000 1 -1.33489191E+13 3 0.00 2 0.00 4 0.00 2 0.00 2 0.00 5 0.00
+DQRTIC 5000 .1 -1.33490240E+12 3 0.00 2 0.00 6 0.00 4 0.00 2 0.00 2 0.00
+EDENSCH 2000 10 -9.44259112E+05 5 0.00 3 0.00 13 0.00
+EDENSCH 2000 1 -9.90061935E+04 5 0.00 3 0.00 5 0.00 2 0.00 6 0.00 7 0.00
+EDENSCH 2000 .1 -9.94642228E+03 5 0.00 3 0.00 7 0.00 4 0.00 6 0.00 3 0.00
+EG2 1000 10 -1.73066127E+02 0 0.00 2 0.00 1 0.00
+EG2 1000 .1 -4.97676498E+01 1 0.00 0 0.00 4 0.00 2 0.00 0 0.00 1 0.00
+EG2 1000 .01 -5.35553453E+00 1 0.00 1 0.00 6 0.00 4 0.00 0 0.00 0 0.00
+EIGENALS 2550 10 -2.15497836E+04 8 0.50 4 1.12 75 0.08
+EIGENALS 2550 1 -1.01221801E+03 8 0.00 7 0.46 7 0.91 3 1.07 37 0.04 25 0.03
+EIGENALS 2550 .1 -9.10252017E+01 8 0.00 7 0.00 10 0.90 6 1.14 37 0.04 12 0.02
+[…]
+EIGENBLS 2550 1 -3.52937908E+01 49 0.00 8 0.55 116 0.91 3 1.07 301 0.34 25 0.03
+EIGENBLS 2550 .1 -3.70678144E+00 49 0.00 8 0.00 119 0.90 6 0.96 301 0.32 12 0.02
+EIGENCLS 2652 10 -9.33034837E+03 9 0.57 4 1.26 83 0.10
+EIGENCLS 2652 1 -4.70591220E+02 9 0.00 9 0.54 7 0.99 3 1.08 41 0.05 31 0.04
+EIGENCLS 2652 .1 -4.70009302E+01 9 0.00 9 0.00 9 0.77 5 0.81 41 0.05 15 0.02
+ENGVAL1 5000 10 -7.80687659E+04 5 0.00 3 0.01 15 0.00
+ENGVAL1 5000 1 -8.67081566E+03 5 0.00 3 0.00 5 0.01 2 0.01 7 0.00 9 0.00
+ENGVAL1 5000 .1 -8.75720987E+02 5 0.00 3 0.00 7 0.01 4 0.01 7 0.00 4 0.00
+EXTROSNB 1000 10 -2.49243249E+05 6 0.00 3 0.00 27 0.00
+EXTROSNB 1000 1 -3.66203611E+04 6 0.00 4 0.00 6 0.00 3 0.00 13 0.00 11 0.00
+EXTROSNB 1000 .1 -3.77900359E+03 6 0.00 4 0.00 8 0.00 5 0.00 13 0.00 5 0.00
+FLETBV3M 5000 10 -4.27009603E+02 4 0.00 3 0.01 11 0.00
+FLETBV3M 5000 1 -4.36166522E+01 4 0.00 3 0.00 5 0.01 2 0.01 5 0.00 7 0.00
+FLETBV3M 5000 .1 -4.37083490E+00 4 0.00 3 0.00 7 0.01 4 0.01 5 0.00 3 0.00
+FLETCBV2 5000 10 -1.81067238E-05 0 0.00 2 0.01 4869 0.47
+FLETCBV2 5000 1 -3.88209445E-06 8 0.00 8 0.01 6 0.01 4 0.01 4868 0.44 9737 1.47
+FLETCBV2 5000 .1 -4.25968981E-07 16 0.00 16 0.00 10 0.01 8 0.01 4868 0.45 4868 0.47
+FLETCBV3 5000 10 -4.27014602E+02 4 0.00 3 0.01 11 0.00
+FLETCBV3 5000 1 -4.36172752E+01 4 0.00 3 0.00 5 0.01 2 0.01 5 0.00 7 0.00
+FLETCBV3 5000 .1 -4.37089841E+00 4 0.00 3 0.00 7 0.01 4 0.01 5 0.00 3 0.00
+FLETCHBV 5000 10 -2.64954410E+10 4 0.00 3 0.01 13 0.00
+FLETCHBV 5000 1 -2.74868337E+09 4 0.00 3 0.00 5 0.01 2 0.01 6 0.00 9 0.00
+FLETCHBV 5000 .1 -2.75861358E+08 4 0.00 3 0.00 7 0.01 4 0.01 6 0.00 4 0.00
+FLETCHCR 1000 10 -1.08811881E+01 0 0.00 2 0.00 2 0.00
+FLETCHCR 1000 1 -1.08786732E+01 1 0.00 1 0.00 5 0.00 3 0.00 1 0.00 3 0.00
+FLETCHCR 1000 .1 -5.31285550E+00 1 0.00 1 0.00 8 0.00 6 0.00 1 0.00 1 0.00
+FMINSRF2 5625 10 -2.48385848E+00 130 0.37 4 0.04 331 0.07
+FMINSRF2 5625 1 -2.74210232E-01 130 0.01 42 0.03 8 0.03 4 0.04 165 0.04 107 0.02
+FMINSRF2 5625 1 -2.74210232E-01 130 0.01 42 0.00 9 0.02 8 0.03 165 0.04 53 0.01
+FMINSURF 5625 10 -2.48582974E+00 129 9.32 4 8.90 331 0.09
+FMINSURF 5625 1 -2.74396123E-01 129 0.01 42 4.93 8 8.57 4 9.22 165 0.04 107 0.02
+FMINSURF 5625 1 -2.74396123E-01 129 0.01 42 0.00 9 3.25 8 9.10 165 0.04 53 0.01
+FREUROTH 5000 10 -5.53358711E+05 4 0.00 3 0.01 11 0.00
+FREUROTH 5000 1 -5.51793805E+04 4 0.00 3 0.00 6 0.01 3 0.01 5 0.00 7 0.00
+FREUROTH 5000 .1 -5.51640621E+03 4 0.00 3 0.00 8 0.01 5 0.01 5 0.00 3 0.00
+GENHUMPS 5000 10 -1.22237034E+05 13 0.01 3 0.01 21 0.01
+GENHUMPS 5000 1 -6.64118303E+03 13 0.00 6 0.00 6 0.01 3 0.01 10 0.00 13 0.00
+GENHUMPS 5000 .1 -6.08296147E+02 13 0.00 6 0.00 8 0.01 5 0.01 10 0.00 6 0.00
+[…]
+[…]
+[…]
+[…]
+INDEFM 100000 1 -4.20748078E+04 184 1.57 70 0.38 120 0.35 11 0.37 5 0.01 11 0.03
+INDEFM 100000 1 -4.20748078E+04 184 0.09 70 0.01 121 0.08 15 0.18 5 0.01 5 0.01
+[…]
+[…]
+JIMACK 3549 1 -4.37368349E-01 24 0.05 24 0.05 33 0.04 34 0.15 1723 0.85 1600 0.96
+LIARWHD 5000 10 -2.76920956E+06 1 0.01 4 0.01 3 0.00
+LIARWHD 5000 1 -4.61798034E+05 1 0.00 1 0.01 7 0.00 3 0.00 1 0.00 3 0.00
+LIARWHD 5000 .1 -4.80286236E+04 1 0.00 1 0.00 9 0.00 5 0.00 1 0.00 1 0.00
+MODBEALE 20000 10 -2.88071792E+06 3 0.01 3 0.03 9 0.01
+MODBEALE 20000 1 -3.04981491E+05 3 0.00 2 0.01 5 0.02 2 0.03 4 0.00 7 0.02
+MODBEALE 20000 .1 -3.06672478E+04 3 0.00 2 0.00 7 0.02 4 0.03 4 0.00 3 0.01
+MOREBV 5000 10 -7.72418940E-14 0 0.00 2 0.01 295 0.02
+[…]
+MOREBV 5000 .1 -7.72418940E-14 8 0.00 8 0.00 119 0.02 117 0.02 294 0.02 294 0.02
+MSQRTALS 1024 10 -7.41223812E+03 16 0.14 4 0.15 85 0.03
+MSQRTALS 1024 1 -3.70040042E+02 16 0.00 4 0.12 8 0.14 4 0.15 42 0.02 27 0.01
+MSQRTALS 1024 .1 -3.36456076E+01 16 0.00 4 0.00 11 0.12 7 0.13 42 0.02 13 0.00
+MSQRTBLS 1024 10 -7.39900251E+03 16 0.14 4 0.15 85 0.03
+MSQRTBLS 1024 1 -3.69312117E+02 16 0.00 4 0.12 8 0.15 4 0.16 42 0.02 27 0.01
+MSQRTBLS 1024 .1 -3.35861466E+01 16 0.00 4 0.00 11 0.12 7 0.13 42 0.02 13 0.01
+NCB20 5010 10 -2.55084157E+03 13 0.02 4 0.04 57 0.09
+NCB20 5010 1 -2.77922014E+02 13 0.00 6 0.02 8 0.04 4 0.05 28 0.05 27 0.05
+NCB20 5010 .1 -2.81547764E+01 13 0.00 6 0.00 11 0.04 7 0.04 28 0.05 13 0.02
+NCB20B 5000 10 -2.55116152E+03 13 0.02 4 0.04 57 0.14
+NCB20B 5000 1 -2.77950996E+02 13 0.00 6 0.02 8 0.04 4 0.05 28 0.07 27 0.07
+NCB20B 5000 .1 -2.81576212E+01 13 0.00 6 0.00 11 0.04 7 0.06 28 0.07 13 0.03
+NONCVXU2 5000 10 -3.33546996E+07 3 0.18 2 0.35 5 0.00
+NONCVXU2 5000 1 -3.33554888E+06 3 0.00 2 0.18 4 0.31 2 0.35 2 0.00 5 0.00
+NONCVXU2 5000 .1 -3.33555677E+05 3 0.00 2 0.00 6 0.31 4 0.33 2 0.00 2 0.00
+NONCVXUN 5000 10 -3.55994124E+07 3 0.01 2 0.02 5 0.00
+NONCVXUN 5000 1 -3.56003262E+06 3 0.00 2 0.01 4 0.02 2 0.02 2 0.00 5 0.00
+NONCVXUN 5000 .1 -3.56004176E+05 3 0.00 2 0.00 6 0.02 4 0.02 2 0.00 2 0.00
+[…]
+NONDIA 5000 1 -1.49970308E+06 1 0.02 1 0.02 14 0.00 3 0.00 1 0.00 3 0.00
+NONDIA 5000 1 -1.49970308E+06 1 0.01 1 0.00 15 0.00 5 0.00 1 0.00 1 0.00
+NONDQUAR 5000 10 -3.33983507E+03 40 0.02 9 0.02 2693 0.23
+NONDQUAR 5000 1 -3.33683482E+03 44 0.00 44 0.02 15 0.01 10 0.03 1346 0.07 117 0.01
+NONDQUAR 5000 .1 -1.70026980E+03 44 0.00 44 0.00 18 0.01 13 0.01 1346 0.07 58 0.01
+NONMSQRT 4900 10 -3.41606798E+04 33 0.07 19 0.08 161 0.62
+NONMSQRT 4900 1 -1.15116171E+03 33 0.00 8 0.05 25 0.03 5 0.03 80 0.30 41 0.16
+NONMSQRT 4900 .1 -1.10670950E+02 33 0.00 8 0.00 28 0.02 8 0.02 80 0.30 20 0.08
+OSCIGRAD 100000 10 -8.09081071E+09 10 0.37 5 0.28 53 0.13
+OSCIGRAD 100000 1 -5.34690566E+08 10 0.00 8 0.34 9 0.24 5 0.29 26 0.07 43 0.12
+OSCIGRAD 100000 .1 -1.90912328E+08 10 0.00 8 0.00 12 0.21 8 0.23 26 0.06 21 0.06
+PENALTY1 1000 10 -2.43780058E+14 1 0.04 2 0.07 1 0.00
+PENALTY1 1000 1 -2.43960328E+13 1 0.00 1 0.04 4 0.07 2 0.07 0 0.00 1 0.00
+PENALTY1 1000 .1 -2.43978355E+12 1 0.00 1 0.00 6 0.07 4 0.07 0 0.00 0 0.00
+POWELLSG 5000 10 -1.20598070E+05 2 0.00 3 0.01 7 0.00
+POWELLSG 5000 1 -1.57803913E+04 2 0.00 2 0.00 6 0.00 3 0.01 3 0.00 7 0.00
+POWELLSG 5000 .1 -1.61760603E+03 2 0.00 2 0.00 8 0.00 5 0.00 3 0.00 3 0.00
+POWER 10000 10 -9.47415692E+14 8 12.41 3 30.58 19 0.00
+POWER 10000 1 -1.13407283E+14 8 0.00 5 11.73 6 29.81 3 32.33 9 0.00 11 0.00
+POWER 10000 .1 -1.15281904E+13 8 0.00 5 0.00 8 21.72 5 22.82 9 0.00 5 0.00
+QUARTC 5000 10 -1.33478697E+14 3 0.00 2 0.00 5 0.00
+QUARTC 5000 1 -1.33489191E+13 3 0.00 2 0.00 4 0.00 2 0.00 2 0.00 5 0.00
+QUARTC 5000 .1 -1.33490240E+12 3 0.00 2 0.00 6 0.00 4 0.00 2 0.00 2 0.00
+SBRYBND 5000 10 -9.63022762E+13 22 0.01 17 0.03 193 0.05
+SBRYBND 5000 1 -9.63032704E+11 22 0.00 20 0.01 20 0.02 16 0.05 96 0.03 179 0.05
+SBRYBND 5000 .1 -9.63132614E+09 22 0.00 20 0.00 23 0.02 19 0.02 96 0.03 89 0.02
+SCHMVETT 5000 10 -5.62455481E+02 10 0.00 3 0.01 113 0.03
+SCHMVETT 5000 1 -7.28426352E+01 10 0.00 10 0.00 6 0.01 3 0.01 56 0.02 35 0.01
+SCHMVETT 5000 .1 -7.45027318E+00 10 0.00 10 0.00 9 0.01 6 0.01 56 0.02 17 0.01
+SCOSINE 5000 10 -8.25379286E+12 18 0.01 13 0.02 195 0.02
+SCOSINE 5000 1 -8.25379341E+10 18 0.00 13 0.01 15 0.01 12 0.01 97 0.01 187 0.02
+SCOSINE 5000 .1 -8.25380118E+08 18 0.00 13 0.00 18 0.01 15 0.01 97 0.01 93 0.01
+SCURLY10 10000 10 -1.18559957E+31 1 0.02 3 0.05 19 0.00
+SCURLY10 10000 1 -1.28077199E+30 1 0.00 1 0.02 6 0.04 3 0.05 9 0.00 11 0.00
+SCURLY10 10000 .1 -1.29043382E+29 1 0.00 1 0.00 8 0.04 5 0.04 9 0.00 5 0.00
+SCURLY20 10000 10 -1.51685064E+32 1 0.04 3 0.09 17 0.01
+SCURLY20 10000 1 -1.63614356E+31 1 0.00 1 0.03 6 0.09 3 0.09 8 0.00 11 0.00
+SCURLY20 10000 .1 -1.64823202E+30 1 0.00 1 0.00 8 0.08 5 0.08 8 0.00 5 0.00
+SCURLY30 10000 10 -6.94024457E+32 1 0.05 3 0.13 17 0.01
+SCURLY30 10000 1 -7.47542209E+31 1 0.00 1 0.05 6 0.13 3 0.13 8 0.00 9 0.00
+SCURLY30 10000 .1 -7.52956757E+30 1 0.00 1 0.00 8 0.11 5 0.11 8 0.00 4 0.00
+SINQUAD 5000 10 -5.10574190E+05 3 0.02 4 0.01 7 0.00
+SINQUAD 5000 1 -7.12672063E+03 3 0.00 2 0.02 8 0.00 4 0.00 3 0.00 5 0.00
+SINQUAD 5000 .1 -5.12198852E+02 3 0.00 2 0.00 11 0.00 7 0.00 3 0.00 2 0.00
+SPARSINE 5000 10 -2.64613490E+07 4 0.47 3 0.99 25 0.01
+SPARSINE 5000 1 -2.92330968E+06 4 0.00 4 0.44 6 0.96 3 1.02 12 0.00 11 0.00
+SPARSINE 5000 .1 -2.95128316E+05 4 0.00 4 0.00 8 0.67 5 0.71 12 0.00 5 0.00
+SPARSQUR 10000 10 -8.10327063E+06 15 2.93 4 6.62 47 0.03
+SPARSQUR 10000 1 -1.19555581E+06 15 0.00 7 2.67 6 3.38 2 3.45 23 0.01 15 0.01
+SPARSQUR 10000 .1 -1.23655088E+05 15 0.00 7 0.00 8 3.37 4 3.47 23 0.01 7 0.00
+SPMSRTLS 4999 10 -1.27251641E+03 5 0.01 5 0.01 73 0.02
+SPMSRTLS 4999 1 -8.16740777E+01 7 0.00 7 0.01 8 0.01 3 0.01 36 0.01 21 0.00
+SPMSRTLS 4999 .1 -7.77805138E+00 7 0.00 7 0.00 11 0.01 6 0.01 36 0.01 10 0.00
+[…]
+SROSENBR 5000 1 -1.15690489E+02 9 0.00 9 0.00 7 0.00 4 0.00 3 0.00 5 0.00
+SROSENBR 5000 1 -1.15690489E+02 7 0.00 9 0.00 8 0.00 7 0.00 3 0.00 2 0.00
+SSBRYBND 5000 10 -5.94528641E+08 18 0.01 12 0.03 159 0.04
+SSBRYBND 5000 1 -5.98063290E+06 18 0.00 17 0.01 15 0.02 12 0.02 79 0.02 145 0.04
+SSBRYBND 5000 .1 -7.28753976E+04 18 0.00 17 0.00 19 0.02 16 0.02 79 0.02 72 0.02
+SSCOSINE 5000 10 -5.13393463E+07 13 0.01 14 0.01 163 0.01
+SSCOSINE 5000 1 -5.13429696E+05 13 0.00 12 0.01 17 0.01 16 0.01 81 0.01 155 0.01
+SSCOSINE 5000 .1 -5.15840821E+03 13 0.00 12 0.00 20 0.01 19 0.01 81 0.01 77 0.01
+TOINTGSS 5000 10 -4.14177394E+03 5 0.01 3 0.02 15 0.00
+TOINTGSS 5000 1 -4.23179011E+02 5 0.00 4 0.00 5 0.01 2 0.01 7 0.00 9 0.00
+TOINTGSS 5000 .1 -4.24079188E+01 5 0.00 4 0.00 7 0.01 4 0.01 7 0.00 4 0.00
+TQUARTIC 5000 10 -2.37420750E-01 1 0.02 4 0.01 3 0.00
+TQUARTIC 5000 1 -2.91590249E-02 1 0.00 1 0.01 7 0.00 3 0.00 1 0.00 3 0.00
+TQUARTIC 5000 .1 -6.55745471E-03 1 0.00 1 0.00 10 0.00 6 0.01 1 0.00 1 0.00
+TRIDIA 10000 10 -1.08067135E+07 15 0.01 4 0.02 37 0.00
+TRIDIA 10000 1 -1.14762126E+06 15 0.00 7 0.01 7 0.01 3 0.02 18 0.00 15 0.00
+TRIDIA 10000 .1 -1.15438160E+05 15 0.00 7 0.00 9 0.01 5 0.01 18 0.00 7 0.00
+VAREIGVL 5000 10 -9.00927260E+04 4 2.09 3 5.49 13 0.00
+VAREIGVL 5000 1 -1.03187418E+04 4 0.00 3 1.95 6 4.97 3 5.18 6 0.00 9 0.00
+VAREIGVL 5000 .1 -1.04528079E+03 4 0.00 3 0.00 9 4.99 6 5.15 6 0.00 4 0.00
+WOODS 4000 10 -4.64705754E+06 2 0.00 3 0.00 7 0.00
+WOODS 4000 1 -5.13132992E+05 2 0.00 2 0.00 6 0.00 3 0.00 3 0.00 7 0.00
+WOODS 4000 .1 -5.17983606E+04 2 0.00 2 0.00 8 0.00 5 0.00 3 0.00 3 0.00
+YATP1LS 10200 10 -2.40174770E+07 1 4.28 3 10.85 3 0.01
+YATP1LS 10200 1 -2.47600401E+06 1 0.00 1 4.04 6 10.67 3 11.27 1 0.00 3 0.00
+YATP1LS 10200 .1 -2.48349990E+05 1 0.00 1 0.00 8 7.17 5 7.71 1 0.00 1 0.00
+YATP2LS 10200 10 -5.23502800E+05 1 6.74 3 11.27 3 0.00
+YATP2LS 10200 1 -6.17000891E+04 1 0.00 1 6.03 6 11.29 3 11.87 1 0.00 3 0.00
+YATP2LS 10200 .1 -6.26350766E+03 1 0.00 1 0.00 8 7.82 5 8.22 1 0.00 1 0.00
+Table D.1: Subproblem solves for CUTEst examples: Complete results (continued)